A Complete Bibliography of Computer Graphics Forum: 2020–2029

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
             WWW URL: http://www.math.utah.edu/~beebe/

02 June 2023 Version 3.01

Title word cross-reference

1 [FMKK21]. 2 [BBB22, EBB23, HS20, MOW22, MRWE20, OPP^{+}21, RKW20, RGR20, RCDB23, SW22, SMM22, VLLA22, WWW20, YLC^{+}22, ZWW^{+}21]. 3 [BCOM^{+}22, BTCL22, BYK^{+}21, BGMC22, BBB22, CETC22, CSQ^{+}22, CSAN23, CZG^{+}22b, CRW^{+}20, CL21, CXW^{+}21, CLC^{+}22, CL22b, DYDZ22, DLX^{+}22, FLH^{+}20a, FAG^{+}20, FRH^{+}21, GGMP21, GVF^{+}21, GCZ^{+}20, GLLL20, GCLB21, GRGH20, HWDL21, HTT^{+}20, HS21, HLJ^{+}20, HZSQ20, JJZ^{+}22, KHW^{+}22, KFS^{+}22, LSP^{+}21, LTY^{+}22, LLDT20, LAK20, LPW^{+}20b, LCI^{+}21, LSSW22, LFD^{+}20, LHG^{+}23, MCT21, MMRC20, MRWE20, MMC^{+}22, NYB^{+}20, ONH^{+}22, PMG^{+}20, PCDS20, QZL^{+}20, RSB^{+}21, SW22, SLL20, SWB^{+}20, TKR^{+}23, TCL20, WTO^{+}20, WYM^{+}20, WXL^{+}21, WYA^{+}20, WBB^{+}20, WS20, YT22, YTK23, ZLWT22, ZWM^{+}20]. L_{∞} [BBB22]. S^{2} [MBB22]. n [NBH20]. Z [KK21].

-Equivarianceg [PZL^{+}22]. -Pose [PZL^{+}22]. -Thickness [KK21]. -vector [NBH20].

2010s [vTP21]. 2019 editorial v2 [HB20].

3D [PLL^{+}20]. 3D-QR-Codes [PLL^{+}20].

4-axis [NTM^{+}21]. 4-Space [LZ22].

6D [PZL^{+}22].

A-ULMPM [SXHA22]. Aberrations [CK21]. Abstract [KFS^{+}22, YPZ^{+}22]. Abstraction [CSB^{+}20]. accelerate [JG21]. accelerated [vdRE21]. Accelerating [GZL^{+}20a, QMSL20]. Acceleration [OPY^{+}20, RSB^{+}21, SMOW20]. accessibility [EBM22]. Accessible [KJRK21, ZLL^{+}22b, EBM22]. Accuracy [RSGP21]. Accurate [KC22, WBB^{+}20]. Acquiring [FJM^{+}20]. Across [OM20]. Action [ZYP^{+}21]. Activations [RCPW21]. Actuated [JSVB22]. Adaptation [CHC^{+}23, II22, KC22, TKE21, WZX^{+}22]. Adapted [PKO22]. Adaptive [BP20, HMS^{+}20, LFS21, LSY21, SPK23, SYG20, SSKEA21, WH20, WMU^{+}20, WCU^{+}20, ZLL20, ZHHK22, ADP20]. Adaptively [SXHA22]. Additional [Csé23]. Additive [SMS^{+}20]. Adjacency [BSU^{+}22]. Adjustable [WZB20]. Adjustment [GOG21]. ADMM [OPY^{+}20]. Advancements [vTP21]. Advances [TTM^{+}22]. Advection [GHMR^{+}20, OL21]. Adversarial [GWM22, MCBR20, MA23, SWW^{+}22a, WWZ^{+}22, ZLY^{+}20, ZK21a]. Aerial [SLJ^{+}20, SXZS21]. aerodynamic [RCDB23]. Aesthetics [ABK^{+}20, CL22b]. Affective [BSB^{+}20]. Agent [JH22]. Agglomerations [QLL^{+}22]. Aggregation [NQHH20]. Agnostic [TST^{+}22]. Agonizing [AMA22b]. Air [QLL^{+}22, ZWLY22, ZWVL22]. AirLens [QLL^{+}22]. Algebra [MWS^{+}20b]. Algebraic [McN21]. Algorithm [WM22, FCD20]. Algorithms [BDK^{+}22, GGS21, SR20, vTP21, DPB^{+}22a]. Aliased [YVK^{+}23]. aliasing [Özt20]. Alignment [LWL^{+}20, MVNT21, ZJA21]. ALLSTEPS [XLKvdP20]. along [Ano23d]. Alternating [BSH22]. Ambient [IIHG^{+}20, PHP22]. Amortized [NMN20]. Analogies [LJ21]. analogue [IWH23]. Analogy [BMBW21, WMS^{+}23]. analyse [IWH23]. Analyses [Ano22b, ML22]. Analysis [ANI^{+}20, BTB^{+}22, BDL^{+}20, CRZ^{+}22, ELW^{+}22, FSS^{+}21, FMKK21, FSK^{+}22, GCL22, HLJ^{+}22, HL22, KSHL20, KMP^{+}21, LMBM20, LXYM20, LHX^{+}22, MTL^{+}21, McN21, MRKEA22, MRSF20, dQNSVE20, SJS20, SHR20, SFMS^{+}21, SLB^{+}21, SSC^{+}20, SFB20, TCG^{+}22, XOW^{+}20, BS22, PFCR20]. Analytic [Csé23, MMBP22, d’E21]. analytical [CLH^{+}20]. Analytics [ABK^{+}20, BSP^{+}22, CMKK21, CPK23, DPR^{+}21, FSS^{+}21, FAP^{+}22, HNWC23, HDP^{+}21, KFS^{+}22, LBB^{+}23, LLT^{+}20b, MFH^{+}22, dQNSVE20, PMCEA^{+}22, PL20, STX^{+}22, SAK^{+}21, USK21, XXeL^{+}20, ZWLY22, ZWVL22].

Analyzing [CM21b, SLJ^{+}20]. Anatomical [GFZ^{+}20, RGW20, SKRW22]. anatomy [FCD20]. Ancient [MMP20]. Anderson [OPY^{+}20]. Animal [CSB^{+}20]. Animate [PYAP22]. Animated [LTH20, ROCP20, TLLS20, WGH^{+}21]. Animating [SW20]. Animation [AAC22, BSS20, CZG^{+}22a, ECC^{+}20, EBB23, KADM22, KAK^{+}22, MHL^{+}22, NRS22, RTK^{+}21, Ron21, SPDM20, SW22, YY20, ZXS^{+}21, ZSS22, LLDT20]. Animations [ALTT22, CWW^{+}22, GZL^{+}20b, LSK23, PCS21]. Anisotropic [CLR^{+}20, EMAK20, FVMK21, LHL^{+}21, LBK20, SW20]. Annotated [HSvO^{+}22, LFH^{+}22, BAMG21]. Annotations [GCZ^{+}20, HCS20, PCDS20]. Annular [WTO^{+}20]. Answering [HKM22]. Anti [Özt20, YVK^{+}23]. Anti-Aliased [YVK^{+}23]. Anti-aliasing [Özt20]. Antialiasing [YLS20]. Anticipation [BLC20]. antigen [Ano23d]. Any [ZWCM21]. Anyway [McN21]. Aortic [MRB^{+}21]. Apollonius [WYM^{+}20]. Appearance [DLC^{+}22, HHG^{+}22, SL23, YDPG21]. Application [SMS^{+}20, VLLA22]. Applications [AJLN22, BPD20, BGA^{+}22, CBL^{+}20, YMS^{+}21]. apply [DPB^{+}22a]. Applying [GMOB22]. Approach [AL21, GMOB22, HNWC23, LLT^{+}20b, MDBH22, dQNSVE20, PMC^{+}22b, SSX^{+}21, SHV21, TCT22, WCM22, BAMG21]. Approximating [HT20, LSK21]. Approximation [BVHSH21, HYW^{+}22, LWH^{+}20, MMO^{+}21, Mit22]. approximations [LBFS21]. Arbitrarily [wKH20]. Arbitrarily-shaped [wKH20]. Arbitrary [BTCL22, SRC^{+}20, ZLL^{+}22a]. Architecture [BGA^{+}22, ZWM^{+}20]. Areas [RLHS20a]. Argument [KCW^{+}21]. Arrays [QFWH21]. Arrows [BDK^{+}22]. Art [AMA^{+}22a, BGA^{+}22, BYK^{+}21, BCD20, BYH^{+}20, CMG21, CMJ^{+}20, CNS22, HHS20, HKM22, II22, LBB^{+}23, MPE23, TFT^{+}20, TKR^{+}23, WSP21, WL22, YVK^{+}23, PMG^{+}20]. artefact [IWH23]. Artifacts [FAG^{+}20]. Artist [KHW^{+}22]. Artistic [CSG21, KFDA21, PCS21]. As-Few-As-Possible [CZGM20]. Assemblies [LHX^{+}22, WSP21]. Assembly [DZN^{+}21, QLZ^{+}20]. Assessing [JLB22]. Assessment [NBPM20, SBMS23]. Assignment [BHJS23]. Assignments [HTI22]. Assisted [NBPM20]. Associated [CLL^{+}22]. Astrophysics [LYA^{+}21]. Asynchronous [KMA20, LXYM20]. At-Most-Hexa [BTL22]. Atmosphere [Hil20]. Atmospheric [SGAG22]. Attention [RLHS20b]. Attentive [YLL^{+}21, ZYP^{+}21]. Attitude [LPW20a]. Attribute [PSK^{+}20, XL20]. Attributes [DLC^{+}22, LSK23, ME23, SL23]. Audio [ALTT22]. Audio-driven [ALTT22]. Auditory [WJK22]. Augmented [BC20, LY20, SGBI22, KP22]. Augmenting [MFH^{+}22, PSK^{+}20]. Authoring [GPM^{+}22, LBC^{+}22, TLLS20, TZP^{+}22]. AutoClips [SSX^{+}21]. autocovariance [LSD23]. Automated [AJLN22, BLC20]. Automatic [GKG^{+}22, KF21, LMF^{+}20, LWH^{+}20, MMRC20, NCB^{+}21, NTM^{+}21, PMG^{+}20, SBB^{+}22, SSX^{+}21, ZOM^{+}22]. Autonomous [CBL^{+}20, SLJ^{+}20]. auxetic [BHM21]. Auxiliary [FKL^{+}21]. Avatar [CSAN23, CHC^{+}23]. Avatars [PYAP22, ROCP20]. Averaging [GvdVS23]. Aware [BGGSSG20, CSG21, CTJ^{+}20, DMT21, FGZ^{+}22, GGS21, LSK23, LLDT20, LADH22, LHG^{+}23, NZCS21, RSB^{+}21, YTK23, ZLY^{+}20, ZPH^{+}23, CCG^{+}22, DX22, EHL21, GSWS21, HZSQ20, JHC^{+}20, TACS21, YJS^{+}22]. Axes [Ano22b, ML22, RSLSRÁ21]. Axis [DLX^{+}22, LBKP21, ZDS20, NTM^{+}21].

Band [LWH^{+}20, MCBR20, OP22]. Band-Limited [LWH^{+}20, MCBR20]. Baraff [Kim20]. Baraff-Witkin [Kim20]. BareSkinNet [YT22]. Barrio [TCG^{+}22]. Bas [WSN^{+}20, ZWW^{+}21]. Bas-Relief [WSN^{+}20, ZWW^{+}21]. Based [ABPP23, AAC20, AWKK21, BSS20, BC20, CDMH^{+}22, CLL^{+}20, CYD^{+}22, CMG21, CTJ^{+}20, CMM^{+}20, GK22, GMOB22, GCLB21, HZ22, HZSQ20, JSL22, JZGF20, KJS^{+}21, KPLD21, LTY^{+}22, MAA20, MHL^{+}22, MMC^{+}20, MMCJ22, OM20, OPY^{+}20, PHP22, QZL^{+}20, QZLC22, RDG20, RMSQ20, SAK20, TAA^{+}21, TK21, WGT^{+}22, WMU^{+}20, WBL20, WSN^{+}20, XHZ^{+}20, ZCC22, ZLWT22, AMA22b, CSDD20, CSSI21, CvTZ^{+}22, DLC^{+}22, DKVV23, FLH^{+}20b, FTA22, GVF^{+}21, GFH^{+}23, GvdVS23, GZG^{+}22, HXA21, HTT^{+}20, HTI22, HHH22, HJL^{+}20, HPKI22, JJZ^{+}22, KQL^{+}21, KUKH23, KPKH20, KMP^{+}21, KFDA21, KMTM22, LWY^{+}21, LCI^{+}21, LSE21, MCC23, NZL^{+}21, NHSS21, SPDM20, TL21, TKAS23, TWF22, TBSB20, WCSH21, XK21, YYH20, YBRP21, YZ22, YUIH21, YY20, ZXS^{+}21, ZLL^{+}22a, ZSY^{+}22, ADP20]. Bases [AL21, MN20, PKO22]. Basis [GAD^{+}20, SS21]. Bayesian [GHYZ20]. Behavior [CSB^{+}20, SYHL21]. Behaviors [CLL^{+}22, CvTZ^{+}22]. Belief [MCK^{+}22]. Benchmark [LSP^{+}21, OPP^{+}21]. Bending [PRCO20]. Between [DCMO22]. Beyond [Cav21, Csé23, XTS^{+}22]. Bézier [KMC22, YLC^{+}22]. BGP [USK21]. bi [KP22]. bi-cubic [KP22]. Bicycle [DFV21]. Bidirectional [CSQ^{+}22, FJF20, GGS21, JHD20, NID20, QSBJ22]. Big [MHT^{+}20]. Bijective [YSLF20]. Billion [SMOW20]. Bin [AWKK21]. Binary [CDK20, GZG^{+}22, LL20]. Biological [SKRW22]. Bipartite [PPF23]. Biphasic [TCT22]. Bisection [JC22]. Blend [CSG21]. Blend-Aware [CSG21]. Blending [IA21, KK21]. Blendshape [CZGM20]. Blind [PBM^{+}22]. Blood [EMKL21]. Blue [vOSRR21]. Bodies [PT23, MMRC20]. Body [ARC22, ANK^{+}23, CARK20, GDG^{+}23, JPK20, MMC^{+}20, TKE21, WHZ^{+}23, ZWG^{+}21, SCCZ21, YKL21, YLHX22]. Bomb [AWB20]. Bombalytics [AWB20]. Book [BYK^{+}21, CRW^{+}20, FLH^{+}20a, FAG^{+}20, GCZ^{+}20, GLLL20, GRGH20, HTT^{+}20, HS20, HLJ^{+}20, HZSQ20, LLDT20, LAK20, LPW^{+}20b, LFD^{+}20, MMRC20, MRWE20, NYB^{+}20, PMG^{+}20, PCDS20, QZL^{+}20, RKW20, RGR20, SLL20, SWB^{+}20, TCL20, WWW20, WTO^{+}20, WYM^{+}20, WYA^{+}20, WBB^{+}20, WCSH21, WS20, ZWM^{+}20, LZ22]. Boosting [ZGP^{+}20]. Bootstrap [GvdVS23]. Bound [ZSL^{+}22]. Boundaries [LAK20]. Boundary [CLH^{+}20, JHC^{+}20, JSL22, LCG^{+}20, SGL23, SBH22, VRBT21, ZLL^{+}22a]. Boundary-aware [JHC^{+}20]. Boundary-based [ZLL^{+}22a]. Bounding [MOB^{+}21, VEPG23]. Bounds [GGPP20, MYT^{+}21, UKPW21]. Box [VEPG23]. Boxer [GBYH20]. Bracket [WBM21]. Brain [SMRB22, TCG^{+}22]. Branch [WLG22]. BRDF [BP20, CSDD20, HGC^{+}20, LBFS21, Pet21, SRRW21, ZSWL21, d’E21]. BRDFs [LFR23]. Breast [NBPM20]. Bridging [WYZ^{+}22]. Brittle [CMSK20]. Broad [SR20]. Broad-Phase [SR20]. Broadmark [SR20]. Browser [LSE21]. Browsing [MZAD^{+}20]. BTFs [RGJW20]. Bubble [Ano22a]. Builder [XSZX22]. Building [AACF22, HWC21]. Buildings [OM20, WHPG22]. Bundles [DCMO22]. Buoyancy [BPD20]. Business [LLT^{+}20b]. BVH [TDDB23].

Cable [LMF^{+}20]. Cable-Tensioned [LMF^{+}20]. Caches [HSG^{+}22]. CAD [CSQ^{+}22]. Cages [CTL^{+}20]. Calibration [ANK^{+}23, WBB^{+}20]. Camera [BLC20, HMGU20, HACOG22, KKN^{+}21, LHG^{+}23, LPW^{+}20b]. Cameras [CLCO^{+}22]. Can [BRK^{+}22]. Cancer [NBPM20]. Canis [GZL^{+}20b]. Can’t [Hei20]. Capture [BRMO22, KADM22, LHG^{+}23, MRR^{+}22, ZM22]. captured [ÇBSM23]. Capturing [AKA^{+}20, FVMK21, RPPMCG23]. car [RCDB23]. Cardiovascular [EMML22]. Care [BSP^{+}22]. Carlo [FWHB21, FFJ22, GWH20, GE21, LWY^{+}21, LZBD21, SSC^{+}20, ZMV^{+}21]. Carotid [EMKL21]. Carpentry [NZCS21]. Cartograms [McN21]. Cartoonization [LWZ^{+}20]. Case [LHX^{+}22, ZZF^{+}22, ZWVL22]. Cast [GRP22, NRS22]. Categorical [GCH^{+}21]. Catmull [DV21]. Caustics [JC22]. Cave [PGP^{+}21]. CDF [Hei20]. Cell [HTI22, Ano23d]. Cellular [SMS^{+}20]. Center [ZJZ^{+}22]. Centered [SEAG^{+}21]. Central [Csé23]. Chain [GWH20]. Challenges [KJRK21]. Channel [GLLL20]. Character [ECC^{+}20, GWE^{+}20, KAK^{+}22, NRS22, YY20, ZXS^{+}21]. Character-Scene [ZXS^{+}21]. Characteristic [NA22]. Characterization [DPR^{+}21]. Characters [ARC22, FRH^{+}21, FTA22]. Chart [GZL^{+}20b, HCS20, HKM22]. Chartability [EBM22]. Charts [BDL^{+}20, TB21, WJK22]. Chebyshev [LGW^{+}22, LLZ^{+}20]. Checkpoint [KF21]. Chinese [LG22, TL21]. Chocks [Mit22]. Choices [NQHH20]. Cinematography [AJLN22]. Circles [BDF^{+}22]. Circuit [GWB^{+}21]. Circular [XFS^{+}20]. Cities [MMP20]. Civic [BJSM22]. CLA [ZLY^{+}20]. CLA-GAN [ZLY^{+}20]. Clamped [LWH^{+}20]. Clark [DV21]. Class [EK20, GMH22]. Class- [EK20]. Classification [BNR20, DXG^{+}20, MWZ^{+}20, WTO^{+}20]. Classifier [BNR20, GBYH20, GCDT22]. Classifier-Guided [BNR20]. Cleanup [MHLH22]. Clip [AJLN22, YVK^{+}23]. Clip-Art [YVK^{+}23]. Clipart [SLS^{+}21]. ClipFlip [SLS^{+}21]. Close [GMH22]. Closed [BCOM^{+}22, IA21]. Closure [WGT^{+}22]. Cloth [CCR^{+}20, CMM^{+}20, Kim20, LTY^{+}22, PRCO20, LTY^{+}22]. Clothing [JZGF20, LSK23]. Cloud [CLC^{+}22, CCG^{+}22, FLH^{+}20b, GZLL23, GKK^{+}20, HCL^{+}22, HXW^{+}22, LFH^{+}22, MWZ^{+}20, SDR^{+}22, SCZ21, SXZS21, YZ22]. Clouds [BGGSSG20, GCLB21, GZG^{+}22, HLJ^{+}20, JG21, LMBM20, MHG^{+}22, QZLC22, RLG^{+}20, SOW20, SKW21, LCBM21, LAK20, QZL^{+}20]. Cluster [BZP^{+}20]. Clustering [NZL^{+}21, TBL^{+}22, HLJ^{+}20]. Clusters [GCH^{+}21]. ClusterSets [GCH^{+}21]. CNN [SLB^{+}21]. Co [LB20, NKA^{+}23, SSKEA21, SLL20]. Co-Adaptive [SSKEA21]. Co-creating [LB20]. Co-Speech [NKA^{+}23]. Co-training [SLL20]. Coaching [PMB^{+}23]. Coarse [HWJ22, ZGP^{+}20]. Coarse-to-Fine [HWJ22]. Coarsening [HPO22]. Coatings [FVMK21]. Code [PHP22, TSK^{+}20]. Coded [AFTH21]. Codes [PLL^{+}20]. Cognitive [JH22]. Coherence [PMC^{+}22b]. Coherent [GVF^{+}21, HSS21, PGP^{+}21]. Cohort [ANI^{+}20]. Collaboration [AWB20]. Collaborative [EBD21, LXYM20, VRBT21]. Collection [CBC20]. Collections [Fre22, MRKEA22, YGY^{+}21]. Collective [CSB^{+}20]. Collision [CDK20, CARK20, SR20, WFJ^{+}22]. Collisions [WTB^{+}21]. Collocated [GHMR^{+}20]. Color [GSDT22, RSGP21, RSB^{+}21, WRZ20]. Color-mapped [GSDT22]. Colored [FYFL20]. Colorization [AMT20, CMG21]. Colormaps [NCB^{+}21]. Colorspaces [NCB^{+}21]. Colour [GCLB21, QFWH21]. Colouring [PMC22a]. Colourization [XHZ^{+}20]. Combination [CDMH^{+}22, LTH20]. Combinatorial [BSK21, ZTZC20]. Combined [CYD^{+}22, BP21]. ComBiNet [PPF23]. Combining [PYAP22]. Comets [SLJ^{+}20]. CommAID [FSS^{+}21]. Common [MWM^{+}22]. Communication [FSS^{+}21, KMH21]. Compact [NSP^{+}21, ZM22, ZLW^{+}22]. Comparative [BDL^{+}20, LLLB22, NBPM20, PMB^{+}23]. Comparison [GBYH20, KMTH22, MvdEV22, PPF23, TCG^{+}22, VSC^{+}20, YMS^{+}21]. Compassion [MAS22]. Compensation [SLCL22]. Competition [AWB20]. Completion [HCL^{+}22, LSSW22, WH20]. Complex [BGMC22, DCMO22, GGN20, MHT^{+}20]. Complexity [LCI^{+}21]. Components [DBP20]. Composing [HPM^{+}20]. Compositing [GVF^{+}21, LFS21]. Composition [LADH22]. Compositional [ZMV^{+}21]. Compound [MRSF20]. Comprehensive [NKA^{+}23, Özt20]. Compressed [BGT20, CBE20, ME23]. Compressive [HMGU20, LJLB21]. Computation [MLP21, MHT^{+}20, SYKE20, TBK21, WYM^{+}20, ZSL^{+}22]. Computational [AMA^{+}22a, JSVB22, VZTH22, WSP21, XFS^{+}20]. Computations [MO23, WH20]. Compute [SKW21]. Computer [BD23, KAC^{+}21, KBST22, RGJ^{+}23, Ron21]. Computing [BDF^{+}22, FJF20, XTS^{+}22]. ComVis [PMB^{+}23]. ComVis-Sail [PMB^{+}23]. Conceiving [Cav21]. Concentric [BDF^{+}22]. ConceptGraph [KSHL20]. Conditional [GMH22]. Conditions [CLH^{+}20, LCG^{+}20]. Cone [TCT22]. Cones [EMAK20]. Confidence [HZSQ20]. Conflict [MLD^{+}21]. Conflicting [LXYM20]. Conformal [DCMO22]. Conforming [WWW20]. connected [DR22]. Connections [RLHS20b]. Connectivity [MWS^{+}20b, DKVV23]. Connectomics [BTB^{+}22]. Consensus [CLCO^{+}22, GvdVS23]. Consensus-based [GvdVS23]. Conservative [UKPW21]. Considerations [EBD21]. Consistency [LCP^{+}23]. Consistent [CM21a, HRWO20]. Constitutive [WDK^{+}20]. Constrained [LCK21a, TWF22, WZB20, ZWG^{+}22]. Constraints [KCRH20, SMM22, ZCC22]. Constructed [MWS^{+}20a]. Constructing [BBB22]. Construction [AL21, DXG^{+}20, HCS20, TDDB23, WMS^{+}23, ZTZC20, ZYCF20, BHM21, JG21]. Consumer [PYAP22]. Consumer-Grade [PYAP22]. Contact [MHLH22, PT23, RLR^{+}20, STX^{+}22, ZCC22]. Containment [KOYSS21]. Content [BC20, CLL^{+}22, CSSI21]. Content-based [CSSI21]. Context [CDMH^{+}22, CTJ^{+}20, KMTM22, ZLY^{+}20]. Context-Aware [CTJ^{+}20]. Context-based [KMTM22]. Contextual [CLC^{+}22]. Contextualization [SKB^{+}22]. Contextualized [SSKEA21]. Continuous [NCB^{+}21, WFJ^{+}22]. Contour [LWL^{+}20]. Contours [PMC22a]. Contrast [NBPM20, YJS^{+}22, WCS^{+}22]. Contrast-Enhanced [NBPM20]. Contrastive [WWF^{+}22]. Control [BSH22, BLC20, GAM^{+}21, HWJ22, HACOG22, LFD^{+}20, SPDM20, TKAS23, XK21, ZWLY22, ZWVL22]. Controllable [AHKB20, CCOCM21, End22, WKA^{+}20]. Controlled [Lev22, ZSS22, BCOM^{+}22]. Controllers [PWBD21, YY20]. Controlling [HHG^{+}22]. Conversation [HDP^{+}21]. Conversations [YYH20]. Convolution [MRCK21]. Convolutional [CMM^{+}20, HW22, SRC^{+}20, VSGC20, WTO^{+}20]. Convolutions [KPKH20]. Cooperative [SRK21]. Coordinates [BvOdK^{+}22, BZP^{+}20, ML20, RSLSRÁ21]. Core [SOW20, CW22]. Corners [CRH^{+}21]. CorpusVis [MRKEA22]. Correctable [HTI22]. corrected [LRTC20]. Correction [BNR20, GWB^{+}21]. Corrector [TKAS23]. Correlated [CMSK20, SSC^{+}20]. Correlation [EHL21, GGS21, SYG20]. Correlation-Aware [GGS21]. Correspondence [CBC20, LXY^{+}22]. Correspondences [LCP^{+}23]. Cosines [DGJ^{+}20]. Cosserat [WCU^{+}20, ZLW^{+}22]. Cost [Csé23, LBK20, SCHC20]. Cost-Effective [SCHC20]. Counterfactual [SGL23]. Counts [OM20]. Coupled [CTL^{+}20, HXA21]. Coupling [SW22]. Course [BJSM22]. Coverage [DLX^{+}22]. CPU [WMU^{+}20]. Create [AKA^{+}20]. creating [LB20]. Creation [BC20, VRBT21]. Creative [GAM^{+}21]. Creature [LLL^{+}21]. Crime [dQNSVE20]. Critical [BJSM22]. Cross [FAP^{+}22, GCZ^{+}20]. Cross-Dimension [GCZ^{+}20]. Cross-Virtuality [FAP^{+}22]. Crowd [CDMH^{+}22, CvTZ^{+}22, MWM^{+}22, MBA22, vTP21]. Crowds [LBC^{+}22]. CSG [TK21]. CubeGAN [MA23]. cubes [HWC21]. cubic [KP22]. Cues [SWB^{+}20]. Culling [LJSL21, UKPW21]. Cultural [BBBD20]. Curriculum [XLKvdP20, ZCH^{+}23]. Curriculum-driven [XLKvdP20]. Cursive [TL21]. Curvature [BSH22, YUIH21, LRTC20]. Curve [CW22, Hei20, LCI^{+}21, MOW22, OPP^{+}21, HLJ^{+}20]. Curved [HHH^{+}21, XCC22]. Curves [BSH22, LCI^{+}21, YLC^{+}22, BCOM^{+}22, IA21]. Customizable [TCG^{+}22]. Customized [HXW^{+}22, YGY^{+}21]. Cut [Hei20, MLP21, ZYCF20]. Cutting [SWB^{+}20]. cVAEs [ALTT22]. CVFont [LG22]. Cyclostationary [LSD21].

D [BYK^{+}21, CRW^{+}20, FLH^{+}20a, FAG^{+}20, GCZ^{+}20, GLLL20, GRGH20, HTT^{+}20, HS20, HLJ^{+}20, HZSQ20, LLDT20, LAK20, LPW^{+}20b, LFD^{+}20, MMRC20, MRWE20, NYB^{+}20, PMG^{+}20, PCDS20, QZL^{+}20, RKW20, RGR20, SLL20, SWB^{+}20, TCL20, WWW20, WTO^{+}20, WYM^{+}20, WYA^{+}20, WBB^{+}20, WS20, ZWM^{+}20, BCOM^{+}22, BTCL22, BGMC22, BBB22, CETC22, CSQ^{+}22, CSAN23, CZG^{+}22b, CL21, CXW^{+}21, CLC^{+}22, CL22b, DYDZ22, DLX^{+}22, EBB23, FRH^{+}21, FMKK21, GGMP21, GVF^{+}21, GCLB21, HWDL21, HS21, JJZ^{+}22, KKN^{+}21, KHW^{+}22, KFS^{+}22, LSP^{+}21, LTY^{+}22, LCI^{+}21, LSSW22, LHG^{+}23, MCT21, MOW22, MMC^{+}22, OPP^{+}21, ONH^{+}22, RSB^{+}21, RCDB23, SW22, SMM22, TKR^{+}23, VLLA22, WXL^{+}21, WLNM21, YT22, YLC^{+}22, YTK23, ZWW^{+}21, ZLWT22]. DAGs [DSA20]. Daisen [SZM^{+}21]. damage [IWH23]. Damping [WDK^{+}20]. Dance [PWBD21]. Danmu [CLL^{+}22]. DanmuVis [CLL^{+}22]. Dashboard [DWH^{+}22]. Data [ANI^{+}20, ADP20, Ano22b, ATL^{+}23, AL21, BKL21, BvOdK^{+}22, BCD20, BBBD20, BDL^{+}20, BSP^{+}22, CWW^{+}22, CM20, CM21b, DSMA^{+}21, Fre22, GZL^{+}20a, GZL^{+}20b, GCH^{+}21, HNWC23, JM22, JZGF20, KKM^{+}20, LCB21, LPH^{+}20, LLLB22, LLC^{+}20, LPW20a, LLDT20, MZAD^{+}20, MCK^{+}22, MWM^{+}22, ML20, ML22, MLD^{+}21, NQHH20, NKA^{+}23, PRD21, RSLSRÁ21, SLJ^{+}20, SGOC20, SFMS^{+}21, SSX^{+}21, TC21, TLLS20, WMU^{+}20, WJK22, WCM22, XTL^{+}21, YGY^{+}21, ZLL^{+}22b, BAMG21]. Data-Driven [AL21, GZL^{+}20a, GZL^{+}20b, JZGF20, LLC^{+}20, LLDT20, NKA^{+}23, LCB21, SGOC20]. Data-type [Ano22b, ML22]. Dataset [AJLN22, LSP^{+}21, WBM21, LBFS21]. De-lighting [YT22]. De-makeup [YT22]. Deblurring [SLCL22, YVK^{+}23]. Decade [CSK^{+}23]. Decision [SGL23, ZWLY22, ZWVL22]. Decision-Making [ZWVL22]. Decisions [JVW20]. Decomposed [ECC^{+}20]. Decomposition [BGMC22, FAG^{+}20, SRM^{+}21, SS21, WLG22, YTK23]. Decomposition-Independent [WLG22]. Decontouring [JZZ^{+}20]. Decoupling [ZSY^{+}22]. Decoupling-based [ZSY^{+}22]. Decremental [YPZ^{+}22]. Deep [AFTH21, AAC22, BRMO22, DX22, DBP20, EGH21, GWE^{+}20, GRGH20, HWDL21, HGC^{+}20, HSG^{+}22, HL22, JVW20, JZZ^{+}20, JZN^{+}22, KJS^{+}21, KHW^{+}22, KJB^{+}22, LBB^{+}23, LPH^{+}20, LSK23, LG22, MHL^{+}22, MHLH22, NODI21, QZL^{+}20, RCPW21, SLL20, YDPG21, ZMV^{+}21, ZWCM21, ZA21, ZXJ^{+}20]. DeepBRDF [HGC^{+}20]. Deeper [LCB21]. Definitions [MDBH22]. Deformable [GGMP21, LLJ^{+}20, MMC^{+}22, WDY^{+}22]. Deformables [SBH22]. Deformation [CMM^{+}20, CTL^{+}20, LSK21, LTY^{+}22, LHZ^{+}22, MFJ^{+}21, SHR20, TKAS23, ZLL20]. Deformation-based [TKAS23]. Deformations [TRPO21]. Deforming [Fre21]. Defragmentation [MCT21]. Delaunay [PMC22a, SLMV21, XABR21]. Delaunay-Graph [SLMV21]. Demand [GRGH20]. Demonstrations [HACOG22]. Demosaicking [QFWH21]. Dendritic [COKE20]. Denoise [RLG^{+}20]. Denoising [FWHB21, FFJ22, HMS^{+}20, HXW^{+}22, JZN^{+}22, LWY^{+}21, MH20, WCR^{+}23, ZMV^{+}21, ZOM^{+}22, ZSWL21]. Dense [FLH^{+}20a, Fre21, MO23, MB20, RLG^{+}20, TBSB20, TB21, TKR^{+}23, XHZ^{+}20, ZQW^{+}20]. Densities [KHW^{+}22, MTL^{+}21]. Density [YUIH21, ZXJ^{+}20]. Density-based [YUIH21]. Dental [ABK^{+}20]. Dependencies [EST20]. Dependent [BYH^{+}20, HS20, HS21, VSC^{+}20, VGdS^{+}20]. Depiction [MVNT21].

Depixelization [MPE23]. Depth [ANK^{+}23, FGZ^{+}22, HZ22, LADH22, NSP^{+}21, RPPMCG23, SWB^{+}20, ZK23]. Depth-Aware [FGZ^{+}22, LADH22]. Derivatives [Csé23, KMP^{+}21]. Descent [MEM^{+}20]. Description [MRCK21]. Descriptor [LHL^{+}21, ZWM^{+}20]. Descriptors [MRB^{+}21, NODI21, YMS^{+}21]. Design [AMA^{+}22a, CLR^{+}20, CMMK20, EBD21, EMJ^{+}22, FYFL20, JSVB22, KQL^{+}21, KCW^{+}21, KJRK21, KMH21, LMF^{+}20, LNRH20, MCK^{+}22, MAS22, SYPB21, SLS^{+}21, TC21, TLLS20, TZP^{+}22, UPE22, VZTH22, VRBT21, WSP21, WNT^{+}20, XFS^{+}20, ZPH^{+}23, ZWLY22, RCDB23]. Designer [HJZ^{+}21]. Designing [BDL^{+}20, MWS^{+}20a, SLJ^{+}20, WHZ^{+}23]. Designs [LLC^{+}20]. Desktop [BC20]. Destination [TC21]. Detail [Fre22, LSK23, TBL^{+}22, ZWCM21, ZA21]. Detail-Aware [LSK23]. Detailed [KADM22, MMC^{+}20, NA22, SZM^{+}21]. Details [LWB^{+}21]. Detection [CDK20, JPK20, MHLH22, SR20, Wan20, WFJ^{+}22, WYZ^{+}22, WMS^{+}23, ZGP^{+}20, ZSY^{+}22]. Detector [ZQW^{+}20]. Deterministic [Mit22]. Developable [BVHSH21, WCSH21]. Developing [LYA^{+}21]. Devices [PYAP22]. DFR [WS20]. Diagnosis [MvdEV22]. Diagrams [BAR^{+}21, BBB22, LZ22, PSK^{+}20, WYM^{+}20]. Dice [RGW20]. did [BvOdK^{+}22]. Differences [CBC20, Csé23, LCO20]. Differentiable [ANK^{+}23, CSQ^{+}22, DBA^{+}21, DPB^{+}22b, GKG^{+}22, KJB^{+}22, LZBD21, RSP22, TST^{+}22, WQL^{+}20, WS20]. differential [LCBM21]. Differentiation [SBB^{+}22]. Diffraction [CCFM23, FJF20, BP21]. Diffuse [MMK^{+}22]. Diffusion [GvdVS23, HXA21, LFH^{+}22, MSY^{+}21, SHV21]. DiffusionPointLabel [LFH^{+}22]. Digital [GFZ^{+}20, LFD^{+}20, MFH^{+}22, MRKEA22, WHPG22]. Dilated [KPKH20]. Dimension [GCZ^{+}20]. Dimensional [Cav21, FIDF20, GZLL23, Sub21, XTL^{+}21]. Dimensionality [RKW20]. Dimensioned [BSU^{+}22]. Dimensions [BJSM22, Cav21, WJK22]. Direct [DBP20, KRF22, WHW22, ZFWW20]. Directed [BDK^{+}22]. Directing [Ron21]. Direction [DLLK21]. Directional [DPÖM22, PLL^{+}20]. Directionality [ZPH^{+}23]. Directionality-Aware [ZPH^{+}23]. Directions [HKM22]. Dirichlet [SWJG20]. Discontinuities [HZ22]. Discovering [LYA^{+}21]. Discovery [AACF22]. Discrete [BB23, LGW^{+}22, LLZ^{+}20, MLP21, RMWO21, WDH20, ZAB^{+}21, PFCR20]. Discretization [SWJG20, WCU^{+}20]. Discretizations [PRCO20]. Discriminative [GLLL20]. Disentangled [CZG^{+}22a, WDY^{+}22, WLT^{+}22]. Disentanglement [JLB22]. Disk [Mit22]. Displacement [CMSK20, FSK^{+}22]. Displacement-Correlated [CMSK20]. Display [AKA^{+}20, ROG20]. Displays [CM21b, RLR^{+}20, YJS^{+}22]. Dissections [MRB^{+}21]. Distance [BHJS23, SYKE20, TBK21, ZSL^{+}22]. Distances [GH21, WLG22]. Distant [CARK20]. Distinguished [HS20]. Distributed [QMSL20]. Distributions [BDL^{+}20, SHJD22]. Districts [LLT^{+}20b]. Divergence [JLDH20]. Divergence-free [JLDH20]. Diverse [PWBD21]. Diversifying [EK20]. DNA [Ano23d]. Do [LGS^{+}22]. DOFs [ZCC22].

Domain [ATL^{+}23, LWH^{+}20, TKE21, TACS21, WZX^{+}22, XSW^{+}20]. Domains [DR22, HHH^{+}21, YLC^{+}22, ZAB^{+}21, ZTZC20]. Dominated [WM22]. DONeRF [NSP^{+}21]. Double [WWRT21]. Douglas [OPY^{+}20]. Douglas-Rachford [OPY^{+}20]. Drape [RPPMCG23]. Draping [HPG^{+}23]. Drawing [AMA22b, JSL22, ZWW^{+}21]. Drawings [AMT20, NODI21, ZLL^{+}22a]. Dressi [TST^{+}22]. Dressing [FRH^{+}21]. Driven [AHKB20, AL21, BJG20, DPB^{+}22b, GZL^{+}20a, GZL^{+}20b, GDG^{+}23, JZGF20, LLC^{+}20, LLDT20, LJ21, LTH20, NKA^{+}23, RKL^{+}21, SWW^{+}22b, TB21, ZSL^{+}20, ZYP^{+}21, ALTT22, LCB21, SGOC20, XLKvdP20, WYA^{+}20]. Driving [CBL^{+}20]. DRLViz [JVW20]. Drone [AJLN22]. dual [ÇBSM23, MEM^{+}20]. dual-exposure [ÇBSM23]. Duality [VLLA22]. Ductile [ZHL^{+}20]. During [KSHL20, STX^{+}22]. Dynamic [AB20, ARC22, AMA22b, CDMH^{+}22, CL22a, CLCO^{+}22, HACOG22, JL21, KKN^{+}21, MMK^{+}22, MB20, MSK23, MWS^{+}20b, NBPM20, ONH^{+}22, PPF23, SHVR21, SBH22, TAA^{+}21, VCT21, WYZ^{+}22, ÇBSM23, RGR20]. Dynamical [BGG20]. Dynamics [CLL^{+}22, FSS^{+}21, MEM^{+}20, MWM^{+}22, MMC^{+}20, PCS21, RMSQ20, SGOC20, WGT^{+}22, WZB20, WBL20, WTB^{+}21, XXeL^{+}20, XK21].

Earth [LYA^{+}21]. ECHO [MRCK21]. Eclipses [SGAG22]. Ecological [ZWVL22]. Economic [ZSL^{+}22]. Edge [GWM22, HZSQ20]. Edge-Based [HZSQ20]. Edge-Enhanced [GWM22]. Edges [CRH^{+}21]. Edit [WLG22]. Editing [ABPP23, BSS20, CSQ^{+}22, DLC^{+}22, EK20, End22, FVMK21, LGW^{+}22, LADH22, ME23, NHSS21, SOL^{+}22, SRM^{+}21, SL23, ZXS^{+}21, ZFWW20]. Editorial [AH23, HB21, HA22]. Educating [LYA^{+}21]. Edutainment [RGW20, SKRW22]. Effect [FVMK21]. Effective [KK21, LXY^{+}22, SCHC20, SSV^{+}22, SSS20b, WZX^{+}22, PCDS20]. Effects [LHW21, NHSW20, NQHH20, SGAG22, ZWVL22]. Efficient [BMPP21, CK21, FIDF20, GGN20, GEE20, HSS21, HTFW23, HRWO20, KRF22, LHJ21, MO23, MDVP22, MFJ^{+}21, OL21, SYKE20, SXHA22, SWW^{+}22b, TACS21, YSLF20, ZLW^{+}22, LCBM21, MRWE20]. EGGS [TSK^{+}20]. Egocentric [SAK^{+}21, SYHL21]. EHR [WL22]. EL-GAN [GWM22]. Elastic [BBBD20, WDK^{+}20]. Elasticity [CLPO22]. Elastoplastic [SW20]. Electrical [EMJ^{+}22]. Electronic [MTL^{+}21, TBL^{+}22]. Element [Kim20, MCC23, SBH22]. Elicitation [MCK^{+}22]. Elliptic [MRB^{+}21]. elliptical [BP21]. Embedding [BSK21, CM20, GMH22, PHP22, SKB^{+}22, XTL^{+}21]. Embeddings [EBB23, MH20]. Embroidery [ZPH^{+}23]. Emergency [LLC^{+}20, STX^{+}22]. Empathy [LPW20a]. Empty [AMT20]. EMU [MFJ^{+}21]. Enabling [SHVR21]. Encoding [LL20, RGJW20, XHZ^{+}20, DKVV23]. End [QLWQ21]. End-to-end [QLWQ21]. Energy [SWJG20]. Engine [WBM21]. Engineering [CMMK20, NZCS21]. Engines [EMJ^{+}22]. Enhance [EMJ^{+}22]. Enhanced [GWM22, NBPM20, TB21]. Enhancement [HWDL21, TBSB20, WYZ^{+}22, ZWCM21]. Enhancing [CMJ^{+}20]. Enrichment [CO22]. Ensembles [EHL21, HL22, TBL^{+}22]. Entries [MRSF20]. Envelope [KCRH20]. Environment [JPK20, WM22]. EnviRonments [RZZ^{+}22, ARC22, AKA^{+}20, PMCEA^{+}22, TAA^{+}21, PMG^{+}20]. Equal [RLHS20a]. Equations [LCG^{+}20]. equilibrium [JL21]. Equivalent [SW22]. Equivarianceg [PZL^{+}22]. Erosion [GGMP21]. Erratum [Ano21a, Ano22b]. Error [CRZ^{+}22]. Errors [GWB^{+}21, HCS20]. Estimating [BKL21]. Estimation [CXW^{+}21, EGH21, GHYZ20, GEE20, LHJ21, MHLH22, PZL^{+}22, ROCP20, RSLSRÁ21, SMM22, SRC^{+}20, ZK23, ZSL^{+}22, ZK21a, ZXJ^{+}20, HWD22, ZA21]. Estimator [GGN20]. estimators [LCBM21]. Ethical [EBD21]. Euclidean [NCB^{+}21]. Eulerian [GHMR^{+}20, JSG^{+}22, KMA20]. Evacuation [LLC^{+}20]. Evaluating [BZP^{+}20, LPW20a, ML22, XXeL^{+}20, Ano22b, EBM22]. Evaluation [BHJS23, KQL^{+}21, KCW^{+}21, LB20, LLLB22, SLJ^{+}20, VGdS^{+}20, ZWLY22]. Evaluations [CBL^{+}20, SSV^{+}22, SEAG^{+}21]. Event [AMA22b, GEE20, MLD^{+}21, XSZX22, HWD22]. Event-based [AMA22b]. Events [MSY^{+}21]. Evocube [DPM^{+}22]. Evoke [MAS22]. Evolution [MMP20, QLL^{+}22]. Evolutionary [CMKK21, ZWG^{+}22]. Evolving [KP23]. Exact [WFJ^{+}22]. Example [GFH^{+}23, GCLB21, HTT^{+}20, KJB^{+}22, XHZ^{+}20]. Example-Based [GCLB21, XHZ^{+}20, GFH^{+}23, HTT^{+}20]. Examples [HHG^{+}22]. Execution [SZM^{+}21]. Existing [WXL^{+}21]. Experience [GK22]. Experiences [ZLL^{+}22b]. Experiments [BDK^{+}22]. Expert [ZWVL22]. eXplainable [LBB^{+}23]. Explanations [SKB^{+}22]. Explicit [CYD^{+}22, CMG21, JSG^{+}22, BS22, SSS20b]. Exploration [AB20, AEC^{+}22, ATL^{+}23, BBBD20, EMJ^{+}22, Fre21, Fre22, JGvK21, JH22, KCW^{+}21, LLLB22, LCI^{+}21, LSE21, MLD^{+}21, MWLH^{+}20, NZL^{+}21, QLL^{+}22, SAK^{+}21, TBL^{+}22, TZP^{+}22, RMH^{+}22]. Exploring [CLC^{+}22, EEE^{+}20, MYT^{+}21, MAS22, NQHH20, RCPW21, XTL^{+}21, XSZX22, ZWVL22, BAMG21]. Exponent [SJS20]. exposure [ÇBSM23]. Expression [BRMO22, CZGM20, SOL^{+}22, WBZB20, ZYP^{+}21]. Expressions [CZGM20]. Extended [MRCK21, MMC^{+}20, RMSQ20]. External [NHSS21]. Extracting [MPSM21]. Extraction [CW22, GLLL20, GZG^{+}22, HS21, HLJ^{+}20, LL20, QZL^{+}20, TU22, YTK23, GCZ^{+}20, HS20]. Extractor [LXY^{+}22]. Eye [GK22, KADM22]. Eye-Tracking-Based [GK22]. Eyeball [WBB^{+}20]. Eyebrow [WZX^{+}22].

Fabric [JSVB22, RPPMCG23]. Fabricable [LLL^{+}22, PLL^{+}20]. Fabrication [II22, LLZ^{+}20, NZCS21, NTM^{+}21, SCHC20]. Fabrication-Aware [NZCS21]. Face [ABPP23, DBA^{+}21, JJZ^{+}22, LHZ^{+}22, NHSW20, ONH^{+}22, YT22]. Faces [CSK^{+}23, GFZ^{+}20, GLLL20]. Facets [WM22]. Facial [ALTT22, BRMO22, BSS20, CZG^{+}22a, GCZ^{+}20, GLLL20, KJS^{+}21, LLDT20, WBZB20, ZM22, ZYP^{+}21]. Facilitate [KCW^{+}21, dQNSVE20]. Facilitating [LLT^{+}20b]. Factorization [ZSWL21]. Factorizations [HPO22]. Factors [YGW^{+}21]. Facts [SSX^{+}21]. FAKIR [FCD20]. FARM [MMRC20]. Fascia [RMSQ20]. Fashion [FRH^{+}21]. Fast [AZ21, CDK20, CLSK20, CDAS20, FBG^{+}20, HPO22, LCG^{+}20, MDVP22, SOW20, TK20, WH20, WFJ^{+}22, WHW22, ZJA21]. FE [ZCC22]. Feature [GLLL20, GZG^{+}22, HLJ^{+}20, LL20, LTH20, LSE21, MWZ^{+}20, ZGP^{+}20, ZOM^{+}22, ZHHK22, WCS^{+}22]. Feature-Adaptive [ZHHK22]. Feature-based [LSE21]. Features [TLC20, Ano21a]. feedback [RCDB23]. Fetch [Csé23]. Few [CZGM20, SDR^{+}22]. Fiber [BRP^{+}20]. Fiblets [SMRB22]. Fidelity [ZWG^{+}21]. Field [BGP^{+}22, HMGU20, HS20, HS21, LTH20, MDVP22, PSS21, TACS21, YMS^{+}21, ZSS22, ZHL^{+}20, CDAS20]. Field-Controlled [ZSS22]. Fields [CL22a, CvTZ^{+}22, DLLK21, HBK^{+}21, LJLB21, NSP^{+}21, TU22, WDY^{+}22, XTS^{+}22, GSDT22, NBH20]. filling [BCOM^{+}22]. Film [Ron21, IWH23]. Filtered [WDH20]. Filtering [XK21]. Filters [CO22, HMSS22, SYG20]. Financial [ATL^{+}23]. Fine [DDB20, HWJ22, ZGP^{+}20, ZJZ^{+}22, vBGF22]. Fine-Grained [ZJZ^{+}22, vBGF22]. Fine-Tuning [DDB20]. Finger [HPSK21]. Finite [Kim20, MCC23, SJS20]. Finite-Time [SJS20]. First [LB20]. fisheye [CXWZ22]. Fishing [SFB20]. Fit [SDR^{+}22]. Fitting [BP20, HPM^{+}20, HWC21]. Flash [YDPG21]. Flash-lit [YDPG21]. Flexible [TZP^{+}22]. flight [LJJ22]. Flip [LZ22]. Flip-book [LZ22]. Floor [MPSM21, WXL^{+}21]. Floorplan [BSU^{+}22]. Flow [BYH^{+}20, CLCO^{+}22, DX22, EMKL21, FSK^{+}22, SLB22, WKA^{+}20, WWRT21, ZSL^{+}20, CLCO^{+}22]. Flower [ZWW^{+}21, Ano22a]. Flows [AHKB20, CLH^{+}20]. Fluid [GHMR^{+}20, JSG^{+}22, LCG^{+}20, NHSS21, OP22, QMSL20, TKAS23, WKA^{+}20]. Fluids [GGMP21, JLDH20, JL21, LWB^{+}21, PT23, TACS21]. Fluidymation [PCS21]. Fluorescence [JHD20]. Fluorescent [HTFW23]. Focal [QFWH21]. Follower [KF21]. Fonts [LG22]. Foot [MHLH22, PYL20]. Footskate [MHLH22]. Force [MHT^{+}20, MHLH22, NHSS21, TACS21]. Forest [EMJ^{+}22]. Formal [KSHL20]. Formulation [HXA21, Kim20, QSBJ22]. Fortified [MMP20]. Forward [OL21]. Four [Cav21]. Fourier [MMO^{+}21, MRB^{+}21, SSC^{+}20]. Foveated [FFM^{+}21]. Fracture [CLR^{+}20, CMSK20, HXA21, MCC23, SXHA22, ZHL^{+}20]. Fractured [LBB22]. Fragment [KK21]. Frame [MWS^{+}20a, PSS21, ÇBSM23]. Framework [BSS20, CTJ^{+}20, DLC^{+}22, DPM^{+}22, FVMK21, GHYZ20, JZGF20, Özt20, SR20, SZM^{+}21, WBZB20]. Free [Bd22, CLH^{+}20, HPG^{+}23, DSA20, JLDH20, MCC23]. Freeform [WCSH21]. Frequency [GOG21, TACS21]. Frequency-aware [TACS21]. Friction [EMAK20, PT23]. Front [Ano20a, Ano20b, Ano20c, Ano20d, Ano20e, Ano20f, Ano20g, Ano20h, Ano21b, Ano21c, Ano21d, Ano21e, Ano21f, Ano21g, Ano21h, Ano21i, Ano22c, Ano22d, Ano22e, Ano22f, Ano22g, Ano22h, Ano22i, Ano22j, Ano23a, Ano23b, CMMK20]. Full [GDG^{+}23, JPK20, TKE21, YLHX22]. Full-Body [GDG^{+}23, JPK20, TKE21, YLHX22]. Fully [VSGC20]. Function [CRGFT21, CK21, SAK20, WS20]. Functional [AL21, DCMO22, MO23, MMRC20, PKO22, CL21, PFCR20]. Functionality [RKL^{+}21]. Functionality-Driven [RKL^{+}21]. Functions [DXG^{+}20, GAD^{+}20, KPKH20, LBB22, MBB22]. Furniture [CL21, VPGV20, WYA^{+}20]. Further [Csé23]. Fusion [MWM^{+}22, RLHS20a, FLH^{+}20a]. Future [HKM22]. Fuzzy [LWL^{+}20].

Gabor [HMSS22]. Game [AWB20]. Games [Ron21, WHPG22]. Gamut [TWF22]. GAN [GWM22, GMH22, ZLY^{+}20, ZYP^{+}21]. GANs [JLB22]. Gaps [PMC22a]. Garment [BKL21, ZWCM21]. Garments [CETC22, VZTH22, WHZ^{+}23]. Gaseous [QLWQ21]. Gauss [BVHSH21, KFDA21]. Gaussian [CDAS20, HJL^{+}20, LSD21, MBB22, XWHM20]. Gaussian-based [HJL^{+}20]. Gaze [WBB^{+}20]. Gears [XFS^{+}20]. General [BB23]. Generalizations [NQHH20]. Generate [YLHX22]. Generating [ARC22, CSAN23, HYC^{+}23, MCBR20, WXL^{+}21, ZXC^{+}20, GSDT22]. Generation [BHJS23, CCOCM21, DLLK21, DZN^{+}21, GWM22, GFH^{+}23, GAM^{+}21, IIHG^{+}20, JGvK21, LFH^{+}22, NKA^{+}23, SHVR21, SOW20, SSX^{+}21, TSK^{+}20, WS20, WHPG22, ZA22, ZJZ^{+}22, ZLWT22, ZPIS22]. Generative [BSS20, CRW^{+}20, CSSI21, DLC^{+}22, GWM22, MA23, QLWQ21, SWW^{+}22a, WDY^{+}22, WWZ^{+}22, YUIH21, ZLY^{+}20, ZA21]. Genetic [DPM^{+}22]. Genus [BTCL22]. Geo [MZAD^{+}20, WBL20]. Geo-Semantic [WBL20]. Geo-Text [MZAD^{+}20]. Geodesic [SHR20, TBK21]. Geographic [ADP20, LCB21]. Geologically [PGP^{+}21]. Geometric [BHM21, GE21, JZZ^{+}20]. Geometrical [ZJZ^{+}22]. Geometries [CYD^{+}22]. Geometry [ONH^{+}22, OKN21, SBB^{+}22, DKVV23]. Geospatial [SYPB21, SFB20]. Gesture [AHKB20, GFH^{+}23, NKA^{+}23]. gestures [KMTM22]. Get [FTA22]. Get-Up [FTA22]. GI [OLK^{+}21]. Gist [KQL^{+}21]. Gist-based [KQL^{+}21]. Glass [LMF^{+}20, ZSY^{+}22]. GlassNet [ZSY^{+}22]. Glints [CSDD20, WDH20]. Global [CDAS20, DMT21, DBP20, MMK^{+}22, PMC^{+}22b, RLHS20b, WBZB20]. Global-Local [WBZB20]. Globally [CL22b, OKN21, WWW20]. Glossy [HJL^{+}20]. go [BvOdK^{+}22]. Golden [HYC^{+}23]. Good [McN21]. GPGPU [vBGF22]. GPU [ADP20, BAR^{+}21, JG21, MWS^{+}20b, SH21, SZM^{+}21, ZHL^{+}20]. GPU-based [ADP20]. GPUs [MRSF20]. Grade [PYAP22]. Gradient [GPM^{+}22, XSW^{+}20, ZHHK22, MMBP22]. Gradient-Domain [XSW^{+}20]. Grained [ZJZ^{+}22, vBGF22]. Grammar [CLL^{+}20, GMOB22]. Grammar-Based [CLL^{+}20, GMOB22]. Graph [AMA22b, BDK^{+}22, BSU^{+}22, CJMM22, FLH^{+}20a, FLH^{+}20b, HW22, LMBM20, LHJ^{+}22, MCC23, SAK20, SLMV21, VSGC20, WBL20, WMS^{+}23, ZJZ^{+}22, DPB^{+}22a]. Graph-Based [SAK20, WBL20, FLH^{+}20b, MCC23]. Graphical [JM22, QMSL20]. Graphics [BD23, Cav21, CCOCM21, KJB^{+}22, KBST22, LSY21, RGJ^{+}23, TLLS20].

Graphs [BBS20, LSK21, MSY^{+}21]. Grasping [HPSK21]. Greedy [ZCH^{+}23, ZYCF20]. Grid [BHJS23, BBS20, Fre22, SBBB20]. Grids [BDF^{+}22]. Ground [MHLH22, WWRT21]. Grounded [SLJ^{+}20]. Growth [GKK^{+}20]. GTMapLens [MZAD^{+}20]. guaranteed [HT20]. guaranteed-quality [HT20]. Guarding [KMC22]. Guidance [CSSI21, HWDL21, NHSS21, PMCEA^{+}22, SSKEA21, WM22]. Guide [KP23]. Guided [BTW^{+}22, BNR20, DDB20, GCDT22, JGvK21, SRK21, VCT21, VPGV20, WWF^{+}22, CL21]. Guider [KF21]. Guider-Follower [KF21]. Guiding [DGJ^{+}20, DPÖM22, SHJD22]. Gyre [WWRT21].

Hair [HHH22, BP21]. Hairy [SWB^{+}20]. Half [SCCZ21]. Half-body [SCCZ21]. Halfedge [DV21]. Hamiltonian [PFCR20]. Hand [KC22, SMM22, WRZ20, ZCH^{+}23]. Handicraft [II22]. Handle [CPK23, WMS^{+}23]. Handling [PT23]. Handwriting [TL21]. Hardness [CPK23]. HardVis [CPK23]. Hardware [LSY21, TST^{+}22]. Hardware-Agnostic [TST^{+}22]. Harmonic [DR22]. Harmonics [MDVP22, MMBP22]. Hausdorff [ZSL^{+}22]. HDR [AFTH21, YJS^{+}22, YLL^{+}21, ZA21]. HDRNet [GZLL23]. Headsets [FFM^{+}21]. Health [PL20]. Heat [KMP^{+}21]. Heritage [BBBD20]. Heterogeneity [Ano22b, ML22]. Heterogeneous [EEE^{+}20]. Heuristic [LCI^{+}21]. Hex [BRK^{+}22]. Hexa [BTL22]. Hexahedral [MPZS20]. Hidden [KUS^{+}21]. Hierarchical [CJMM22, GWE^{+}20, GMH22, LJSL21, MLD^{+}21, SMOW20, TBL^{+}22, ZLL20, ZHHK22]. Hierarchies [MOB^{+}21, VEPG23]. Hierarchy [JG21]. Hierarchy-Traversal [JG21]. High [BTB^{+}22, CLPO22, FIDF20, GZLL23, GZL^{+}20b, JZN^{+}22, LHJ^{+}22, LSY21, MRR^{+}22, ME23, NHSW20, SMM22, WM22, YLC^{+}22, ZXC^{+}20, ZMV^{+}21, ZWG^{+}21, ZXS^{+}21, ÇBSM23, IWH23]. High-Dimensional [FIDF20, GZLL23]. High-Fidelity [ZWG^{+}21]. High-Level [GZL^{+}20b, SMM22, ZXS^{+}21]. High-Order [CLPO22, LSY21, YLC^{+}22]. High-Precision [WM22]. High-Quality [JZN^{+}22, ZXC^{+}20, ZMV^{+}21]. High-Resolution [BTB^{+}22, MRR^{+}22, NHSW20, LHJ^{+}22, ME23, IWH23]. Higher [Cav21, LLJ^{+}20]. Higher-Order [LLJ^{+}20]. Highlights [WWZ^{+}22]. Hinted [ML20]. Histogram [MRCK21]. History [ELW^{+}22]. Hollowed [TCL20]. Homogeneous [Bd22]. Homology [BSCK21, FIDF20]. Homology-Preserving [FIDF20]. Honey [TACS21]. Horizon [PHP22]. Hornero [DPR^{+}21]. Hot [GMH22]. Hotspot [dQNSVE20]. Human [CSAN23, GDG^{+}23, GLLL20, HSvO^{+}22, KJS^{+}21, LHG^{+}23, MMRC20, MHL^{+}22, PYL20, QZL^{+}20, SYHL21, SEAG^{+}21, TKE21, WQL^{+}20, WNT^{+}20, ZLL20, ZWG^{+}21]. Human-Centered [SEAG^{+}21]. Humans [CLCO^{+}22, JPK20, KAC^{+}21, SGOC20, TKE21]. Hybrid [BMPP21, BDL^{+}20, GHMR^{+}20, HLJ^{+}20, SGBI22, SCHC20, ZK21a]. Hyperbolic [HS20, WWRT21, IA21]. Hyperelastic [KUKH23]. Hypergraphs [BDF^{+}22, DPB^{+}22a]. HyperNP [AEC^{+}22]. Hyperparameter [CMKK21]. Hyperparameters [AEC^{+}22].

Hyperspectral [LTDG20].

Iconized [CLL^{+}20]. Identification [BZP^{+}20, GWB^{+}21, SFB20]. Identifying [MDBH22]. Identity [CZG^{+}22a]. If [BRK^{+}22]. II [SWB^{+}20]. Illumination [BJG20, CDAS20, DMT21, EGH21, MMK^{+}22, VPGV20, YTK23]. Illumination-Aware [DMT21, YTK23]. Illumination-Driven [BJG20]. Illumination-Guided [VPGV20]. Image [AAC20, BNR20, BMBW21, BVHSH21, BBA21, CSAN23, CCOCM21, CXWZ22, CSSI21, DLC^{+}22, EK20, End22, GWM22, HWW^{+}22, JZN^{+}22, KJB^{+}22, KF21, MRR^{+}22, MA23, PMC22a, QLZ^{+}20, RLHS20a, RLHS20b, ROG20, SXZS21, SMM22, WWF^{+}22, WDY^{+}22, WYZ^{+}22, WWZ^{+}22, XCC22, XL20, XSW^{+}20, YTK23, YLL^{+}21, ZQW^{+}20, ZA22, ZYP^{+}21, ZSY^{+}22, ZK21a, ZWY23, ZFWW20, GRP22, ZK21b, WYA^{+}20]. Image-Based [AAC20, DLC^{+}22]. Image-Driven [WYA^{+}20]. Image-to-Image [XL20]. Images [BSB^{+}20, CM20, FFJ22, GOG21, HYC^{+}23, IIHG^{+}20, LFS21, RPPMCG23, WRZ20, ZXC^{+}20, NYB^{+}20, WS20]. Imaging [AFTH21, GSWS21, KWK20, NBPM20, SHV21]. IMAT [LBKP21]. Img2Logo [HYC^{+}23]. Immersive [KFS^{+}22, SAK^{+}21]. IMoS [GDG^{+}23]. Impairments [WJK22]. Imperfect [HSG^{+}22]. Implicit [CYD^{+}22, CDK20, LHZ^{+}22, LCP^{+}23, MRB^{+}21, ZAB^{+}21, ZLWT22, YZ22]. Importance [GGS21, HTFW23, Pet21, SSC^{+}20, SRRW21, TB21, CGG^{+}22, LSD23]. Importance-Driven [TB21]. improve [IWH23]. Improved [BHJS23, ELW^{+}22, GZL^{+}20a]. Improvement [MvdEV22, NCB^{+}21]. Improves [SDR^{+}22]. In-Hand [ZCH^{+}23]. In-the-wild [SL23]. Inbetweening [JSL22]. Incident [KQL^{+}21]. Incorporation [SGAG22]. Incremental [ATL^{+}23]. Independent [CZG^{+}22b, WLG22]. Indicators [SBMS23]. Individual [LCO20]. Indoor [FYFL20, MPSM21, PMG^{+}20]. Industrial [WM22]. Inextensible [ZLW^{+}22]. Inference [BSCK21, GHYZ20, RSGP21]. Inferring [LG22]. Influence [LPW20a]. Infographics [TZP^{+}22, WGH^{+}21]. Infomages [CM20]. InfoMotion [WGH^{+}21]. Information [Ano20i, Ano20j, Ano21k, Ano21j, Ano22l, Ano22k, Ano23c, QLZ^{+}20]. Informative [KMP^{+}21]. Informed [TKAS23]. Infused [LSK23]. Initiative [PMCEA^{+}22, SSKEA21]. Injective [OKN21]. Inner [DLX^{+}22]. Inpainting [ZQW^{+}20]. Inputs [SYHL21, TAA^{+}21]. Inquiry [BJSM22]. Insights [LXYM20]. Instance [CPK23, LPW^{+}20b]. Instance-level [LPW^{+}20b]. InstanceFusion [LPW^{+}20b]. Instant [MHG^{+}22]. Integer [SBBB20]. Integer-Grid [SBBB20]. Integral [AZ21, Bd22, LHJ21]. Integrals [Sub21]. Integrated [MWLH^{+}20]. Integrating [MWM^{+}22]. Integration [GE21, LLJ^{+}20, MLD^{+}21, ZSWL21]. Intensive [BSP^{+}22]. Intent [GDG^{+}23]. Intent-Driven [GDG^{+}23]. Intentions [AKA^{+}20]. Interacting [SYPB21]. Interaction [CTJ^{+}20, CvTZ^{+}22, GMOB22, GMH22, HPKI22, SFMS^{+}21, ZXS^{+}21]. Interactions [GDG^{+}23, JPK20, XOW^{+}20]. Interactive [Ano22b, AEC^{+}22, CLL^{+}20, CSSI21, DXG^{+}20, EMJ^{+}22, FSS^{+}21, FYFL20, GCL22, GBYH20, GFZ^{+}20, HTI22, KFDA21, LPH^{+}20, LLLB22, LGW^{+}22, MZAD^{+}20, MPE23, ML22, MvdEV22, MDBH22, RCDB23, SRM^{+}21, SLB^{+}21, SMS^{+}20, TÇE^{+}21, VGL^{+}20, WL22, WYA^{+}20, XSZX22, YBRP21, YGW^{+}21, BAMG21, RGR20]. Interactively [CBE20, JLB22]. Interface [CL21]. Interfaces [BYK^{+}21, ZWVL22]. INterior [RZZ^{+}22]. Interplay [CWW^{+}22, LCB21]. Interpolants [CLPO22]. Interpolated [LRTC20]. Interpolating [BSH22]. Interpolation [Csé23, DR22, LSY21, OL21, ÇBSM23]. Interpolatory [HYW^{+}22]. Interpretation [KSHL20]. Interpreting [BYH^{+}20]. Interreflections [HJL^{+}20]. Intrinsic [SRM^{+}21]. Intuitive [BSS20, CvTZ^{+}22]. Invariant [AZ21]. Inverse [AWKK21, CYD^{+}22, KJB^{+}22, LTDG20, PMC^{+}22b, XK21]. Invert [Hei20]. Invertible [LNRH20]. Investigating [CWW^{+}22]. Investigation [CCFM23]. Investment [YGW^{+}21]. iQUANT [YGW^{+}21]. Isochrone [VRBT21]. Isolated [ZCC22]. Isometric [PKO22]. Isoperimetric [ZDS20]. Isosurface [KRF22, NMN20, Wan20]. isosurfaces [HT20]. Isovalue [Wan20]. Issue [Ano20i, Ano20j, Ano21k, Ano21j, Ano22l, Ano22k, Ano23c]. Iterative [LBKP21].

Jet [WBM21]. Joint [LWL^{+}20, MTB^{+}22, SMM22]. Joints [LHX^{+}22]. JOKR [MTB^{+}22]. Journalism [MCK^{+}22].

KD [LCI^{+}21]. KD-trees [LCI^{+}21]. Kernel [FWHB21, KMP^{+}21, ZXJ^{+}20]. Kernels [MWS^{+}20b, vBGF22]. Key [WCM22]. Keyframe [HWJ22]. Keyframes [PWBD21]. Keypoint [MTB^{+}22]. Kinesthetic [VZTH22]. Knot [LZ22]. Knowledge [AACF22, ATL^{+}23, CLC^{+}22, NBPM20]. Knowledge-Assisted [NBPM20]. known [DKVV23].

Label [SHVR21, ZSY^{+}22]. labeling [NRS22]. Labelling [DPM^{+}22]. Labels [BNR20]. Lagrangian [GHMR^{+}20, HXA21, SXHA22]. Lagrangian/Eulerian [GHMR^{+}20]. Lambertian [d’E21]. Landmark [GCZ^{+}20, GH21, PKO22, ZM22]. Landmark-Adapted [PKO22]. Landscape [TÇE^{+}21]. Language [GZL^{+}20b, SKB^{+}22]. Laplace [AHKSH20]. Laplacian [BHKB20, SC20]. Laplacians [BB23]. Lapped [GGMP21]. Large [AKA^{+}20, CW22, DDB20, FBG^{+}20, Fre22, JM22, RMH^{+}22, SFMS^{+}21, SW22, YSLF20, ZCC22, ZZF^{+}22]. Large-Scale [DDB20, FBG^{+}20, SFMS^{+}21, ZCC22, ZZF^{+}22, RMH^{+}22]. Latent [ABPP23, End22, GWE^{+}20, JGvK21, WKA^{+}20, XL20]. Lattice [WWW20]. Lattices [LHJ21]. Layer [DX22, EK20, HJZ^{+}21, LXY^{+}22, MH20]. Layer-aware [DX22]. Layer-wise [EK20, HJZ^{+}21]. Layered [GGN20, XWHM20, DPB^{+}22a]. Layerizing [WTO^{+}20]. Layers [DVP22, ECC^{+}20]. Layerwise [SKB^{+}22]. Laying [AWB20]. Layout [BTW^{+}22, BSK21, End22, GWM22, KOYSS21, LG22, LWL^{+}20, MRSF20, VPGV20, WNT^{+}20, DPB^{+}22a]. Layout-to-Image [GWM22]. Layouts [BHJS23, BDF^{+}22, BDK^{+}22, BSU^{+}22, Fre22, LCK21a, LCK21b, ZM22]. LBVH [JG21]. LBVH-Construction [JG21]. Learn [LFR23, LGS^{+}22]. Learned [KJB^{+}22, LBB22]. Learning [AFTH21, BTCL22, CMJ^{+}20, CTJ^{+}20, CXW^{+}21, CL22b, CJMM22, DLLK21, DZN^{+}21, FLH^{+}20b, GK22, HW22, HACOG22, HSvO^{+}22, HTI22, HPKI22, JVW20, JZN^{+}22, KJS^{+}21, KAK^{+}22, LBB^{+}23, LPH^{+}20, LHW21, LFR23, MYT^{+}21, MMC^{+}22, MHL^{+}22, MHLH22, ONH^{+}22, PZL^{+}22, QZL^{+}20, RLG^{+}20, RCPW21, SHVR21, SDR^{+}22, SYG20, SCZ21, SYHL21, SSKEA21, SEAG^{+}21, TL21, WDK^{+}20, WCS^{+}22, WYZ^{+}22, WS20, XLKvdP20, YBF22, YLHX22, ZCH^{+}23, ZWM^{+}20, CRW^{+}20, LAK20]. Learning-Based [CTJ^{+}20, KJS^{+}21, QZL^{+}20]. Least [FBG^{+}20, ZJA21]. Least-Squares [ZJA21]. Lens [MZAD^{+}20]. lenses [BAMG21]. Level [CTJ^{+}20, Fre22, GZL^{+}20b, PRCO20, QLL^{+}22, TBL^{+}22, LPW^{+}20b, SMM22, ZXS^{+}21]. Level-of-Detail [Fre22]. Leveraging [ELW^{+}22, MSY^{+}21]. Levitating [KUS^{+}21]. Ligand [SFMS^{+}21]. Light [CDAS20, DLS21, GWH20, GEE20, HMGU20, LFD^{+}20, MVNT21, NID20, PLL^{+}20, SS21]. Light-field [CDAS20]. Light-Weight [DLS21]. Lightcuts [HJL^{+}20]. Lighting [DBP20, HWDL21, SCCZ21, UPE22, WSN^{+}20, YT22]. Lightness [ZLY^{+}20]. Lights [CRZ^{+}22, MDVP22, Pet21, WH20, MMBP22]. Lightweight [SLCL22, WBB^{+}20]. Like [EMML22, HHS20, LL20, RPPMCG23, TL21]. Likert [SSV^{+}22]. Limited [FKL^{+}21, HACOG22, LWH^{+}20, MCBR20, QLWQ21]. Line [AMT20, ADP20, CMG21, LWZ^{+}20, NODI21, TB21, WCR^{+}23, ZWW^{+}21, ZLL^{+}22a, VLLA22, TB21]. Lineage [HTI22]. LineageD [HTI22]. Linear [BHJS23, Mit22, MWS^{+}20b, Pet21, QFWH21, YY20, EBB23]. Linearly [DGJ^{+}20]. Lines [BvOdK^{+}22]. Link [DCMO22, PSK^{+}20]. Lipschitz [GGPP20]. Liquid [GZL^{+}20a, KMA20, NA22, SPDM20, SW22]. Lists [BGT20]. lit [YDPG21]. Live [USK21]. LMFingerprints [SKB^{+}22]. LoBSTr [YKL21]. Local [BSH22, CM21b, GGPP20, GZG^{+}22, HPO22, HS21, KC22, LBK20, MN20, MVNT21, MRCK21, SRC^{+}20, WBZB20, YZ22]. LOCALIS [ADP20]. Localized [KMP^{+}21, PMC^{+}22b, SSS^{+}20a, WTB^{+}21]. Locally [ADP20, CL22b, NBH20, ADP22]. Locally-adaptive [ADP20]. Location [LLT^{+}20b, OM20]. Location-Based [OM20]. Locomotion [GK22]. Locus [MLP21]. Logos [HYC^{+}23]. Logs [GMOB22, SLJ^{+}20]. Longitudinal [ANI^{+}20, JM22]. Loop [LNRH20, WMS^{+}23]. LOOPS [ADP22]. Loss [KPKH20, SWW^{+}22b, WCS^{+}22]. Low [Sub21]. Low-dimensional [Sub21]. Lower [YKL21]. Lower-body [YKL21]. Luminance [MDVP22, YLL^{+}21]. Lyapunov [SJS20].

Machine [CMJ^{+}20, GK22, HTI22, SEAG^{+}21, WNT^{+}20]. Machines [NTM^{+}21]. Made [BHKB20, DZN^{+}21, SBB^{+}22]. Magnetic [wKH20, NBPM20]. Makeup [YTK23, YT22]. Making [ARC22, EST20, JHC^{+}20, ZWVL22]. Man [DZN^{+}21]. Man-Made [DZN^{+}21]. Manhattan [HWC21]. Manhattan-world [HWC21]. Manifold [JC22, LHL^{+}21, MMMR22]. Manipulating [HGC^{+}20]. Manipulation [RLR^{+}20, ZFWW20]. Manipulations [ZCH^{+}23]. Manufacturing [SMS^{+}20]. Many [AKA^{+}20, WH20, BRP^{+}20]. Map [AL21, BTW^{+}22, BSCK21, EMML22, HHS20, HRWO20, LXY^{+}22, MZAD^{+}20, MO23, SLB22, TZX^{+}22, WNT^{+}20]. Map-Like [EMML22, HHS20]. mapped [GSDT22, TZX^{+}22]. Mapping [AWKK21, CM21a, GEE20, MAA20, WJK22, ZXJ^{+}20, WCS^{+}22]. Maps [BBS20, DCMO22, DPM^{+}22, EMML22, NA22, PKO22, PSK^{+}20, SPK23, VRBT21, PFCR20]. Mark [GVF^{+}21]. Mark-based [GVF^{+}21]. Markov [GWH20]. masking [WCS^{+}22]. Masks [AFTH21]. Massive [SMRB22, SOW20, RZZ^{+}22]. Match [GMH22]. Matching [BRMO22, CBC20, MMCJ22, NODI21, PKO22, PYAP22, RMWO21, WM22]. Matchstick [EMAK20]. MaterIA [MRR^{+}22]. Material [CLR^{+}20, CRGFT21, DLC^{+}22, DDB20, GHYZ20, HHG^{+}22, MRR^{+}22, SXHA22, SL23, UPE22, WDK^{+}20, YDPG21, ZFWW20]. Materials [FJM^{+}20, GGN20, SW20, TCT22, XWHM20, d’E21]. Mathematical [BYH^{+}20]. Matrices [MRSF20, RLR^{+}20]. Matrix [WH20]. Matter [Ano20a, Ano20b, Ano20c, Ano20d, Ano20e, Ano20f, Ano20g, Ano20h, Ano21b, Ano21c, Ano21d, Ano21e, Ano21f, Ano21g, Ano21h, Ano21i, Ano22c, Ano22d, Ano22e, Ano22f, Ano22g, Ano22h, Ano22i, Ano22j, Ano23a, Ano23b]. Matting [DLS21, QLZ^{+}20, WZX^{+}22]. Maximal [Mit22]. MCFS [GLLL20]. Me [BRK^{+}22]. Mean [LWH^{+}20]. Mean-Variance [LWH^{+}20]. Meaningfulness [BYH^{+}20]. Measured [HGC^{+}20]. Measurement [GGD^{+}20]. Measures [Wan20, LRTC20]. Mechanical [LHX^{+}22]. Mechanics [HBK^{+}21, RPPMCG23, ROCP20]. Mechanisms [GAM^{+}21]. Media [DVP22, FSK^{+}22, PCS21]. Medial [LBKP21, ZDS20]. Medical [GSWS21]. Membranes [HXA21]. Memento [SSS^{+}20a]. Memory [JVW20, YSLF20, vBGF22]. Memory-Efficient [YSLF20]. MendNet [LBB22]. Merge [WLG22]. Merging [KK21]. Mesh [BJG20, CMM^{+}20, DLLK21, HPG^{+}23, KFDA21, LLT^{+}20a, LTY^{+}22, LHJ^{+}22, LBK20, LCK21a, MPZS20, MWS^{+}20b, SBMS23, WMU^{+}20, ZSL^{+}20, DKVV23]. Mesh-Based [LTY^{+}22]. Meshes [AHKSH20, BTL22, BB23, MO23, PHP22, SC20, UKPW21, ZSL^{+}22, ZHHK22, HT20]. MeshFormer [LHJ^{+}22]. Meshing [XABR21, YLC^{+}22]. Meshlet [UKPW21]. Meshlets [NMSS22]. Meso [VGL^{+}20]. Meso-scale [VGL^{+}20]. Meta [TL21]. Metal [MWS^{+}20a]. metamaterials [BHM21]. Method [CLR^{+}20, FLH^{+}20b, GHMR^{+}20, GZL^{+}20a, GCZ^{+}20, HZSQ20, JSG^{+}22, MCC23, MWM^{+}22, MMRC20, SW20, SW22, SXHA22, SBH22, WBL20, ZHL^{+}20, vdRE21]. Methods [BGA^{+}22, KBST22, KAK^{+}22, LYA^{+}21, MEM^{+}20, DPB^{+}22a]. Metric [BP20, TL21]. Metric-based [TL21]. Metrics [DFB^{+}21, LBFS21]. Metro [BBS20, BTW^{+}22]. Micro [QMSL20, GSDT22]. Micro-partitioning [QMSL20]. micro-patterns [GSDT22]. Microfacet [AWKK21, HHH22].

Microfacet-based [HHH22]. Microfacets [Bd22, WDH20]. Microscopic [vTP21]. Microstructure [CLPO22, CRGFT21, FJF20, TZX^{+}22, LFD^{+}20]. Microsurface [AKDW22]. Milling [NTM^{+}21]. MINERVAS [RZZ^{+}22]. Minimization [TK20]. Minimizing [LBK20]. Minimum [SYKE20]. Mining [CLC^{+}22]. MIP [TZX^{+}22]. MIP-Mapped [TZX^{+}22]. Misinformative [LGS^{+}22]. Misinformed [LGS^{+}22]. Missing [ANI^{+}20, BvOdK^{+}22, JM22]. Mitral [WGT^{+}22]. Mix [GMH22]. Mixed [BTW^{+}22, CTJ^{+}20, ML20, PMCEA^{+}22, SSKEA21]. Mixed-Initiative [PMCEA^{+}22, SSKEA21]. Mixing [CCR^{+}20]. Mixture [DPÖM22, JLDH20, JL21, SWW^{+}22a]. Mobile [EEE^{+}20, LLLB22]. Mobility [SCZ21]. MoCo [CLCO^{+}22]. MoCo-Flow [CLCO^{+}22]. Model [AWKK21, BPD20, DWH^{+}22, EMAK20, GWE^{+}20, GvdVS23, HHH22, JLDH20, JL21, JH22, KC22, KSHL20, LFH^{+}22, LLDT20, MvdEV22, MN20, NHSS21, PRCO20, QLWQ21, SKB^{+}22, SSKEA21, TBSB20, VBRKW22, BP21]. Model-based [TBSB20]. Model-reduced [MN20]. Modeling [CL21, FSS^{+}21, GKG^{+}22, GFZ^{+}20, KOYSS21, KFDA21, MRB^{+}21, ROCP20, SGOC20, SMS^{+}20, WRZ20, WYA^{+}20]. Modelling [CLL^{+}20, CSSI21, CRGFT21, PYL20, PMC^{+}22b, WSN^{+}20, YJS^{+}22, ZWW^{+}21]. Models [CLR^{+}20, CZG^{+}22b, CBL^{+}20, CMJ^{+}20, CRW^{+}20, CW22, DPÖM22, EGH21, FJM^{+}20, HSvO^{+}22, HJZ^{+}21, JJZ^{+}22, MCT21, RGJ^{+}23, WDK^{+}20, XABR21, YSLF20, HTT^{+}20, PCDS20]. ModelWise [MvdEV22]. Modern [KMTH22]. Modified [WWRT21]. Modifying [CBE20, TÇE^{+}21]. MODNet [HXW^{+}22]. Modulated [ZPIS22]. Mold [MMMR22]. Molecules [MTL^{+}21]. Mollification [JHD20]. MoMaS [MMMR22]. Moment [TWF22]. Moment-based [TWF22]. Monitoring [SFB20]. Monochrome [QFWH21]. Monocular [BRMO22, CETC22, CLCO^{+}22, HZSQ20, NHSS21, TKE21, TKR^{+}23, WQL^{+}20]. Monolithic [PT23]. Monoscopic [CXWZ22]. Monotone [BSH22]. Monte [FWHB21, FFJ22, GWH20, GE21, LWY^{+}21, LZBD21, SSC^{+}20, ZMV^{+}21]. MoreGAN [SWW^{+}22a]. Morphable [JJZ^{+}22]. Morphologically [CHC^{+}23]. Morphologically-similar [CHC^{+}23]. Morphology [LLL^{+}21]. Most [BTL22]. Motion [ARC22, CZG^{+}22a, CL21, CLCO^{+}22, CHC^{+}23, ECC^{+}20, GWE^{+}20, GDG^{+}23, HPSK21, HPKI22, KPKH20, LHG^{+}23, MBA22, PYAP22, SLCL22, XK21, YYH20, YLHX22, ZLY^{+}21]. Motion-guided [CL21]. MotionGlyphs [CSB^{+}20]. Motions [FTA22]. Motor [MYT^{+}21]. Motorcycle [BGMC22]. Mould [AMA^{+}22a]. Movement [WHZ^{+}23]. Moving [Fre21, MRWE20, SS21, WLNM21]. MP [CL22a]. MP-NeRF [CL22a]. MPM [BS22, SSS20b]. MRI [GvdVS23]. Multi [CL22a, CL21, DFV21, FWBI22, GMH22, GLLL20, HXW^{+}22, JZN^{+}22, KK21, LMBM20, LSK21, LSK23, LY20, LLL^{+}22, LHG^{+}23, MVNT21, NODI21, PYL20, PRD21, QLZ^{+}20, QLL^{+}22, SLS^{+}21, WRZ20, XTL^{+}21, XL20, ZWG^{+}21, vdRE21]. Multi-Attribute [XL20]. Multi-Channel [GLLL20]. Multi-dimensional [XTL^{+}21]. Multi-Fragment [KK21]. Multi-functional [CL21]. Multi-Hot [GMH22].

Multi-Human [LHG^{+}23]. Multi-Level [QLL^{+}22]. Multi-offset [HXW^{+}22]. multi-pass [vdRE21]. Multi-Person [LY20, CL22a]. Multi-Phase [PRD21]. Multi-Resolution [LSK21]. Multi-Scale [JZN^{+}22, LLL^{+}22, MVNT21, HXW^{+}22, LMBM20, QLZ^{+}20]. Multi-Segment [PYL20]. Multi-Source [LSK23]. Multi-Stage [DFV21]. Multi-View [FWBI22, WRZ20, ZWG^{+}21, NODI21, SLS^{+}21]. Multidimensional [AEC^{+}22, VGdS^{+}20]. Multifragment [VVP20]. Multilinear [WBZB20]. Multimodal [LLLB22, SLL20]. Multiparametric [MWLH^{+}20]. Multiphase [JLDH20, JL21]. Multiple [Bd22, GGS21, LHW21, LWL^{+}20, NID20, WRZ20]. Multiple-Scattering [LHW21]. Multiply [DR22]. Multiply-connected [DR22]. Multirate [HYW^{+}22]. MultiResGNet [LSK21]. Multiscale [AWKK21, GKV^{+}22, MB20]. Multivariate [BSP^{+}22, EEE^{+}20, PPF23, XSZX22]. Multiverse [HLJ^{+}22]. Muscle [MFJ^{+}21, RMSQ20]. Musculature [RKL^{+}21]. Music [MFH^{+}22, MRKEA22]. Musical [KKM^{+}20]. my [BvOdK^{+}22, EBM22].

N [LTY^{+}22]. N-Cloth [LTY^{+}22]. Nameability [RSGP21]. Nanoscale [TCG^{+}22]. Narrations [CWW^{+}22]. Narrative [DFV21]. Narrow [OP22]. Narrow-Band [OP22]. Natural [PCS21]. Navigation [KF21, LFS21]. Near [CRZ^{+}22, WSN^{+}20]. Near-Lighting [WSN^{+}20]. Neighborhood [LCI^{+}21, TCG^{+}22]. Neighborhood-based [LCI^{+}21]. Neighbour [BGT20]. NeRF [BGP^{+}22, CL22a]. NeRF-Tex [BGP^{+}22]. Nested [SKRW22, ZCLJ22]. NET [Sub21, MWZ^{+}20, RLHS20b]. Nets [LGW^{+}22, LLZ^{+}20]. Network [CCG^{+}22, CMM^{+}20, CJMM22, FLH^{+}20a, FWHB21, GWM22, GZLL23, HW22, HCL^{+}22, HXW^{+}22, HL22, JZZ^{+}20, LWY^{+}21, MWZ^{+}20, MHT^{+}20, OL21, RLHS20a, RLHS20b, SWW^{+}22a, SAK^{+}21, Sub21, WTO^{+}20, WWF^{+}22, WWZ^{+}22, ZGP^{+}20, ZLY^{+}20, ZSY^{+}22, ZWM^{+}20]. Networks [CSB^{+}20, CSAN23, EEE^{+}20, LTY^{+}22, MA23, NSP^{+}21, PGP^{+}21, PPF23, SYPB21, SLMV21, SRC^{+}20, VSGC20, XL20, YLL^{+}21, GRGH20]. Neural [BGP^{+}22, CL22a, CCOCM21, CLCO^{+}22, CMM^{+}20, EGH21, GWB^{+}21, HMS^{+}20, HPG^{+}23, HL22, KPLD21, LHZ^{+}22, LWY^{+}21, LJLB21, MMH^{+}21, MH20, NHSW20, NSP^{+}21, RGJW20, RBRD22, RSB^{+}21, SLB22, Sub21, SLMV21, SRC^{+}20, SRRW21, TKAS23, TFT^{+}20, TTM^{+}22, VSGC20, WHW22, XTS^{+}22, ZWW^{+}21, ZSS22, ZSY^{+}22, ZPIS22]. Neural3Points [YLHX22]. Neurosymbolic [RGJ^{+}23]. Next [GEE20, HWD22]. NN [JG21]. Nodal [PRCO20]. Node [PSK^{+}20]. Node-Link [PSK^{+}20]. Noise [CRGFT21, vOSRR21, GSDT22, LSD21]. Noisy [BNR20]. Non [Ano21a, DYDZ22, EBB23, JL21, OKN21, PKO22, TLC20, TKR^{+}23, XFS^{+}20]. Non-Circular [XFS^{+}20]. Non-equilibrium [JL21]. Non-Injective [OKN21]. Non-Isometric [PKO22]. Non-linear [EBB23]. Non-Rigid [DYDZ22, TKR^{+}23]. Non-Uniform [TLC20, Ano21a]. Nonconvex [OPY^{+}20]. Nonlinear [FBG^{+}20, LSK21, ROCP20, SGOC20, SHR20].

Nonmanifold [SC20]. Noon [TÇE^{+}21]. Normal [LJ21, TZX^{+}22, WSN^{+}20, ZSL^{+}20]. Normal-Based [WSN^{+}20]. Normal-Driven [LJ21]. Normalising [AHKB20]. Normals [KFDA21]. NOVA [KAC^{+}21]. Novel [RLHS20a, Wan20, WXL^{+}21, ZHL^{+}20, ZWY23, BAMG21]. Novice [ZWVL22]. NSTO [ZPIS22]. nuclear [Ano23d]. Numerical [HPO22, SXHA22]. NURBS [SH21, XABR21].

Object [FAG^{+}20, GDG^{+}23, HZSQ20, HPSK21, HPKI22, PZL^{+}22, SMM22, WYZ^{+}22, ZGP^{+}20, ZWY23]. Objective [DXG^{+}20]. Objects [BTCL22, DZN^{+}21, wKH20, KKN^{+}21, KUS^{+}21, VRBT21, WDY^{+}22, WLNM21, TCL20]. Occluder [WHPG22]. Occlusion [HZSQ20, IIHG^{+}20, LJSL21, PHP22]. Occlusion-aware [HZSQ20]. Occlusions [BLC20]. Occupancy [LBB22]. Octilinear [BBS20]. Octree [SOW20]. Ocular [CK21]. Ocupado [OM20]. Offs [KMH21]. offset [HXW^{+}22, YZ22]. Omnidirectional [KMTH22, MA23]. Onboarding [DWH^{+}22]. Once [AKA^{+}20, HWD22]. Once-more [HWD22]. One [Csé23, FLH^{+}20b, SMOW20]. One-Shot [FLH^{+}20b]. Online [CLL^{+}22, CHC^{+}23, TL21, XXeL^{+}20]. Only [SBH22]. onto [PHP22]. Operations [SFB20]. operator [WCS^{+}22]. Operators [AHKSH20, PSS21]. Opportunities [BCD20, GKV^{+}22, KJRK21]. Optical [FJM^{+}20]. Optimal [BD23, GKK^{+}20, QZLC22, RSLSRÁ21, XCC22]. Optimised [WDH^{+}21]. Optimization [ABPP23, BMPP21, BSK21, CMKK21, CSSI21, CLSK20, FBG^{+}20, HWJ22, KUKH23, KOYSS21, KPLD21, LHX^{+}22, OKN21, RMWO21, RSP22, SKW21, SOL^{+}22, TACS21, VPGV20, XFS^{+}20, ZK23, ZWG^{+}22, ZZF^{+}22, ZHL^{+}20, ZPIS22, YPZ^{+}22]. Optimization-based [KUKH23]. Optimizations [SRK21]. Optimized [ADP22, WTB^{+}21]. Optimizing [Fre22, GCH^{+}21, JG21, SH21, FAG^{+}20]. Oracle [NSP^{+}21]. Orchard [EEE^{+}20]. Order [CLPO22, JSG^{+}22, LSY21, LLJ^{+}20, SKW21, TCT22, TB21, YLC^{+}22]. Orientation [CXW^{+}21, PYAP22, BCOM^{+}22]. Orientations [MRCK21]. Oriented [VEPG23, LCBM21]. Origin [TC21]. Origin-Destination [TC21]. Orthognathic [DPB^{+}22b]. Orthogonalized [MMO^{+}21]. Ostensibly [CDK20]. Ostensibly-Implicit [CDK20]. Our [LYA^{+}21]. Out-of-Core [SOW20, CW22]. OutCast [GRP22]. Outcome [DPB^{+}22b]. Outcome-Driven [DPB^{+}22b]. Outdoor [GRP22]. Outliers [RLG^{+}20]. Overcomplete [SCCZ21]. Overlap [CCG^{+}22, ZJZ^{+}22]. Overlapping [AB20]. Oversampling [CPK23]. Overview [GKV^{+}22].

Packing [CZGM20, HSS21, TST^{+}22]. Page [KOYSS21]. Pailou [CLL^{+}20]. Pain [AMA22b]. Painting [COKE20, PMC22a, YPZ^{+}22]. Paintings [CNS22]. Pair [BMBW21]. Pairing [FKL^{+}21]. Palette [TU22]. Panorama [CXWZ22, YLL^{+}21]. Papercrafts [SKRW22]. Paradigm [Ano20k, GFZ^{+}20]. Paradigms [TLLS20]. Paradoxic [LNRH20]. Parallel [Ano22b, BvOdK^{+}22, BZP^{+}20, DFB^{+}21, DV21, ML22, VEPG23, XL20]. Parameter [DLS21, EST20, GHYZ20, PBM^{+}22, RSP22, ZTZC20]. Parameterization [Hei20, LL20, SWW^{+}22b]. Parameterizations [YSLF20, ZYCF20]. Parameterized [RW21]. Parametric [SGOC20, TRPO21, VSGC20, PFCR20]. Parametrization [HPG^{+}23, Lev22, ZSL^{+}20]. Parametrization-Free [HPG^{+}23]. Pareto [CMMK20]. Parity [WFJ^{+}22]. ParSetgnostics [DFB^{+}21]. Part [ABPP23, DZN^{+}21, LAK20, NYB^{+}20, SCZ21, WQL^{+}20]. Part-Based [ABPP23]. Part-wise [NYB^{+}20]. Partial [CCG^{+}22, MMC^{+}22, QZLC22, RDG20]. Participating [DVP22]. Particle [HXA21, MSY^{+}21, SPDM20]. Particle-based [HXA21, SPDM20]. partitioning [QMSL20]. Partitions [VLLA22]. Parts [WSP21]. pass [CGG^{+}22, vdRE21]. Patch [GGMP21, ZYP^{+}21]. Patches [HXW^{+}22]. Path [Bd22, DGJ^{+}20, DPÖM22, FJF20, LWY^{+}21, LHJ21, NID20, OLK^{+}21, SHJD22, WDH^{+}21]. Path-based [LWY^{+}21]. Pathline [NZL^{+}21]. paths [NID20]. Patient [MRB^{+}21]. Patient-Specific [MRB^{+}21]. Pattern [AACF22, GAM^{+}21, HMSS22, LPH^{+}20, LSP^{+}21]. Patterns [AACF22, BKL21, CLR^{+}20, GZG^{+}22, KMH21, Özt20, RSP22, ZPH^{+}23, GSDT22]. PAVED [CMMK20]. PD [YY20]. PDViz [HNWC23]. Peax [LPH^{+}20]. Pedestrian [KF21]. People [WJK22]. Per-View [KPLD21]. Perceptual [DLC^{+}22, LBFS21, PMC22a, SL23, SWW^{+}22b]. Perceptual-Loss-on-Screen [SWW^{+}22b]. Performance [BRMO22, PMB^{+}23, ZM22]. PERGAMO [CETC22]. Perlin [CRGFT21]. Persistence [LMBM20]. Person [LY20, CL22a]. Personality [MWM^{+}22]. Personalized [CETC22, WRZ20, WHZ^{+}23]. Perspectives [WNT^{+}20]. Perturbations [MCBR20]. Phase [BGG20, FSK^{+}22, PRD21, SR20, ZHL^{+}20]. Phase-Field [ZHL^{+}20]. Phenomena [FMKK21, JM22, QLWQ21]. Phones [EEE^{+}20]. Photo [MCT21, SRM^{+}21]. Photo-Reconstructed [MCT21]. Photograph [DBP20, LWZ^{+}20]. Photographic [LADH22]. Photographs [TÇE^{+}21]. Photometric [CRZ^{+}22, HZ22, WSN^{+}20]. Photon [CM21a, ZXJ^{+}20, LJJ22]. Photorealistic [ZFWW20]. Physical [BYH^{+}20, DSMA^{+}21]. Physicalization [DSMA^{+}21, RGW20]. Physically [CSDD20, MMCJ22, RDG20, WCSH21, YLHX22]. Physically-based [WCSH21]. Physics [CYD^{+}22, CJMM22, FTA22, NZL^{+}21, SXHA22, TKAS23, YY20]. Physics-Based [CYD^{+}22, FTA22, NZL^{+}21, YY20]. Physics-Informed [TKAS23]. Physiology [GKV^{+}22]. Piecewise [HYW^{+}22, WCR^{+}23]. Piecewise-Smooth [WCR^{+}23]. Pipeline [CCOCM21]. Pixel [II22, JZGF20, MPE23, ZQW^{+}20]. Pixel-Based [JZGF20]. Pixel-wise [ZQW^{+}20]. Placement [SBGC20]. Placing [BDK^{+}22]. Plan [LLC^{+}20]. Planar [DR22, GCH^{+}21, LMBM20, WM22]. Plane [QFWH21]. Planes [SWB^{+}20]. Planning [DPB^{+}22b]. Plans [MPSM21, WXL^{+}21]. Plant [GKK^{+}20, HTI22]. Plastic [JSVB22, ZHL^{+}20]. Players [AWB20]. Plots [RSLSRÁ21, TBSB20, vOSRR21, BDL^{+}20]. Point [BGGSSG20, CLR^{+}20, CRZ^{+}22, CCG^{+}22, CK21, DLX^{+}22, FLH^{+}20b, GZLL23, GCH^{+}21, GKK^{+}20, GCLB21, GAD^{+}20, GZG^{+}22, HLJ^{+}20, HCL^{+}22, HXW^{+}22, HMSS22, JG21, KP22, KPLD21, LMBM20, LFH^{+}22, LAK20, MWZ^{+}20, MHG^{+}22, QZL^{+}20, QZLC22, RLG^{+}20, RKW20, SOW20, SKW21, SDR^{+}22, SCZ21, SXHA22, VLLA22, ZWM^{+}20, FCD20, LCBM21, YZ22]. Point-augmented [KP22]. Point-Based [KPLD21]. Point-line [VLLA22]. Point-Pattern [HMSS22]. Point-Spread [CK21]. PointCleanNet [RLG^{+}20]. Points [CLC^{+}22, OPP^{+}21, SMOW20]. Pointset [WCR^{+}23]. PointSkelCNN [QZL^{+}20]. Poisson [KCRH20, Mit22]. Poisson-Disk [Mit22]. Polarization [QFWH21]. Policies [CDMH^{+}22]. Policy [HNWC23, STX^{+}22]. Political [BJSM22]. Polycube [DPM^{+}22]. Polycube-Maps [DPM^{+}22]. Polygon [ADP22, BHKB20]. Polygonal [BB23, VLLA22, LRTC20]. Polygons [GH21]. Polynomials [MMO^{+}21]. pOp [RSP22]. Porous [FSK^{+}22]. Portrait [DLS21, HWDL21, SOL^{+}22, SCCZ21]. Pose [AAC22, MSK23, PZL^{+}22, SMM22, WQL^{+}20, YKL21, ZLL20, FCD20, PZL^{+}22]. Position [Bd22, MMC^{+}20, RMSQ20, TÇE^{+}21, WGT^{+}22]. Position-Based [WGT^{+}22]. Position-Free [Bd22]. Possible [CZGM20]. Post [CM21a]. Post-Reconstruction [CM21a]. PosterChild [CSG21]. Posterization [CSG21, TU22]. Postures [WRZ20]. Power [BAR^{+}21]. Practical [DBA^{+}21, DGJ^{+}20, GGD^{+}20, HPM^{+}20, LLZ^{+}20, MLP21, SW20]. Pre [WDH20, WWZ^{+}22, ZSWL21]. Pre-Filtered [WDH20]. Pre-integration [ZSWL21]. Pre-trained [WWZ^{+}22]. Precise [KMTH22, YLC^{+}22]. Precision [WM22]. Precomputed [CDAS20, RBRD22, SS21]. Precondition [ZCC22]. Precondition-Based [ZCC22]. Predicting [LTY^{+}22]. Prediction [CCG^{+}22, FWHB21, GK22, MSK23, PYAP22, YKL21]. Predictions [WKA^{+}20]. Predicts [RSGP21]. Preferences [HSvO^{+}22, SSKEA21]. Preferred [KFDA21]. Prefilters [ROG20]. Presentation [WGH^{+}21]. Preservation [BHJS23]. Preserving [BJG20, BCD20, FIDF20, LSD23]. Pressure [JSG^{+}22, LCG^{+}20]. Prevention [EMKL21]. Preview [FYFL20]. PriFit [SDR^{+}22]. Primal [MEM^{+}20]. Primal/Dual [MEM^{+}20]. Primitive [HPSK21]. Primitives [SDR^{+}22, LJJ22]. Principal [SHR20]. Printing [JSVB22, LFD^{+}20, RSB^{+}21, TCL20, BCOM^{+}22, FAG^{+}20]. Prior [JZN^{+}22]. Priority [DKVV23]. Priority-based [DKVV23]. Privacy [BCD20]. Privacy-Preserving [BCD20]. Probabilistic [GWE^{+}20, TK20]. ProBGP [USK21]. Problem [XXeL^{+}20]. Problem-solving [XXeL^{+}20]. Problems [FBG^{+}20, ZCC22]. Procedural [CSDD20, GKG^{+}22, GAD^{+}20, GHYZ20, JHC^{+}20, ZA22, GSDT22, MMMR22]. Procedures [EMJ^{+}22, WGT^{+}22]. Process [DWH^{+}22, DSMA^{+}21, GAD^{+}20, MBB22]. Processes [FSK^{+}22, WCR^{+}23, ZWVL22]. Processing [SBB^{+}22, WTB^{+}21, LFD^{+}20]. Product [DGJ^{+}20, HPM^{+}20, XWHM20]. Production [Hil20]. Profile [SRK21]. Profiles [ZDS20, RCDB23]. Profiling [MWLH^{+}20, vBGF22]. Program [YBF22]. Programming [TCT22]. Programs [CSQ^{+}22]. Progressive [CM21a, FFJ22, KWK20, SMOW20, SHV21, USK21, ZAB^{+}21]. Projection [AEC^{+}22, BGG20, GHMR^{+}20, JSG^{+}22, MDVP22, VGdS^{+}20]. Projections [FMKK21, VCT21]. Projective [VLLA22, WTB^{+}21]. Proliferating [Ano23d]. Propagation [HJZ^{+}21, SW22, TBK21, RGR20]. Properties [AHKSH20, BYH^{+}20, FJM^{+}20, YMS^{+}21]. Protein [SFMS^{+}21]. Protein-Ligand [SFMS^{+}21]. Provenance [XOW^{+}20]. Proxies [Sub21]. Pseudo [ZK23]. Public [CM21b, PL20]. Pupils [AMT20]. Pursuit [AL21]. Pyramids [XHZ^{+}20].

Q [Sub21]. Q-NET [Sub21]. QEF [TK20]. QR [PLL^{+}20, PHP22]. Quad [CXWZ22, DLLK21, LBK20, LCK21a, LCK21b]. Quad-fisheye [CXWZ22]. Quadric [TK21]. Quadric-Based [TK21]. Quadrics [TK20]. Quality [BHJS23, DFB^{+}21, JZN^{+}22, QLL^{+}22, SBMS23, HT20, LBFS21, ZXC^{+}20, ZMV^{+}21]. Quantitative [VSC^{+}20, VGdS^{+}20, YGW^{+}21]. Quantization [LCK21a]. Quasi [CRZ^{+}22]. Quasi-Point [CRZ^{+}22]. Queries [JG21]. Query [PPF23]. Question [HKM22]. QUESTO [DXG^{+}20].

Races [DFV21]. Rachford [OPY^{+}20]. RadEx [MWLH^{+}20]. Radial [RSLSRÁ21]. Radiance [CL22a, HSG^{+}22, MBB22, NSP^{+}21, RBRD22, TU22, VBRKW22, WDY^{+}22]. Radiomic [MWLH^{+}20]. Rain [SWW^{+}22a]. Raindrop [RLHS20a]. Random [EMJ^{+}22, HMSS22]. Range [VBRKW22, ÇBSM23]. Rank [LHJ21]. Rank-1 [LHJ21]. Rapid [QLWQ21]. RAS [LLDT20]. Raster [LJSL21, PMC22a, YVK^{+}23]. Rasterization [KRF22, TST^{+}22]. Ratio [HYC^{+}23]. Rational [KMC22, YLC^{+}22]. raw [FCD20]. Ray [AZ21, CRH^{+}21, DBA^{+}21, KMTH22, MOB^{+}21, SH21, WMU^{+}20, ZLY^{+}21]. RayCaching [NMN20]. Reaction [MHLH22]. Reactive [TST^{+}22]. Reader [ZLL^{+}22b]. Reading [BBBD20]. Ready [Hil20]. Real [ARC22, BPD20, BLC20, CSDD20, CTL^{+}20, FWHB21, GOG21, GCDT22, HTT^{+}20, HSG^{+}22, KJB^{+}22, LSY21, LPW^{+}20b, MB20, NSP^{+}21, OLK^{+}21, RDG20, RTK^{+}21, SMRB22, SGAG22, SMOW20, SLCL22, TZX^{+}22, WDH20, WBB^{+}20, YKL21, ZLY^{+}21, ZCC22, ZSWL21, MMMR22]. Real-Time [ARC22, BPD20, CSDD20, CTL^{+}20, GOG21, LSY21, NSP^{+}21, OLK^{+}21, RDG20, SMRB22, SGAG22, SMOW20, SLCL22, TZX^{+}22, WDH20, ZCC22, BLC20, FWHB21, GCDT22, HTT^{+}20, HSG^{+}22, KJB^{+}22, LPW^{+}20b, RTK^{+}21, WBB^{+}20, YKL21, ZLY^{+}21, ZSWL21, MMMR22]. Real2Sim [ANK^{+}23]. Realistic [BPD20, YLHX22]. Reality [BC20, CTJ^{+}20, FFM^{+}21, LY20, NMN20, SGBI22, YLHX22]. Reasoning [KSHL20, SGL23]. Recognition [FLH^{+}20a, FLH^{+}20b, LSP^{+}21, SLL20]. Recognizable [JJZ^{+}22]. Recolorable [TU22]. Recolouring [TÇE^{+}21]. Recommendation [ELW^{+}22]. Reconfigurable [AFTH21]. Reconstructed [MCT21]. Reconstructing [JJZ^{+}22]. Reconstruction [BTCL22, BMPP21, BGGSSG20, CXWZ22, CM21a, Csé23, DBA^{+}21, GGD^{+}20, HSG^{+}22, KCRH20, KKN^{+}21, KHW^{+}22, LHZ^{+}22, LPW^{+}20b, MOW22, NHSS21, NYB^{+}20, OPP^{+}21, PMG^{+}20, SLB22, SLMV21, TKR^{+}23, VLLA22, WQL^{+}20, WLNM21, XSW^{+}20, YT22, YLL^{+}21, ZWG^{+}21, ZLL^{+}22a, ZAB^{+}21, HWC21, YZ22, ZA21]. Recovering [YDPG21]. Recovery [LZBD21, WWZ^{+}22]. Recursive [MMBP22]. Reduce [FAG^{+}20]. Reduced [TRPO21, MN20]. Reducing [BSB^{+}20]. Reduction [BJG20, GvdVS23, RKW20, TACS21]. Reeb [WMS^{+}23]. Ref [HWW^{+}22]. Ref-ZSSR [HWW^{+}22]. Reference [HWW^{+}22, ZK23]. Refinement [CSAN23, DV21, KWK20, LWB^{+}21, LBK20, SSKEA21, WMU^{+}20]. Reflectance [BGP^{+}22, GGD^{+}20, HTFW23, YDPG21]. Reflection [DX22]. Reflections [BCD20, VRBT21]. Region [CMG21, Hei20, KADM22, ZJZ^{+}22]. Regional [EHL21]. Regions [WBL20]. Registration [CCG^{+}22, DYDZ22, GZLL23, MMRC20, QZLC22]. Regression [GZLL23, YGW^{+}21]. Regularisation [WDH^{+}21]. Reinforcement [CXW^{+}21, JVW20, KAK^{+}22]. Rejection [Mit22]. Relationship [MWZ^{+}20]. Relationships [CLC^{+}22]. Relaxation [MPZS20]. Relaxed [LCK21b]. Relevance [HJZ^{+}21]. Reliable [HZ22, ZLY^{+}21]. Relief [WSN^{+}20, ZWW^{+}21]. Relighting [EGH21, GRP22, SCCZ21, TKE21, TÇE^{+}21]. Remeshing [MCC23, ZWG^{+}22]. Remeshing-free [MCC23]. Removal [FGZ^{+}22, RLHS20a, SWW^{+}22a, SXZS21, ZLY^{+}20]. Remove [RLG^{+}20]. Rendered [FFJ22]. Renderer [TST^{+}22]. Renderers [KJB^{+}22]. Rendering [CYD^{+}22, CSDD20, CK21, DX22, DSMA^{+}21, FWBI22, FFM^{+}21, GGS21, GCDT22, Hil20, JC22, KAC^{+}21, KK21, KPLD21, LZBD21, MDVP22, MB20, NSP^{+}21, NMN20, OP22, PRD21, RDG20, SMRB22, SGAG22, SMOW20, SKW21, SAK20, TZX^{+}22, TFT^{+}20, TTM^{+}22, TAA^{+}21, TB21, TÇE^{+}21, VVP20, WH20, WQL^{+}20, WDH20, WHW22, WS20, XWHM20, XSW^{+}20, ZMV^{+}21, BP21, HWD22, LJJ22]. Renderings [ZOM^{+}22]. Reordering [BZP^{+}20]. Repair [MPZS20, WGT^{+}22]. Repairing [XABR21]. Repetitive [LSP^{+}21]. Reports [HLJ^{+}22, KQL^{+}21]. Representation [GLLL20, HGC^{+}20, LPH^{+}20, MTB^{+}22, SCCZ21, SRRW21, XCC22, YTK23, YUIH21, ZLW^{+}22]. Representations [AAC22, CBE20, HW22, LJLB21, WLT^{+}22, WHW22]. Representative [Wan20]. Resampling [MMK^{+}22, NID20, OLK^{+}21, CGG^{+}22]. Research [BCD20]. Researchers [LB20]. Residual [JZZ^{+}20]. Resolution [BTB^{+}22, LSK21, LFS21, LSSW22, MRR^{+}22, NHSW20, ZCC22, IWH23, LHJ^{+}22, ME23]. Resolution-switchable [LSSW22]. Resolving [LXYM20]. Resonance [NBPM20]. Response [CARK20, STX^{+}22]. Responses [BSB^{+}20]. Responsive [KMH21]. ReSTIR [OLK^{+}21]. Restoration [LBB22, RLHS20a, RLHS20b, WYZ^{+}22, IWH23]. Restorations [RDG20]. Restricted [BAR^{+}21]. restrictions [SSS20b]. Results [GBYH20]. Retargeting [KJS^{+}21, MTB^{+}22, RKL^{+}21]. Retargetting [KPKH20]. Reusing [GCL22]. revealing [FCD20]. Revectorization [MAA20].

Revectorization-Based [MAA20]. Reverse [NZCS21]. Review [BYK^{+}21, CRW^{+}20, FLH^{+}20a, FAG^{+}20, GCZ^{+}20, GLLL20, GRGH20, HTT^{+}20, HS20, HLJ^{+}20, HZSQ20, LLL^{+}21, LLDT20, LAK20, LPW^{+}20b, LFD^{+}20, MMRC20, MRWE20, NYB^{+}20, NKA^{+}23, PMG^{+}20, PCDS20, QZL^{+}20, RKW20, RGR20, SLL20, SSV^{+}22, SWB^{+}20, TCL20, WWW20, WTO^{+}20, WYM^{+}20, WYA^{+}20, WBB^{+}20, WS20, ZWM^{+}20]. Revolution [SYKE20]. Reweighting [GE21]. RfX [EMJ^{+}22]. RGB [NYB^{+}20, IIHG^{+}20, KKN^{+}21, SXZS21, SMM22, WLNM21]. RGB-D [KKN^{+}21, WLNM21]. RGB2AO [IIHG^{+}20]. RGBD [LPW^{+}20b]. rh [WCU^{+}20]. rh-Adaptive [WCU^{+}20]. Rhapsody [RMH^{+}22]. Ribbed [TCL20]. Ribbons [JSVB22]. Rich [ZLL^{+}22b]. Rig [ALTT22]. Rigid [CARK20, DYDZ22, KUS^{+}21, MMC^{+}20, PT23, QZLC22, TKR^{+}23, WSP21]. RigidFusion [WLNM21]. Rigidity [LLDT20, LHX^{+}22]. Rigidity-Aware [LLDT20]. Rigidly [WLNM21]. Rigidly-moving [WLNM21]. RNN [BSS20]. Robot [WM22]. Robotically [MWS^{+}20a]. Robotically-Constructed [MWS^{+}20a]. Robust [CBC20, CLSK20, KC22, TK20, UKPW21, WCR^{+}23, ZWG^{+}21, ZSY^{+}22, WYM^{+}20]. Robustness [SLB^{+}21]. Rod [WCU^{+}20]. Rods [KUS^{+}21, ZLW^{+}22]. Role [CWW^{+}22]. Roof [ZA22]. Roominoes [WXL^{+}21]. Rooms [WXL^{+}21]. Root [WFJ^{+}22]. Rotational [ZJA21]. Rough [CCFM23, EBB23]. Rounded [CRH^{+}21]. RT [CRH^{+}21]. Rule [DV21]. Running [CM21b].

Sail [PMB^{+}23]. Sailing [PMB^{+}23]. Salient [ZGP^{+}20]. Sample [GE21, LFR23, MSK23]. Samples [TZX^{+}22]. Sampling [CSAN23, GGS21, HPM^{+}20, HMS^{+}20, HTFW23, JC22, Mit22, Özt20, Pet21, SSC^{+}20, SRRW21, TK21, XWHM20, XK21, LSD23, vdRE21]. Sampling-based [XK21]. Satellite [ZA22, ZK21b]. Satellite-image [ZK21b]. Scalable [BBBD20, Hil20, LCG^{+}20, MO23, SLMV21]. Scalar [FJF20, KRF22, LJLB21, YMS^{+}21]. Scale [AZ21, DDB20, FBG^{+}20, JZN^{+}22, KC22, LWY^{+}21, LLL^{+}22, MVNT21, SFMS^{+}21, SW22, YSLF20, ZCC22, ZZF^{+}22, HXW^{+}22, LMBM20, QLZ^{+}20, RMH^{+}22, VGL^{+}20]. Scale-Invariant [AZ21]. Scales [SSV^{+}22]. Scan [BKL21]. Scanning [YDPG21]. scans [IWH23]. Scatter [TBSB20]. scattered [HWD22]. scatterers [d’E21]. Scattering [Bd22, FJF20, HHH22, LHW21, RSB^{+}21, BP21]. Scattering-Aware [RSB^{+}21]. Scatterplot [RLR^{+}20]. Scene [FLH^{+}20a, LSSW22, LHG^{+}23, WYA^{+}20, WLNM21, ZXS^{+}21, ZJZ^{+}22]. Scene-Aware [LHG^{+}23]. Scenes [FYFL20, GVF^{+}21, HACOG22, ME23, RGR20]. SCGA [RLHS20b]. SCGA-Net [RLHS20b]. Schedule [MRSF20]. Schematic [BDF^{+}22]. Science [LB20]. Scientific [YMS^{+}21]. Scores [SKB^{+}22]. Screen [OP22, SWW^{+}22b, ZLL^{+}22b]. Screen-Space [OP22]. Sculpting [GFZ^{+}20]. SDF [ZLWT22]. SDF-Based [ZLWT22]. SDF-StyleGAN [ZLWT22]. Seamless [Lev22, NTM^{+}21, ZTZC20, ZK21b]. Search [CMKK21, CSSI21, LPH^{+}20, SYHL21]. Second [JSG^{+}22, TCT22]. Second-Order [JSG^{+}22]. Seed [SBGC20]. Seeing [WJK22]. Seeking [AACF22]. SEEVis [LLC^{+}20]. Segment [GGPP20, PYL20]. Segmentation [CMG21, CNS22, EST20, HZ22, LHJ^{+}22, NTM^{+}21, SDR^{+}22, Wan20]. Selection [DLX^{+}22, KF21, LLT^{+}20b, MvdEV22, MSK23, PBM^{+}22, SBGC20, SGBI22, SSS^{+}20a, WCM22, ZOM^{+}22]. Self [JSVB22, NRS22, SCZ21, WCS^{+}22]. Self-Actuated [JSVB22]. Self-Supervised [SCZ21, WCS^{+}22]. Self-supervision [NRS22]. Selfie [LY20]. Semantic [CTJ^{+}20, CLC^{+}22, CNS22, EK20, FLH^{+}20a, LSSW22, WBL20, WWF^{+}22]. Semantic-Guided [WWF^{+}22]. Semantic-Level [CTJ^{+}20]. Semantics [HDP^{+}21, JGvK21]. Semantics-Guided [JGvK21]. Semi [CLH^{+}20, FBG^{+}20, GAD^{+}20, SWW^{+}22a, SWW^{+}22a, SLL20]. Semi-analytical [CLH^{+}20]. Semi-MoreGAN [SWW^{+}22a]. Semi-Procedural [GAD^{+}20]. Semi-Sparse [FBG^{+}20]. Semi-supervised [SWW^{+}22a, SLL20]. Sensitivity [YBRP21, YJS^{+}22]. Sensor [HMGU20]. sensors [ÇBSM23]. Sentiment [HDP^{+}21]. SenVis [YBRP21]. Separating [LCG^{+}20]. Separation [DBP20, PBM^{+}22, RLHS20a]. Separation-Restoration-Fusion [RLHS20a]. SeqDynamics [XXeL^{+}20]. Sequence [MMH^{+}21, SCZ21, YDPG21]. Sequences [GKK^{+}20, XSZX22, ÇBSM23]. Sequential [CSSI21, LPH^{+}20]. Series [BSP^{+}22, EST20, EHL21]. Set [AB20]. Sets [AB20, DFB^{+}21, RKW20, FCD20]. SGLBP [GZG^{+}22]. Shade [NMSS22]. Shader [TST^{+}22, YBF22]. Shaders [LWH^{+}20, SKW21]. Shading [DMT21, HSS21, HYW^{+}22, NMSS22]. Shadow [BJG20, DVP22, FGZ^{+}22, HZ22, MAA20, ZLY^{+}20]. Shadowing [KMTH22]. ShadowPatch [HZ22]. Shadows [GRP22]. Shape [BTW^{+}22, CZG^{+}22b, CL22b, CBC20, DR22, DLX^{+}22, GLLL20, JGvK21, KC22, KMP^{+}21, LJ21, LXY^{+}22, LZBD21, MVNT21, MMCJ22, NYB^{+}20, PKO22, PMC^{+}22b, RCPW21, RMWO21, SLL20, UPE22, WTO^{+}20, WQL^{+}20, WLT^{+}22, ZWG^{+}21, ZCH^{+}23, ZLWT22, ZWM^{+}20]. Shape-Guided [BTW^{+}22]. Shape-to-Shape [WLT^{+}22]. shaped [wKH20]. Shapes [CXW^{+}21, FIDF20, Fre21, JJZ^{+}22, LBB22, MMC^{+}22, YUIH21, ZCLJ22]. Sharing [DLS21, FWHB21]. Sharp [ROG20, TLC20, Ano21a]. Sheet [MFH^{+}22, MRKEA22]. Shells [LMF^{+}20]. Shifting [HLJ^{+}20]. Short [RLR^{+}20]. Short-Contact [RLR^{+}20]. Shot [FLH^{+}20b, HWW^{+}22, LLC^{+}20, SDR^{+}22, GFH^{+}23]. Shrinking [WMS^{+}23]. Shrunk [TACS21]. Siamese [ZWM^{+}20]. SiamesePointNet [ZWM^{+}20]. SIGDT [MOW22]. Signage [JH22]. Signal [MMO^{+}21]. Signals [MSK23, YKL21]. Silhouettes [BJG20]. similar [CHC^{+}23]. Similarity [HL22, XSZX22]. SimilarityNet [HL22]. SimJEB [WBM21]. Simple [BHKB20, SBB^{+}22, SWJG20]. Simpler [LCK21b]. Simplicial [FIDF20]. Simplification [ADP20, ADP22, FIDF20, LLT^{+}20a, SBBB20, VLLA22, ZCLJ22]. Simulated [WBM21]. Simulating [CMSK20, IWH23, MMP20, TCT22, WGT^{+}22]. Simulation [COKE20, CCR^{+}20, CO22, CLPO22, CBL^{+}20, CCFM23, CvTZ^{+}22, DPB^{+}22b, GHMR^{+}20, GZL^{+}20a, GEE20, HXA21, HWJ22, JSG^{+}22, KUKH23, wKH20, KMA20, LWB^{+}21, MMMR22, MCC23, MWM^{+}22, MPE23, MFJ^{+}21, MRWE20, MMC^{+}20, OL21, PYL20, RMSQ20, SW22, SXHA22, VGL^{+}20, WCSH21, ZCC22, ZLW^{+}22, vTP21, SSS20b, RGR20]. Simulations [ANK^{+}23, CARK20, EMKL21, MN20, QMSL20, STX^{+}22]. Simultaneous [XL20]. Single [BMBW21, CSAN23, CGG^{+}22, DBP20, GRP22, HMGU20, HWW^{+}22, KKN^{+}21, KJB^{+}22, LPW^{+}20b, LHG^{+}23, MRR^{+}22, NYB^{+}20, SXZS21, SMM22, WWZ^{+}22, ZA22, ZYP^{+}21, ZK21a, ZWY23]. Single-Image [ZK21a, GRP22]. Single-pass [CGG^{+}22]. Single-View [NYB^{+}20]. Singularities [Lev22, LCK21b]. Situ [ELW^{+}22, HCS20, WCM22]. Situated [CM21b]. Situatedness [MAS22]. Six [DPB^{+}22a]. Skeletal [AAC22, RTK^{+}21]. Skeleton [HLJ^{+}20, KPKH20, MHL^{+}22, QZL^{+}20]. Skeleton-Based [MHL^{+}22]. Skeleton-specific [KPKH20]. Skeletonization [DLX^{+}22]. Skeletons [CW22, CTL^{+}20, TRPO21]. Sketch [BC20, CvTZ^{+}22, SBBB20]. Sketch-based [CvTZ^{+}22]. Sketches [FRH^{+}21, KHW^{+}22]. Sketching [DZN^{+}21, MBA22]. Sketchy [HCS20]. SketchZooms [NODI21]. Skills [MYT^{+}21, XLKvdP20]. Skin [CMM^{+}20, GGD^{+}20, ROCP20]. Skinned [UKPW21]. Skinning [LLDT20, LTDG20, RTK^{+}21]. Skip [RLHS20b]. Sky [Hil20, VBRKW22]. Skyscapes [VGL^{+}20]. Slabs [Bd22]. Slice [RGW20]. sliced [LJJ22]. Slices [SWB^{+}20]. sliding [Ano23d]. Smart [LLC^{+}20]. Smartphones [SRM^{+}21]. Smith [Bd22]. Smoke [KHW^{+}22, MN20, OL21]. Smooth [BSH22, WCR^{+}23]. Smoothing [WWF^{+}22]. Smoothly [ECC^{+}20]. SnakeBinning [HSS21]. Snapshot [AFTH21]. Snapshots [FYFL20]. Social [LB20, PPF23, YYH20]. Soft [ANK^{+}23, MAA20, SGOC20, TST^{+}22, TRPO21, WZB20]. Soft-Tissue [TRPO21, WZB20, SGOC20]. SoftSMPL [SGOC20]. Solar [TÇE^{+}21]. Solid [CLH^{+}20, GRGH20, LCG^{+}20]. Solids [KUKH23, LLJ^{+}20, SYKE20]. Solutions [LHW21]. Solver [KUKH23]. Solvers [LCG^{+}20]. solving [XXeL^{+}20]. Sorting [BHJS23]. sound [RGR20]. Sounds [WJK22]. soundscapes [RMH^{+}22]. Source [LSK23, PBM^{+}22, TBK21]. Sources [SW22]. Space [ABPP23, BGG20, BLC20, JGvK21, KJRK21, LZ22, MCK^{+}22, MFJ^{+}21, NMSS22, OP22, SYPB21, SRC^{+}20, TC21, TLLS20, WDH^{+}21, WKA^{+}20, BCOM^{+}22]. space-filling [BCOM^{+}22]. Spaces [CHC^{+}23, SKB^{+}22]. Spacetime [MYT^{+}21]. Sparse [CBE20, CL22a, DSA20, FBG^{+}20, LHZ^{+}22, MSK23, MRSF20, SHR20, YKL21, YGW^{+}21]. Sparse-View [LHZ^{+}22]. Sparsely [HSvO^{+}22]. Sparsity [TSK^{+}20]. Sparsity-Specific [TSK^{+}20]. Spatial [BDF^{+}22, BYK^{+}21, Cav21, MWZ^{+}20, PBM^{+}22, RSGP21, SGBI22, TCG^{+}22].

Spatially [YDPG21]. Spatially-varying [YDPG21]. Spatio [CSB^{+}20, DPÖM22, EHL21, FMKK21, HWJ22, HL22, SSS^{+}20a]. Spatio-Directional [DPÖM22]. Spatio-Temporal [CSB^{+}20, SSS^{+}20a, EHL21, FMKK21, HWJ22, HL22]. Spatiotemporal [KKN^{+}21]. SPCNet [HCL^{+}22]. Specialized [MWS^{+}20b]. Specific [MRB^{+}21, TSK^{+}20, KPKH20]. Spectral [GGD^{+}20, HRWO20, JHD20, LLT^{+}20a, LHL^{+}21, MMC^{+}22, PMC^{+}22b, TWF22, VBRKW22, vdRE21]. Specular [JC22]. Speech [AHKB20, GFH^{+}23, NKA^{+}23]. Speech-Driven [AHKB20]. SPH [BGT20, KUKH23, KBST22, LWB^{+}21, NHSS21, PRD21, PT23]. Sphere [LHW21, MBB22]. Sphere-Tracing [LHW21]. Spheres [Lev22, QSBJ22]. Spherical [CDAS20, HJL^{+}20, LJ21, MDVP22, MMBP22, d’E21]. Splatting [CK21]. Splitting [OPY^{+}20, RW21]. Spread [CK21]. Spring [MPE23]. Squares [FBG^{+}20, MPZS20, ZJA21]. SREC [CRH^{+}21]. SREC-RT [CRH^{+}21]. SRF [MWZ^{+}20]. SRF-Net [MWZ^{+}20]. SRNet [FLH^{+}20a]. Stability [BS22, RKW20]. Stable [LCBM21, VCT21, WKA^{+}20, YY20, ZLW^{+}22]. Stackings [MB20]. Stage [DFV21, GVF^{+}21, LWZ^{+}20, NID20]. STALP [FKL^{+}21]. Star [ML20, RSLSRÁ21, WL22]. State [AMA^{+}22a, BGA^{+}22, BYK^{+}21, BCD20, BYH^{+}20, CMJ^{+}20, HNWC23, HHS20, HKM22, LBB^{+}23, TFT^{+}20, TKR^{+}23, WSP21, WL22, PMG^{+}20]. State-Of-the-Art [WL22, BGA^{+}22, PMG^{+}20]. Static [BKL21, FLH^{+}20a, MWS^{+}20b, WGH^{+}21]. Stationary [CLCO^{+}22, TAA^{+}21]. Statistics [CM21b, LWH^{+}20, YYH20]. Statistics-based [YYH20]. statues [FCD20]. Steering [CDMH^{+}22, CvTZ^{+}22]. Step [Csé23, OL21, SSS20b]. Stepping [XLKvdP20]. Steps [OKN21, WCM22]. Stepwise [HCL^{+}22]. Stereo [CRZ^{+}22, HZ22, WSN^{+}20, WM22]. Stitching [CXWZ22]. Stochastic [CLSK20, PRD21, TDDB23, WDH20]. Stone [XLKvdP20]. Storage [HTFW23]. Stories [LCB21]. Storytelling [LPW20a]. Strategies [AWB20, BZP^{+}20, HPKI22]. Strategy [BMPP21]. Stratified [GWH20, CGG^{+}22]. stream [ZSY^{+}22]. Streaming [HSS21, WCM22]. Streamline [SBGC20]. Streamlines [SWB^{+}20]. Streams [AB20]. Stress [WWW20]. Stretched [JSVB22]. Stroke [EMKL21, JSL22]. Stroke-Based [JSL22]. Strokes [JSL22]. Structural [BMBW21, CLSK20, GCZ^{+}20]. Structure [CRH^{+}21, LMBM20, LPW20a, MDBH22, TBSB20, WWW20, ZPIS22]. Structured [BGMC22, DBP20, PMG^{+}20]. Structures [LNRH20, LL20, MWS^{+}20a, SMOW20, SMS^{+}20, TCG^{+}22, CRW^{+}20]. Studies [MWLH^{+}20, VRBT21]. Study [ANI^{+}20, EMJ^{+}22, NMSS22, ZWLY22, ZWVL22]. Style [AHKB20, FKL^{+}21, WYA^{+}20, KMTM22]. Style-Controllable [AHKB20]. StyleGAN [BGA^{+}22, End22, ZLWT22]. StylePortraitVideo [SOL^{+}22]. StyleProp [HTT^{+}20]. Stylised [DMT21]. Stylization [BSB^{+}20, GVF^{+}21, HTT^{+}20, KFDA21]. Stylized [FRH^{+}21, RTK^{+}21].

Stylizing [PCS21]. Sub [NID20]. Sub-paths [NID20]. Subdivision [DV21, KP23, MWS^{+}20b, TLC20, WKA^{+}20, ZHHK22, Ano21a, KP22]. Subdivision-Specialized [MWS^{+}20b]. Subgraph [GZG^{+}22]. Subgraph-based [GZG^{+}22]. Sublinear [MHT^{+}20]. Subpixel [YVK^{+}23]. Subset [XTL^{+}21]. Subsets [TDDB23]. Subspace [CSSI21]. Subsurface [LHW21]. Sum [MPZS20]. Sum-of-Squares [MPZS20]. Summarizations [YGY^{+}21]. Summary [KQL^{+}21]. Summed [RW21]. SumRe [KQL^{+}21]. Sunset [TÇE^{+}21]. Sunspot [TBSB20]. Superpixels [ZXC^{+}20]. Superresolution [HWW^{+}22]. Supersampling [GCDT22]. Supervised [NYB^{+}20, SCZ21, SWW^{+}22a, WCS^{+}22, SLL20]. supervision [NRS22]. Support [CMKK21, TCL20, ZWLY22]. supported [NBH20]. Surface [BMPP21, BSCK21, BGGSSG20, CLH^{+}20, HW22, KCRH20, KFDA21, LCP^{+}23, MRCK21, NTM^{+}21, SPK23, SW22, SBH22, SLMV21, ZAB^{+}21, YZ22]. Surface-Only [SBH22]. Surfaces [AZ21, BRP^{+}20, CCFM23, JSVB22, LSP^{+}21, LZ22, MLP21, MCBR20, NA22, SH21, SMS^{+}20, SRC^{+}20, TLC20, TK21, WCR^{+}23, WMS^{+}23, WCSH21, Ano21a, KP22, LRTC20, MRWE20]. SurfNet [HW22]. Surgery [DPB^{+}22b]. Surgical [BSB^{+}20]. Surround [YJS^{+}22]. Surround-aware [YJS^{+}22]. Survey [BTB^{+}22, BC20, BB23, CBL^{+}20, DYDZ22, DSMA^{+}21, EGH21, EMML22, FJM^{+}20, FAP^{+}22, GAM^{+}21, GSWS21, HLJ^{+}22, KKM^{+}20, KBST22, KFS^{+}22, KAK^{+}22, LBC^{+}22, LCO20, MOB^{+}21, MHL^{+}22, OPP^{+}21, PL20, SBGC20, SYPB21, SBMS23, SEAG^{+}21, VVP20, WNT^{+}20, XOW^{+}20, YLS20, BD23]. SVBRDF [LZBD21, WWZ^{+}22, ZK21a]. Swapping [NHSW20, ONH^{+}22]. switchable [LSSW22]. Synchronization [HRWO20]. Synthesis [AHKB20, DX22, EK20, GWE^{+}20, GDG^{+}23, GRGH20, HMSS22, HPSK21, HPKI22, JPK20, MA23, PWBD21, SHR20, WBZB20, WDY^{+}22, YPZ^{+}22, YYH20, ZYP^{+}21, ZWY23, ZK21b, CL22a, LSD21, RZZ^{+}22]. Synthesizing [ECC^{+}20, FTA22, GMH22, LG22, PGP^{+}21, TL21, ZPIS22]. System [EMML22, HTI22, LLC^{+}20, LY20, ZWG^{+}21]. Systematic [SSV^{+}22]. Systems [BGG20].

T [LCK21a]. T-Mesh [LCK21a]. Table [McN21]. Tables [RW21]. Tangent [DCMO22, SRC^{+}20]. tangential [NBH20]. Tangible [SGBI22]. Target [SYG20]. Target-Adaptive [SYG20]. Targeting [UPE22]. Tasks [BNR20, DXG^{+}20, HLJ^{+}22, KAC^{+}21, PMCEA^{+}22]. Taxonomies [GMOB22]. Technique [Hil20]. Techniques [CPK23, KCW^{+}21, KMTH22, SBGC20, VGdS^{+}20, YLS20, RKW20]. Teeth [RDG20]. Template [LCP^{+}23]. Templates [SPDM20]. Temporal [CSB^{+}20, GKK^{+}20, GCDT22, HMS^{+}20, OL21, SSS^{+}20a, YLS20, EHL21, FMKK21, HWJ22, HL22]. Temporally [Fre21, HSS21, LJJ22, ZLY^{+}21]. Tensioned [LMF^{+}20]. Tensor [HBK^{+}21, SHV21, WWW20, YBRP21, NBH20]. Tensor-based [YBRP21]. Terrain [GPM^{+}22]. Test [ZK23]. Test-Time [ZK23]. Testing [EMJ^{+}22, SR20]. Tetrahedral [AHKSH20]. Tex [BGP^{+}22]. Text [BJSM22, HDP^{+}21, LCB21, MZAD^{+}20, ZXS^{+}21]. Text-based [ZXS^{+}21]. Texture [CSAN23, Csé23, GAD^{+}20, GRGH20, KKN^{+}21, LSE21, MCT21, NMSS22, ONH^{+}22, SWW^{+}22b, LSD23]. Texture-Space [NMSS22]. Textured [LSP^{+}21, WM22, ZXC^{+}20]. Textures [BGP^{+}22, CDAS20, GGMP21, GAD^{+}20]. Texturing [DSA20, MMMR22]. their [ARC22, LPW20a]. Them [NMSS22]. Thematic [CM20]. Theory [FJF20, Özt20, LSD21]. Thickness [KK21, ZLL^{+}22a]. Thin [HXA21, HHH^{+}21, SXZS21]. Thin-Volume [HHH^{+}21]. Thinning [BVHSH21]. Three [LWY^{+}21, ZSY^{+}22]. Three-Scale [LWY^{+}21]. Three-stream [ZSY^{+}22]. Thunderstorms [DPR^{+}21]. Tiled [CK21, NA22]. Tiles [LLL^{+}22]. tilings [LSD23]. Time [ARC22, BPD20, BSP^{+}22, BYH^{+}20, CSK^{+}23, CSDD20, CTL^{+}20, EST20, EHL21, FFM^{+}21, GOG21, HS20, HS21, LSY21, LLJ^{+}20, MHT^{+}20, MB20, Mit22, NSP^{+}21, OM20, OLK^{+}21, RDG20, SJS20, SMRB22, SGAG22, SMOW20, SSS^{+}20a, SLCL22, TZX^{+}22, VSC^{+}20, VGdS^{+}20, WDH20, WKA^{+}20, WCM22, YY20, ZCC22, ZK23, BLC20, FWHB21, GCDT22, HTT^{+}20, HSG^{+}22, KJB^{+}22, LJJ22, LPW^{+}20b, MMMR22, RTK^{+}21, SSS20b, WBB^{+}20, YKL21, ZLY^{+}21, ZSWL21]. Time-Dependent [BYH^{+}20, HS20, HS21, VSC^{+}20, VGdS^{+}20]. time-of-flight [LJJ22]. Time-Series [EST20]. Time-Varying [WCM22]. Time-Warped [FFM^{+}21]. Time-Warping [SSS^{+}20a]. TinyAD [SBB^{+}22]. Tissue [TRPO21, WZB20, SGOC20]. Together [WYZ^{+}22]. TogetherNet [WYZ^{+}22]. tokenized [KMTM22]. tone [WCS^{+}22]. Tongue [ALTT22]. Tool [SLJ^{+}20]. Tools [JM22, ZWLY22]. Tooth [MWZ^{+}20]. Topic [SSKEA21]. TopoAct [RCPW21]. Topographic [PSK^{+}20]. Topography [EBD21]. Topological [MSY^{+}21, YMS^{+}21, ZCLJ22]. Topology [BTCL22, BGGSSG20, BYH^{+}20, CZG^{+}22b, HS21, WDY^{+}22, ZZF^{+}22, ZLL^{+}22a, ZPIS22, HS20]. Topology-Aware [BGGSSG20]. Topology-Independent [CZG^{+}22b]. Topology-Varying [WDY^{+}22]. TopoNet [BTCL22]. Toric [BLC20]. Touch [RLR^{+}20, SGBI22]. Touch-Manipulation [RLR^{+}20]. Touch/Tangible [SGBI22]. TourVis [DFV21]. TRAC [SFB20]. Traces [BBBD20, YBF22]. Tracing [AZ21, CRH^{+}21, DBA^{+}21, FJF20, GGPP20, KMTH22, LHW21, LWZ^{+}20, MOB^{+}21, NID20, OLK^{+}21, SH21, STX^{+}22, WMU^{+}20, ZLY^{+}21]. Tracking [GK22, HZSQ20, KC22, MSY^{+}21, MSK23, NA22, NRS22, SYG20, WBB^{+}20, YKL21]. Tractograms [SMRB22]. Tractography [GvdVS23]. Trade [KMH21]. Trade-Offs [KMH21]. Traffic [CBL^{+}20, HWJ22, ZWLY22, ZWVL22]. trained [WWZ^{+}22]. Training [CZGM20, ZK21a, SLL20]. Trajectories [HS20, MPSM21, WWRT21]. Trajectory [SFB20]. Transfer [AFTH21, CZGM20, DDB20, FRH^{+}21, FKL^{+}21, GCLB21, HPG^{+}23, MMO^{+}21, RBRD22, SAK20, ZLL20, ZCH^{+}23, KMTM22]. Transform [LBKP21, XL20]. Transformable [LNRH20]. Transformation [MMH^{+}21, VEPG23]. Transformations [AKDW22]. Transformed [DGJ^{+}20]. Transformer [End22, LHJ^{+}22, PWBD21, YZ22]. Transformers [CZG^{+}22a, CZG^{+}22b]. Transforming [BSU^{+}22, DPB^{+}22a, CSK^{+}23]. Transient [EBB23]. Transit [WNT^{+}20]. Transition [HPKI22, TBL^{+}22]. Transitions [MTL^{+}21]. Translation [WLT^{+}22, XL20]. Translucent [FJM^{+}20]. Transparent [ZWY23]. Transport [BD23, GGN20, GKK^{+}20, GWH20, GEE20, QZLC22, SS21]. Traversal [JG21]. Tree [AAC20, LL20, WMU^{+}20]. Tree-Based [WMU^{+}20]. Tree-Like [LL20]. Treemaps [VSC^{+}20]. Trees [CDK20, LWL^{+}20, VEPG23, WLG22, LCI^{+}21]. Trends [GKV^{+}22]. Triage [KQL^{+}21]. Triangle [CMM^{+}20, Hei20, HSS21, SC20, ZSL^{+}22, DKVV23]. Triangle-Cut [Hei20]. Triangles [CCR^{+}20]. triangular [HT20]. Triangulated [PHP22]. Triangulation [XCC22]. Triangulations [SPK23]. Trilinear [Csé23]. Trimmed [SH21]. Triplet [SHJD22]. Triquadratic [Csé23]. Trust [CMJ^{+}20]. Truth [WWRT21]. Try [KJB^{+}22, VSGC20]. Try-On [VSGC20]. Tumor [MWLH^{+}20]. Tuning [DDB20, MRSF20]. Tunnel [WMS^{+}23]. Turbulent [LWB^{+}21]. Two [FSK^{+}22, LWZ^{+}20, NID20, OL21]. Two-Phase [FSK^{+}22]. Two-stage [LWZ^{+}20, NID20]. Two-step [OL21]. type [Ano22b, ML22]. Typology [PMCEA^{+}22].

ULMPM [SXHA22]. Uncertainty [CCG^{+}22, EHL21, GSWS21, GvdVS23, SHV21]. Uncertainty-aware [CCG^{+}22, EHL21, GSWS21]. Underemphasized [BJSM22]. UnderPressure [MHLH22]. Undersampling [CPK23]. Understandable [EST20]. Understanding [JVW20, JH22, LCB21, MSY^{+}21, TLLS20]. Unified [LZBD21, RGJW20]. Uniform [TLC20, Ano21a]. Unions [MMC^{+}22]. Unit [ZSL^{+}20, ZYP^{+}21]. Universe [LYA^{+}21]. Unobtrusive [PLL^{+}20]. Unsupervised [JZN^{+}22, KJS^{+}21, LPH^{+}20, LXY^{+}22, LCP^{+}23, MTB^{+}22, WLT^{+}22, XSW^{+}20]. Unsynchronized [TAA^{+}21]. Untrained [NQHH20]. Updated [SXHA22]. Updates [USK21, ZJA21]. Uplifting [TWF22]. Upper [ARC22, YKL21, ZSL^{+}22]. Upper-Body [ARC22, YKL21]. Upright [CXW^{+}21]. UprightRL [CXW^{+}21]. Upsampling [GKK^{+}20]. Urban [QLL^{+}22, RMH^{+}22, VLLA22, WBL20, HWC21]. Use [AKA^{+}20, CMJ^{+}20, SSV^{+}22]. User [End22, GK22, NQHH20, SSKEA21, XOW^{+}20]. User-Controllable [End22]. Users [YLHX22]. Using [AHKB20, AFTH21, ABPP23, BHJS23, CZG^{+}22b, CPK23, CRGFT21, DGJ^{+}20, DPÖM22, FJF20, GGPP20, GK22, GFZ^{+}20, GAD^{+}20, GRGH20, KKN^{+}21, LBB22, LPH^{+}20, LWH^{+}20, LWY^{+}21, LPW^{+}20b, MA23, NTM^{+}21, OL21, PCS21, PT23, SW22, TU22, WGT^{+}22, Wan20, WBB^{+}20, WWZ^{+}22, XL20, YGW^{+}21, ZCC22, ZK23, ZWG^{+}22, ZCH^{+}23, ZSWL21, CYD^{+}22, CSAN23, CZG^{+}22a, CSSI21, CK21, DLC^{+}22, DPR^{+}21, DSA20, EBB23, FLH^{+}20a, GWE^{+}20, GMH22, HWJ22, HZSQ20, HMSS22, JG21, LZBD21, LSD23, LCK21b, Mit22, NSP^{+}21, NBPM20, NRS22, PHP22, RGR20, SPDM20, SMM22, SL23, SBH22, TBL^{+}22, TK20, WBZB20, WQL^{+}20]. USTNet [WLT^{+}22]. UTOPIC [CCG^{+}22]. UV [DSA20]. UV-free [DSA20].

v [BDL^{+}20]. v-plots [BDL^{+}20]. VA [SFB20]. VA-TRAC [SFB20]. VAEs [EK20]. Validation [MLD^{+}21]. Value [RSLSRÁ21]. Values [ANI^{+}20]. Valve [WGT^{+}22]. Variable [GWE^{+}20, LFS21]. Variables [BRP^{+}20]. Variance [LWH^{+}20]. Variational [MSK23, Özt20, WMS^{+}23]. Variations [AAC20]. Varying [WDY^{+}22, WCM22, YDPG21]. Vaults [TCL20]. Vector [ADP20, HS20, HS21, LG22, LTH20, RSP22, SWJG20, ZSS22, GSDT22, NBH20]. Vectorization [SBBB20]. Vectorizing [ZLL^{+}22a]. Vectors [ZLY^{+}21]. Vehicle [SLJ^{+}20, MLD^{+}21]. Velocity [GHMR^{+}20, RTK^{+}21]. Verification [SFB20]. Versus [KMTH22]. Vertex [SHJD22, SKW21, ZWG^{+}22]. Very [YSLF20]. Very-Large-Scale [YSLF20]. Vessel [EMML22, SFB20]. Via [BRMO22, SRC^{+}20, XHZ^{+}20, BVHSH21, BSK21, CXW^{+}21, DLS21, DX22, DBA^{+}21, EK20, HW22, LWZ^{+}20, LSK21, LG22, LY20, LWB^{+}21, LCK21a, MCBR20, MPZS20, PWBD21, PKO22, SHR20, SPK23, SRM^{+}21, SLCL22, SLL20, TL21, TBK21, VLLA22, WCS^{+}22, WYZ^{+}22, WLT^{+}22, WSN^{+}20, XTL^{+}21, YPZ^{+}22, YT22, YTK23, ZLL^{+}22a]. VIBE [MCK^{+}22]. VICE [GWB^{+}21]. Video [AJLN22, CETC22, ÇBSM23, HMGU20, MTB^{+}22, NHSS21, ONH^{+}22, SSX^{+}21, SLCL22, TAA^{+}21, YDPG21, ZK23]. Video-Based [TAA^{+}21]. Videos [BSB^{+}20, CLL^{+}22, CWW^{+}22, GOG21, SOL^{+}22]. View [DX22, FWBI22, KPLD21, LHZ^{+}22, NYB^{+}20, WRZ20, ZWG^{+}21, ZWY23, NODI21, SLS^{+}21]. Viewer [CLL^{+}22]. Viewpoint [HSvO^{+}22, SHVR21]. Views [AKA^{+}20, CL22a, QLWQ21]. Violence [EBD21]. VirtuAl [RZZ^{+}22, BC20, FFM^{+}21, JPK20, KAC^{+}21, KJB^{+}22, LLL^{+}21, LBC^{+}22, MDVP22, NMN20, TBK21, VSGC20, YLHX22]. Virtual-Try-On [KJB^{+}22]. Virtuality [FAP^{+}22]. Vis [LLT^{+}20b]. VisEvol [CMKK21]. Visibility [GEE20, TU22, WH20]. Visibility-Weighted [TU22]. Vision [BD23, JH22, KAC^{+}21]. Visual [AB20, ANI^{+}20, ABK^{+}20, AACF22, Ano22b, AEC^{+}22, ATL^{+}23, BNR20, BSP^{+}22, CSB^{+}20, CBL^{+}20, CMKK21, CPK23, DPR^{+}21, FAG^{+}20, FSS^{+}21, FMKK21, FSK^{+}22, GWB^{+}21, HNWC23, HDP^{+}21, HTI22, HJZ^{+}21, KSHL20, KOYSS21, LBB^{+}23, LPH^{+}20, LLLB22, LXYM20, LLT^{+}20b, LPW20a, MTL^{+}21, ML22, MFH^{+}22, MRKEA22, MWLH^{+}20, NHSW20, dQNSVE20, PMCEA^{+}22, PBM^{+}22, PPF23, PL20, QLL^{+}22, SJS20, SFMS^{+}21, SKB^{+}22, SYG20, STX^{+}22, SYHL21, USK21, WJK22, XXeL^{+}20, XTS^{+}22, YGY^{+}21, ZWLY22, ZWVL22, MCK^{+}22]. VisuaLint [HCS20]. Visualization [AWB20, ADP20, BJSM22, BTB^{+}22, BCD20, CMMK20, CM21b, DFV21, EBD21, ELW^{+}22, EHL21, GKV^{+}22, GMOB22, GSWS21, GKK^{+}20, HBK^{+}21, HHH^{+}21, HHS20, JM22, KQL^{+}21, KCW^{+}21, KJRK21, KMH21, LYA^{+}21, LCB21, LLC^{+}20, LL20, LCO20, LGS^{+}22, MHT^{+}20, MHG^{+}22, MAS22, NQHH20, PMB^{+}23, SLJ^{+}20, SHV21, SGL23, SSV^{+}22, TC21, WL22, WWRT21, XOW^{+}20, YMS^{+}21, YBRP21, ZLL^{+}22b, EBM22, BYK^{+}21, WWW20]. Visualization-Text [LCB21]. Visualizations [BBBD20, CMJ^{+}20, EMML22, HLJ^{+}22, KKM^{+}20, KFS^{+}22, LB20, LGS^{+}22, LTH20, RSGP21]. Visualizing [CLL^{+}22, EMKL21, LZ22, OM20, SYPB21, SWB^{+}20, SZM^{+}21, WBL20, BvOdK^{+}22, RKW20]. Visually [EST20, RCPW21]. Vocabulary [MBA22]. Voice2Face [ALTT22]. Volume [BGMC22, HHH^{+}21, MOB^{+}21, PRD21, RW21, SAK20, VEPG23, WHW22, WCM22, HWD22]. Volumes [KRF22]. Volumetric [CW22, FWBI22, LHW21, LJLB21, TU22, XSW^{+}20, ZOM^{+}22]. Voronoi [BBB22, CO22]. Vorticity [LWB^{+}21]. Voxel [CBE20, DSA20, ME23]. VoxPoint [WTO^{+}20]. VR [GK22, PYAP22].

Walk [MPSM21, QSBJ22]. Walk2Map [MPSM21]. Wall [RLR^{+}20]. Wang [LLL^{+}22]. Warehouse [LLT^{+}20b]. Warp [LCP^{+}23]. Warped [FFM^{+}21]. Warping [SSS^{+}20a]. Warps [HPM^{+}20]. Wassersplines [ZSS22]. Water [JHC^{+}20, SW22]. Wave [SW22, RGR20]. Wavefront [CK21]. Wavelet [KMP^{+}21, LHL^{+}21]. Wavelet-based [KMP^{+}21]. Waves [JHC^{+}20]. Way [ARC22]. Weak [ZGP^{+}20]. Weakly [WM22, NYB^{+}20]. Weaver [TB21]. Web [KOYSS21]. Weight [DLS21, FWHB21]. Weighted [TU22]. Where [BvOdK^{+}22]. Wide [TWF22, VBRKW22]. Wild [MRR^{+}22, MAS22, TKE21, SL23]. Will [RPPMCG23]. Wires [KUS^{+}21]. wise [EK20, HJZ^{+}21, NYB^{+}20, ZQW^{+}20]. Within [HL22]. Without [SMOW20, AMA22b, Mit22, SGAG22, SXHA22]. Witkin [Kim20]. Wizard [TZP^{+}22]. Workflow [RGW20]. Workflows [GCL22]. Workplaces [LLLB22]. world [HWC21]. Worlds [Cav21, KAC^{+}21]. Worst [LHX^{+}22, ZZF^{+}22]. Worst-Case [LHX^{+}22, ZZF^{+}22]. Write [TL21, ZXS^{+}21]. Write-An-Animation [ZXS^{+}21]. WTFM [LXY^{+}22].

XFEM [CMSK20].

Yarn [PRCO20]. Yarn-Level [PRCO20]. Yarns [CCR^{+}20].

Z2P [MHG^{+}22]. ZerNet [SRC^{+}20]. Zernike [SRC^{+}20]. Zero [GFH^{+}23, HWW^{+}22]. Zero-Shot [HWW^{+}22, GFH^{+}23]. ZeroEGGS [GFH^{+}23]. Zome-Fab [SCHC20]. Zometools [SCHC20]. Zoom [HRWO20]. ZoomShop [LADH22]. ZSSR [HWW^{+}22].

References

Argudo:2020:IBT

[AAC20] Oscar Argudo, Carlos Andújar, and Antoni Chica. Imagebased tree variations. Computer Graphics Forum, 39(1):174–184, February 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Andreou:2022:PRD

[AAC22] N. Andreou, A. Aristidou, and Y. Chrysanthou. Pose representations for deep skeletal animation. Computer Graphics Forum, 41(8):155–167, December 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Andrienko:2022:SPV

[AACF22] N. Andrienko, G. Andrienko, S. Chen, and B. Fisher. Seeking patterns of visual pattern discovery for knowledge building. Computer Graphics Forum, 41(6):124–148, September 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Agarwal:2020:SSV

[AB20] Shivam Agarwal and Fabian Beck. Set streams: Visual exploration of dynamic overlapping sets. Computer Graphics Forum, 39(3):383–391, June 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Amirkhanov:2020:VAD

[ABK^{+}20] Aleksandr Amirkhanov, Matthias Bernhard, Alexey Karimov, Sabine Stiller, Andreas Geier, M. Eduard Gröller, and Gabriel Mistelbauer. Visual analytics in dental aesthetics. Computer Graphics Forum, 39(7):635–646, October 2020. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Aliari:2023:FEU

[ABPP23] Mohammad Amin Aliari, Andre Beauchamp, Tiberiu Popa, and Eric Paquette. Face editing using part-based optimization of the latent space. Computer Graphics Forum, 42(2):269–279, May 2023. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Amiraghdam:2020:LLA

[ADP20] Alireza Amiraghdam, Alexandra Diehl, and Renato Pajarola. LOCALIS: Locally-adaptive line simplification for GPU-based geographic vector data visualization. Computer Graphics Forum, 39(3):443–453, June 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Amiraghdam:2022:LLO

[ADP22] Alireza Amiraghdam, Alexandra Diehl, and Renato Pajarola. LOOPS: LOcally optimized polygon simplification. Computer Graphics Forum, 41(3):355–365, June 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Appleby:2022:HIV

[AEC^{+}22] G. Appleby, M. Espadoto, R. Chen, S. Goree, A. C. Telea, E. W. Anderson, and R. Chang. HyperNP: Interactive visual exploration of multidimensional projection hyperparameters. Computer Graphics Forum, 41(3):169–181, June 2022. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Alghamdi:2021:TDL

[AFTH21] Masheal Alghamdi, Qiang Fu, Ali Thabet, and Wolfgang Heidrich. Transfer deep learning for reconfigurable snapshot HDR imaging using coded masks. Computer Graphics Forum, 40(6): 90–103, September 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Alliez:2023:E

[AH23] Pierre Alliez and Helwig Hauser. Editorial. Computer Graphics Forum, 42(1):5–6, February 2023. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Alexanderson:2020:SCS

[AHKB20] Simon Alexanderson, Gustav Eje Henter, Taras Kucherenko, and Jonas Beskow. Style-controllable speech-driven gesture synthesis using normalising flows. Computer Graphics Forum, 39(2):487– 496, May 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Alexa:2020:PLO

[AHKSH20] Marc Alexa, Philipp Herholz, Maximilian Kohlbrenner, and Olga Sorkine-Hornung. Properties of Laplace operators for tetrahedral meshes. Computer Graphics Forum, 39(5):55–68, August 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ashtari:2022:DVC

[AJLN22] Amirsaman Ashtari, Raehyuk Jung, Mingxiao Li, and Junyong Noh. A drone video clip dataset and its applications in automated cinematography. Computer Graphics Forum, 41(7):189– 203, October 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Aurisano:2020:MOC

[AKA^{+}20] Jillian Aurisano, Abhinav Kumar, Abeer Alsaiari, Barbara Di Eugenio, and Andrew Johnson. Many at once: Capturing intentions to create and use many views at once in large display environments. Computer Graphics Forum, 39(3):229–240, June 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Atanasov:2022:MT

[AKDW22] A. Atanasov, V. Koylazov, R. Dimov, and A. Wilkie. Microsurface transformations. Computer Graphics Forum, 41(4):105–116, July 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Azencot:2021:DDA

[AL21] Omri Azencot and Rongjie Lai. A data-driven approach to functional map construction and bases pursuit. Computer Graphics Forum, 40(5):97–110, August 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Aylagas:2022:VAD

[ALTT22] Monica Villanueva Aylagas, Hector Anadon Leon, Mattias Teye, and Konrad Tollmar. Voice2Face: Audio-driven facial and tongue rig animations with cVAEs. Computer Graphics Forum, 41(8):255–265, December 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Alderighi:2022:SAC

[AMA^{+}22a] T. Alderighi, L. Malomo, T. Auzinger, B. Bickel, P. Cignoni, and N. Pietroni. State of the art in computational mould design. Computer Graphics Forum, 41(6):435–452, September 2022. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Arleo:2022:EBD

[AMA22b] A. Arleo, S. Miksch, and D. Archambault. Event-based dynamic graph drawing without the agonizing pain. Computer Graphics Forum, 41(6):226–244, September 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Akita:2020:CLD

[AMT20] K. Akita, Y. Morimoto, and R. Tsuruno. Colorization of line drawings with empty pupils. Computer Graphics Forum, 39(7):

601–610, October 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Alemzadeh:2020:VAM

[ANI^{+}20] S. Alemzadeh, U. Niemann, T. Ittermann, H. Völzke, D. Schneider, M. Spiliopoulou, K. Bühler, and B. Preim. Visual analysis of missing values in longitudinal cohort study data. Computer Graphics Forum, 39(1):63–75, February 2020. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Arnavaz:2023:DDR

[ANK^{+}23] K. Arnavaz, M. Kragballe Nielsen, P. G. Kry, M. Macklin, and K. Erleben. Differentiable depth for Real2Sim calibration of soft body simulations. Computer Graphics Forum, 42(1):277–289, February 2023. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2020:FMa

[Ano20a] Anonymous. Front matter. Computer Graphics Forum, 39(2): i–xx, May 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2020:FMb

[Ano20b] Anonymous. Front matter. Computer Graphics Forum, 39(2): i–iv, May 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2020:FMc

[Ano20c] Anonymous. Front matter. Computer Graphics Forum, 39(3): i–xi, June 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2020:FMd

[Ano20d] Anonymous. Front matter. Computer Graphics Forum, 39(3): i–vi, June 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2020:FMe

[Ano20e] Anonymous. Front matter. Computer Graphics Forum, 39(4): i–vi, July 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2020:FMf

[Ano20f] Anonymous. Front matter. Computer Graphics Forum, 39(5): i–xii, August 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2020:FMg

[Ano20g] Anonymous. Front matter. Computer Graphics Forum, 39(7):i– xiii, October 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2020:FMh

[Ano20h] Anonymous. Front matter. Computer Graphics Forum, 39(8):i– viii, December 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2020:IIa

[Ano20i] Anonymous. Issue information. Computer Graphics Forum, 39 (1):1–4, February 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2020:IIb

[Ano20j] Anonymous. Issue information. Computer Graphics Forum, 39 (6):1–4, September 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2020:P

[Ano20k] Anonymous. Paradigm. Computer Graphics Forum, 39(1):672– 673, February 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2021:ENU

[Ano21a] Anonymous. Erratum: Non-uniform subdivision surfaces with sharp features. Computer Graphics Forum, 40(1):424, February 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2021:FMa

[Ano21b] Anonymous. Front matter. Computer Graphics Forum, 40(2): i–xxvi, May 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2021:FMb

[Ano21c] Anonymous. Front matter. Computer Graphics Forum, 40(2): i–iv, May 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2021:FMc

[Ano21d] Anonymous. Front matter. Computer Graphics Forum, 40(3): i–xv, June 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2021:FMd

[Ano21e] Anonymous. Front matter. Computer Graphics Forum, 40(3): i–v, June 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2021:FMe

[Ano21f] Anonymous. Front matter. Computer Graphics Forum, 40(4): i–ix, July 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2021:FMf

[Ano21g] Anonymous. Front matter. Computer Graphics Forum, 40(5): i–xi, August 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2021:FMg

[Ano21h] Anonymous. Front matter. Computer Graphics Forum, 40(7):i– xiii, October 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2021:FMh

[Ano21i] Anonymous. Front matter. Computer Graphics Forum, 40(8):i– xii, December 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2021:IIa

[Ano21j] Anonymous. Issue information. Computer Graphics Forum, 40 (1):1–4, February 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2021:II

[Ano21k] Anonymous. Issue information. Computer Graphics Forum, 40 (6):1–4, September 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2022:BF

[Ano22a] Anonymous. Bubble flower. Computer Graphics Forum, 41(1): 507, February 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2022:EED

[Ano22b] Anonymous. Erratum: Evaluating data-type heterogeneity in interactive visual analyses with parallel axes. Computer Graphics Forum, 41(6):480–482, September 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic). See [ML22].

Anonymous:2022:FMa

[Ano22c] Anonymous. Front matter. Computer Graphics Forum, 41(2): i–xiv, May 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2022:FMb

[Ano22d] Anonymous. Front matter. Computer Graphics Forum, 41(2): i–v, May 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2022:FMc

[Ano22e] Anonymous. Front matter. Computer Graphics Forum, 41(3): i–xvii, June 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2022:FMd

[Ano22f] Anonymous. Front matter. Computer Graphics Forum, 41(3): i–v, June 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2022:FMe

[Ano22g] Anonymous. Front matter. Computer Graphics Forum, 41(4): i–xi, July 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2022:FMf

[Ano22h] Anonymous. Front matter. Computer Graphics Forum, 41(5): i–x, August 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2022:FMg

[Ano22i] Anonymous. Front matter. Computer Graphics Forum, 41(7):i– xvii, October 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2022:FMh

[Ano22j] Anonymous. Front matter. Computer Graphics Forum, 41(8):i– xiv, December 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2022:IIa

[Ano22k] Anonymous. Issue information. Computer Graphics Forum, 41 (1):1–4, February 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2022:II

[Ano22l] Anonymous. Issue information. Computer Graphics Forum, 41 (6):1–4, September 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2023:FMa

[Ano23a] Anonymous. Front matter. Computer Graphics Forum, 42(2): i–xx, May 2023. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2023:FMb

[Ano23b] Anonymous. Front matter. Computer Graphics Forum, 42(2): i–v, May 2023. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2023:II

[Ano23c] Anonymous. Issue information. Computer Graphics Forum, 42 (1):1–4, February 2023. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Anonymous:2023:PCN

[Ano23d] Anonymous. Proliferating cell nuclear antigen sliding along DNA. Computer Graphics Forum, 42(1):356, February 2023. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Alvarado:2022:GUB

[ARC22] Eduardo Alvarado, Damien Rohmer, and Marie-Paule Cani. Generating upper-body motion for real-time characters making their way through dynamic environments. Computer Graphics Forum, 41(8):169–181, December 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Arleo:2023:VEF

[ATL^{+}23] Alessio Arleo, Christos Tsigkanos, Roger A. Leite, Schahram Dustdar, Silvia Miksch, and Johannes Sorger. Visual exploration of financial data with incremental domain knowledge. Computer Graphics Forum, 42(1):101–116, February 2023. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Agarwal:2020:BVC

[AWB20] Shivam Agarwal, Günter Wallner, and Fabian Beck. Bombalytics: Visualization of competition and collaboration strategies of players in a Bomb laying game. Computer Graphics Forum, 39(3):89–100, June 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Atanasov:2021:MMM

[AWKK21] Asen Atanasov, Alexander Wilkie, Vladimir Koylazov, and Jaroslav Křivánek. A multiscale microfacet model based on inverse bin mapping. Computer Graphics Forum, 40(2):103–113, May 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Aydinlilar:2021:FRT

[AZ21] Melike Aydinlilar and Cedric Zanni. Fast ray tracing of scaleinvariant integral surfaces. Computer Graphics Forum, 40(6): 117–134, September 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bettio:2021:NAE

[BAMG21] F. Bettio, M. Ahsan, F. Marton, and E. Gobbetti. A novel approach for exploring annotated data with interactive lenses.

Computer Graphics Forum, 40(3):387–398, June 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Basselin:2021:RPD

[BAR^{+}21] J. Basselin, L. Alonso, N. Ray, D. Sokolov, S. Lefebvre, and B. Lévy. Restricted power diagrams on the GPU. Computer Graphics Forum, 40(2):1–12, May 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bunge:2023:SDL

[BB23] A. Bunge and M. Botsch. A survey on discrete Laplacians for general polygonal meshes. Computer Graphics Forum, 42(2): 521–544, May 2023. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bunge:2021:I

[BBA21] A. Bunge, M. Botsch, and M. Alexa. Image. Computer Graphics Forum, 40(5):217–230, August 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bukenberger:2022:CVD

[BBB22] D. R. Bukenberger, K. Buchin, and M. Botsch. Constructing L_{∞} Voronoi diagrams in 2D and 3D. Computer Graphics Forum, 41 (5):135–147, August 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bludau:2020:RTS

[BBBD20] M.-J. Bludau, V. Brüggemann, A. Busch, and M. Dörk. Reading traces: Scalable exploration in elastic visualizations of cultural heritage data. Computer Graphics Forum, 39(3):77–87, June 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bast:2020:MMO

[BBS20] Hannah Bast, Patrick Brosi, and Sabine Storandt. Metro maps on octilinear grid graphs. Computer Graphics Forum, 39(3):357– 367, June 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bhattacharjee:2020:SSB

[BC20] Sukanya Bhattacharjee and Parag Chaudhuri. A survey on sketch based content creation: from the desktop to virtual and augmented reality. Computer Graphics Forum, 39(2):757–780, May 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bhattacharjee:2020:PPD

[BCD20] Kaustav Bhattacharjee, Min Chen, and Aritra Dasgupta. Privacy-preserving data visualization: Reflections on the state of the art and research opportunities. Computer Graphics Forum, 39(3):675–692, June 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bedel:2022:CSF

[BCOM^{+}22] A. Bedel, Y. Coudert-Osmont, J. Martínez, R. I. Nishat, S. Whitesides, and S. Lefebvre. Closed space-filling curves with controlled orientation for 3D printing. Computer Graphics Forum, 41(2):473–492, May 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bitterli:2022:PFP

[Bd22] Benedikt Bitterli and Eugene d’Eon. A position-free path integral for homogeneous slabs and multiple scattering on Smith microfacets. Computer Graphics Forum, 41(4):93–104, July 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bonneel:2023:SOT

[BD23] Nicolas Bonneel and Julie Digne. A survey of optimal transport for computer graphics and computer vision. Computer Graphics Forum, 42(2):439–460, May 2023. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bekos:2022:CSL

[BDF^{+}22] M. A. Bekos, D. J. C. Dekker, F. Frank, W. Meulemans, P. Rodgers, A. Schulz, and S. Wessel. Computing schematic layouts for spatial hypergraphs on concentric circles and grids. Computer Graphics Forum, 41(6):316–335, September 2022. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Binucci:2022:PAD

[BDK^{+}22] Carla Binucci, Walter Didimo, Michael Kaufmann, Giuseppe Liotta, and Fabrizio Montecchiani. Placing arrows in directed graph layouts: Algorithms and experiments. Computer Graphics Forum, 41(1):364–376, February 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Blumenschein:2020:VPD

[BDL^{+}20] Michael Blumenschein, Luka J. Debbeler, Nadine C. Lages, Britta Renner, Daniel A. Keim, and Mennatallah El-Assady. vplots: Designing hybrid charts for the comparative analysis of data distributions. Computer Graphics Forum, 39(3):565–577, June 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bermano:2022:SAA

[BGA^{+}22] A. H. Bermano, R. Gal, Y. Alaluf, R. Mokady, Y. Nitzan, O. Tov, O. Patashnik, and D. Cohen-Or. State-of-the-art in the architecture, methods and applications of StyleGAN. Computer Graphics Forum, 41(2):591–611, May 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bartolovic:2020:PSP

[BGG20] Nemanja Bartolovic, Markus Gross, and Tobias Günther. Phase space projection of dynamical systems. Computer Graphics Forum, 39(3):253–264, June 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bruel-Gabrielsson:2020:TAS

[BGGSSG20] Rickard Brüel-Gabrielsson, Vignesh Ganapathi-Subramanian, Primoz Skraba, and Leonidas J. Guibas. Topology-aware surface reconstruction for point clouds. Computer Graphics Forum, 39(5):197–207, August 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bruckler:2022:MCS

[BGMC22] Hendrik Brückler, Ojaswi Gupta, Manish Mandad, and Marcel Campen. The 3D motorcycle complex for structured volume decomposition. Computer Graphics Forum, 41(2):221–235, May 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Baatz:2022:NTN

[BGP^{+}22] H. Baatz, J. Granskog, M. Papas, F. Rousselle, and J. Novák. NeRF-Tex: Neural reflectance field textures. Computer Graphics Forum, 41(6):287–301, September 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Band:2020:CNL

[BGT20] Stefan Band, Christoph Gissler, and Matthias Teschner. Compressed neighbour lists for SPH. Computer Graphics Forum, 39 (1):531–542, February 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Barthel:2023:IEG

[BHJS23] K. U. Barthel, N. Hezel, K. Jung, and K. Schall. Improved evaluation and generation of grid layouts using distance preservation quality and linear assignment sorting. Computer Graphics Forum, 42(1):261–276, February 2023. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bunge:2020:PLM

[BHKB20] Astrid Bunge, Philipp Herholz, Misha Kazhdan, and Mario Botsch. Polygon Laplacian made simple. Computer Graphics Forum, 39(2):303–313, May 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bonneau:2021:GCA

[BHM21] Gp. Bonneau, S. Hahmann, and J. Marku. Geometric construction of auxetic metamaterials. Computer Graphics Forum, 40(2): 291–304, May 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bethe:2020:PSS

[BJG20] F. Bethe, J. Jendersie, and T. Grosch. Preserving shadow silhouettes in illumination-driven mesh reduction. Computer Graphics Forum, 39(6):155–166, September 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Baumer:2022:CPC

[BJSM22] Eric P. S. Baumer, Mahmood Jasim, Ali Sarvghad, and Narges Mahyar. Of course it’s political! A critical inquiry into underemphasized dimensions in civic text visualization. Computer Graphics Forum, 41(3):1–14, June 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bang:2021:EGP

[BKL21] Seungbae Bang, Maria Korosteleva, and Sung-Hee Lee. Estimating garment patterns from static scan data. Computer Graphics Forum, 40(6):273–287, September 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Burg:2020:RTA

[BLC20] L. Burg, C. Lino, and M. Christie. Real-time anticipation of occlusions for automated camera control in toric space. Computer Graphics Forum, 39(2):523–533, May 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Benaim:2021:SAS

[BMBW21] S. Benaim, R. Mokady, A. Bermano, and L. Wolf. Structural analogy from a single image pair. Computer Graphics Forum, 40 (1):249–265, February 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bertolino:2021:EHO

[BMPP21] Giulia Bertolino, Marco Montemurro, Nicolas Perry, and Franck Pourroy. An efficient hybrid optimization strategy for surface reconstruction. Computer Graphics Forum, 40(6):215–241, September 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bauerle:2020:CGV

[BNR20] A. Bäuerle, H. Neumann, and T. Ropinski. Classifier-guided visual correction of noisy labels for image classification tasks. Computer Graphics Forum, 39(3):195–205, June 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bieron:2020:ABF

[BP20] J. Bieron and P. Peers. An adaptive BRDF fitting metric. Computer Graphics Forum, 39(4):59–74, July 2020. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Benamira:2021:CSD

[BP21] Alexis Benamira and Sumanta Pattanaik. A combined scattering and diffraction model for elliptical hair rendering. Computer Graphics Forum, 40(4):163–175, July 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bajo:2020:RBM

[BPD20] J. M. Bajo, G. Patow, and C. A. Delrieux. Realistic buoyancy model for real-time applications. Computer Graphics Forum, 39 (6):217–231, September 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Beaufort:2022:HMI

[BRK^{+}22] P.-A. Beaufort, M. Reberol, D. Kalmykov, H. Liu, F. Ledoux, and D. Bommes. Hex me if you can. Computer Graphics Forum, 41 (5):125–134, August 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bailey:2022:MFP

[BRMO22] Stephen W. Bailey, Jérémy Riviere, Morten Mikkelsen, and James F. O’Brien. Monocular facial performance capture via deep expression matching. Computer Graphics Forum, 41(8): 243–254, December 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Blecha:2020:FSM

[BRP^{+}20] C. Blecha, F. Raith, A. J. Präger, T. Nagel, O. Kolditz, J. Maßmann, N. Röber, M. Böttinger, and G. Scheuermann. Fiber surfaces for many variables. Computer Graphics Forum, 39(3):317– 329, June 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bai:2022:SAE

[BS22] Song Bai and Craig Schroeder. Stability analysis of explicit MPM. Computer Graphics Forum, 41(8):19–30, December 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Besancon:2020:RAR

[BSB^{+}20] Lonni Besançon, Amir Semmo, David Biau, Bruno Frachet, Virginie Pineau, El Hadi Sariali, Marc Soubeyrand, Rabah Taouachi, Tobias Isenberg, and Pierre Dragicevic. Reducing affective responses to surgical images and videos through stylization. Computer Graphics Forum, 39(1):462–483, February 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Born:2021:SMH

[BSCK21] J. Born, P. Schmidt, M. Campen, and L. Kobbelt. Surface map homology inference. Computer Graphics Forum, 40(5):193–204, August 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Binninger:2022:SIC

[BSH22] Alexandre Binninger and Olga Sorkine-Hornung. Smooth interpolating curves with local control and monotone alternating curvature. Computer Graphics Forum, 41(5):25–38, August 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Born:2021:LEC

[BSK21] J. Born, P. Schmidt, and L. Kobbelt. Layout embedding via combinatorial optimization. Computer Graphics Forum, 40(2): 277–290, May 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Brich:2022:VAM

[BSP^{+}22] N. Brich, C. Schulz, J. Peter, W. Klingert, M. Schenk, D. Weiskopf, and M. Krone. Visual analytics of multivariate intensive care time series data. Computer Graphics Forum, 41(6): 273–286, September 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Berson:2020:IFA

[BSS20] Eloïse Berson, Catherine Soladié, and Nicolas Stoiber. Intuitive facial animation editing based on a generative RNN framework. Computer Graphics Forum, 39(8):241–251, December 2020. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bisht:2022:TAG

[BSU^{+}22] Sumit Bisht, Krishnendra Shekhawat, Nitant Upasani, Rahil N. Jain, Riddhesh Jayesh Tiwaskar, and Chinmay Hebbar. Transforming an adjacency graph into dimensioned floorplan layouts. Computer Graphics Forum, 41(6):5–22, September 2022. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Beyer:2022:SVA

[BTB^{+}22] Johanna Beyer, Jakob Troidl, Saeed Boorboor, Markus Hadwiger, Arie Kaufman, and Hanspeter Pfister. A survey of visualization and analysis in high-resolution connectomics. Computer Graphics Forum, 41(3):573–607, June 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

BenCharrada:2022:TTL

[BTCL22] Tarek Ben Charrada, Hedi Tabia, Aladine Chetouani, and Hamid Laga. TopoNet: Topology learning for 3D reconstruction of objects of arbitrary genus. Computer Graphics Forum, 41(6): 336–347, September 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bukenberger:2022:MHM

[BTL22] Dennis R. Bukenberger, Marco Tarini, and Hendrik P. A. Lensch. At-most-hexa meshes. Computer Graphics Forum, 41(1):7–28, February 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Batik:2022:SGM

[BTW^{+}22] T. Batik, S. Terziadis, Y.-S. Wang, M. Nöllenburg, and H.-Y. Wu. Shape-guided mixed metro map layout. Computer Graphics Forum, 41(7):495–506, October 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Binninger:2021:DAG

[BVHSH21] Alexandre Binninger, Floor Verhoeven, Philipp Herholz, and Olga Sorkine-Hornung. Developable approximation via Gauss image thinning. Computer Graphics Forum, 40(5):289–300, August 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bauerle:2022:WDM

[BvOdK^{+}22] A. Bäuerle, C. van Onzenoodt, S. der Kinderen, J. Johansson Westberg, D. Jönsson, and T. Ropinski. Where did my lines go? Visualizing missing data in parallel coordinates. Computer Graphics Forum, 41(3):235–246, June 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Bujack:2020:SAT

[BYH^{+}20] Roxana Bujack, Lin Yan, Ingrid Hotz, Christoph Garth, and Bei Wang. State of the art in time-dependent flow topology: Interpreting physical meaningfulness through mathematical properties. Computer Graphics Forum, 39(3):811–835, June 2020. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Besancon:2021:BRS

[BYK^{+}21] Lonni Besançon, Anders Ynnerman, Daniel F. Keefe, Lingyun Yu, and Tobias Isenberg. Book review: The State of the Art of Spatial Interfaces for 3D Visualization. Computer Graphics Forum, 40(1):293–326, February 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Blumenschein:2020:ERS

[BZP^{+}20] Michael Blumenschein, Xuan Zhang, David Pomerenke, Daniel A. Keim, and Johannes Fuchs. Evaluating reordering strategies for cluster identification in parallel coordinates. Computer Graphics Forum, 39(3):537–549, June 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Coevoet:2020:DCR

[CARK20] E. Coevoet, S. Andrews, D. Relles, and P. G. Kry. Distant collision response in rigid body simulations. Computer Graphics Forum, 39(8):113–122, December 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cavallo:2021:HDG

[Cav21] M. Cavallo. Higher dimensional graphics: Conceiving worlds in four spatial dimensions and beyond. Computer Graphics Forum, 40(2):51–63, May 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cohen:2020:RSC

[CBC20] Aharon Cohen and Mirela Ben-Chen. Robust shape collection matching and correspondence from shape differences. Computer Graphics Forum, 39(2):555–568, May 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Careil:2020:IMC

[CBE20] V. Careil, M. Billeter, and E. Eisemann. Interactively modifying compressed sparse voxel representations. Computer Graphics Forum, 39(2):111–119, May 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chao:2020:SVT

[CBL^{+}20] Qianwen Chao, Huikun Bi, Weizi Li, Tianlu Mao, Zhaoqi Wang, Ming C. Lin, and Zhigang Deng. A survey on visual traffic simulation: Models, evaluations, and applications in autonomous driving. Computer Graphics Forum, 39(1):287–308, February 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cogalan:2023:VFI

[ÇBSM23] U. Çoğalan, M. Bemana, Hp. Seidel, and K. Myszkowski. Video frame interpolation for high dynamic range sequences captured with dual-exposure sensors. Computer Graphics Forum, 42(2): 119–131, May 2023. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Clausen:2023:ISD

[CCFM23] O. Clausen, Y. Chen, A. Fuhrmann, and R. Marroquim. Investigation and simulation of diffraction on rough surfaces. Computer Graphics Forum, 42(1):245–260, February 2023. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chen:2022:UUA

[CCG^{+}22] Zhilei Chen, Honghua Chen, Lina Gong, Xuefeng Yan, Jun Wang, Yanwen Guo, Jing Qin, and Mingqiang Wei. UTOPIC: Uncertainty-aware overlap prediction network for partial point cloud registration. Computer Graphics Forum, 41(7):87–98, October 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chen:2021:TNG

[CCOCM21] Xuelin Chen, Daniel Cohen-Or, Baoquan Chen, and Niloy J. Mitra. Towards a neural graphics pipeline for controllable image generation. Computer Graphics Forum, 40(2):127–140, May 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Casafranca:2020:MYT

[CCR^{+}20] Juan J. Casafranca, Gabriel Cirio, Alejandro Rodríguez, Eder Miguel, and Miguel A. Otaduy. Mixing yarns and triangles in cloth simulation. Computer Graphics Forum, 39(2):101–110, May 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Currius:2020:SGL

[CDAS20] R. R. Currius, D. Dolonius, U. Assarsson, and E. Sintorn. Spherical Gaussian light-field textures for fast precomputed global illumination. Computer Graphics Forum, 39(2):133–146, May 2020.

CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chitalu:2020:BOI

[CDK20] Floyd M. Chitalu, Christophe Dubach, and Taku Komura. Binary ostensibly-implicit trees for fast collision detection. Computer Graphics Forum, 39(2):509–521, May 2020. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cabrero-Daniel:2022:DCC

[CDMH^{+}22] B. Cabrero-Daniel, R. Marques, L. Hoyet, J. Pettré, and J. Blat. Dynamic combination of crowd steering policies based on context. Computer Graphics Forum, 41(2):209–219, May 2022. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Casado-Elvira:2022:PPG

[CETC22] Andrés Casado-Elvira, Marc Comino Trinidad, and Dan Casas. PERGAMO: Personalized 3D garments from monocular video. Computer Graphics Forum, 41(8):293–304, December 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ciklabakkal:2022:SPS

[CGG^{+}22] Ege Ciklabakkal, Adrien Gruson, Iliyan Georgiev, Derek Nowrouzezahrai, and Toshiya Hachisuka. Single-pass stratified importance resampling. Computer Graphics Forum, 41(4):41–49, July 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Choi:2023:OAM

[CHC^{+}23] Soojin Choi, Seokpyo Hong, Kyungmin Cho, Chaelin Kim, and Junyong Noh. Online avatar motion adaptation to morphologically-similar spaces. Computer Graphics Forum, 42 (2):13–24, May 2023. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chentanez:2022:LPH

[CJMM22] Nuttapong Chentanez, Stefan Jeschke, Matthias Müller, and Miles Macklin. Learning physics with a hierarchical graph network. Computer Graphics Forum, 41(8):283–292, December 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Csoba:2021:ERO

[CK21] István Csoba and Roland Kunkli. Efficient rendering of ocular wavefront aberrations using tiled point-spread function splatting. Computer Graphics Forum, 40(6):182–199, September 2021. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chen:2021:MGI

[CL21] Minchan Chen and Manfred Lau. A motion-guided interface for modeling 3D multi-functional furniture. Computer Graphics Forum, 40(7):229–240, October 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chao:2022:MNN

[CL22a] X. J. Chao and H. Leung. MP-NeRF: Neural radiance fields for dynamic multi-person synthesis from sparse views. Computer Graphics Forum, 41(8):317–325, December 2022. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chen:2022:LSA

[CL22b] Minchan Chen and Manfred Lau. Learning 3D shape aesthetics globally and locally. Computer Graphics Forum, 41(7):579–588, October 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chen:2022:ECR

[CLC^{+}22] Lianggangxu Chen, Jiale Lu, Yiqing Cai, Changbo Wang, and Gaoqi He. Exploring contextual relationships in 3D cloud points by semantic knowledge mining. Computer Graphics Forum, 41 (7):75–86, October 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chen:2022:MFN

[CLCO^{+}22] Xuelin Chen, Weiyu Li, Daniel Cohen-Or, Niloy J. Mitra, and Baoquan Chen. MoCo-Flow: Neural motion consensus flow for dynamic humans in stationary monocular cameras. Computer Graphics Forum, 41(2):147–161, May 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chang:2020:SAS

[CLH^{+}20] Yue Chang, Shusen Liu, Xiaowei He, Sheng Li, and Guoping Wang. Semi-analytical solid boundary conditions for free surface flows. Computer Graphics Forum, 39(7):131–141, October 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cai:2020:IIG

[CLL^{+}20] Zhong-Qi Cai, Ying-Sheng Luo, Yu-Chi Lai, Chih-Shiang Chan, and Wen-Kai Tai. Interactive iconized grammar-based pailou modelling. Computer Graphics Forum, 39(1):50–62, February 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chen:2022:DVD

[CLL^{+}22] S. Chen, S. Li, Y. Li, J. Zhu, J. Long, S. Chen, J. Zhang, and X. Yuan. DanmuVis: Visualizing danmu content dynamics and associated viewer behaviors in online videos. Computer Graphics Forum, 41(3):429–440, June 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chan-Lock:2022:HOE

[CLPO22] Antoine Chan-Lock, Jesús Pérez, and Miguel A. Otaduy. Highorder elasticity interpolants for microstructure simulation. Computer Graphics Forum, 41(8):63–74, December 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cao:2020:FPD

[CLR^{+}20] Wei Cao, Luan Lyu, Xiaohua Ren, Bob Zhang, Zhixin Yang, and Enhua Wu. Fracture patterns design for anisotropic models with the material point method. Computer Graphics Forum, 39 (7):93–104, October 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cui:2020:FRS

[CLSK20] Qiaodong Cui, Timothy Langlois, Pradeep Sen, and Theodore Kim. Fast and robust stochastic structural optimization. Computer Graphics Forum, 39(2):385–397, May 2020. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Coelho:2020:IED

[CM20] D. Coelho and K. Mueller. Infomages: Embedding data into thematic images. Computer Graphics Forum, 39(3):593–606, June 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Choi:2021:CPR

[CM21a] Hajin Choi and Bochang Moon. Consistent post-reconstruction for progressive photon mapping. Computer Graphics Forum, 40 (7):121–130, October 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Coenen:2021:PDV

[CM21b] J. Coenen and A. Vande Moere. Public data visualization: Analyzing local running statistics on situated displays. Computer Graphics Forum, 40(3):159–171, June 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cao:2021:LAC

[CMG21] Ruizhi Cao, Haoran Mo, and Chengying Gao. Line art colorization based on explicit region segmentation. Computer Graphics Forum, 40(7):1–10, October 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chatzimparmpas:2020:SAE

[CMJ^{+}20] A. Chatzimparmpas, R. M. Martins, I. Jusufi, K. Kucher, F. Rossi, and A. Kerren. The state of the art in enhancing trust in machine learning models with the use of visualizations. Computer Graphics Forum, 39(3):713–756, June 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chatzimparmpas:2021:VVA

[CMKK21] A. Chatzimparmpas, R. M. Martins, K. Kucher, and A. Kerren. VisEvol: Visual analytics to support hyperparameter search through evolutionary optimization. Computer Graphics Forum, 40(3):201–214, June 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chentanez:2020:CSD

[CMM^{+}20] Nuttapong Chentanez, Miles Macklin, Matthias Müller, Stefan Jeschke, and Tae-Yong Kim. Cloth and skin deformation with a triangle mesh based convolutional neural network. Computer Graphics Forum, 39(8):123–134, December 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cibulski:2020:PPF

[CMMK20] Lena Cibulski, Hubert Mitterhofer, Thorsten May, and Jörn Kohlhammer. PAVED: Pareto front visualization for engineering design. Computer Graphics Forum, 39(3):405–416, June 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chitalu:2020:DCX

[CMSK20] Floyd M. Chitalu, Qinghai Miao, Kartic Subr, and Taku Komura. Displacement-correlated XFEM for simulating brittle fracture. Computer Graphics Forum, 39(2):569–583, May 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cohen:2022:SSA

[CNS22] N. Cohen, Y. Newman, and A. Shamir. Semantic segmentation in art paintings. Computer Graphics Forum, 41(2):261–275, May 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Casafranca:2022:VFS

[CO22] Juan J. Casafranca and Miguel A. Otaduy. Voronoi filters for simulation enrichment. Computer Graphics Forum, 41(8):43–51, December 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Canabal:2020:SDP

[COKE20] José A. Canabal, Miguel A. Otaduy, Byungmoon Kim, and Jose Echevarria. Simulation of dendritic painting. Computer Graphics Forum, 39(2):597–606, May 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chatzimparmpas:2023:HVA

[CPK23] A. Chatzimparmpas, F. V. Paulovich, and A. Kerren. HardVis: Visual analytics to handle instance hardness using undersampling and oversampling techniques. Computer Graphics Forum, 42(1): 135–154, February 2023. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Conde-Rodriguez:2021:MMM

[CRGFT21] F. Conde-Rodríguez, Á-.L. García-Fernández, and J. C. Torres. Modelling material microstructure using the Perlin noise function. Computer Graphics Forum, 40(1):195–208, February 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Courtin:2021:SRS

[CRH^{+}21] Simon Courtin, Mickael Ribardière, Sebastien Horna, Pierre Poulin, and Daniel Meneveaux. SREC-RT: a structure for ray tracing rounded edges and corners. Computer Graphics Forum, 40(6):200–214, September 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chaudhuri:2020:BRL

[CRW^{+}20] Siddhartha Chaudhuri, Daniel Ritchie, Jiajun Wu, Kai Xu, and Hao Zhang. Book review: Learning Generative Models of 3D Structures. Computer Graphics Forum, 39(2):643–666, May 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chen:2022:EAP

[CRZ^{+}22] Q. Chen, Y. Ren, Z. Zhao, W. Tao, and H. Zhao. Error analysis of photometric stereo with near quasi-point lights. Computer Graphics Forum, 41(6):149–165, September 2022. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cha:2023:GTH

[CSAN23] Sihun Cha, Kwanggyoon Seo, Amirsaman Ashtari, and Junyong Noh. Generating texture for 3D human avatar from a single image using sampling and refinement networks. Computer Graphics Forum, 42(2):385–396, May 2023. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cakmak:2020:MVA

[CSB^{+}20] E. Cakmak, H. Schäfer, J. Buchmüller, J. Fuchs, T. Schreck, A. Jordan, and D. Keim. MotionGlyphs: Visual abstraction of spatio-temporal networks in collective animal behavior. Computer Graphics Forum, 39(3):63–75, June 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chermain:2020:PPB

[CSDD20] X. Chermain, B. Sauvage, J.-M. Dischler, and C. Dachsbacher. Procedural physically based BRDF for real-time rendering of glints. Computer Graphics Forum, 39(7):243–253, October 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Csebfalvi:2023:OSF

[Csé23] Balázs Csébfalvi. One step further beyond trilinear interpolation and central differences: Triquadratic reconstruction and its analytic derivatives at the cost of one additional texture fetch. Computer Graphics Forum, 42(2):191–200, May 2023. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chao:2021:PBA

[CSG21] Cheng-Kang Ted Chao, Karan Singh, and Yotam Gingold. PosterChild: Blend-aware artistic posterization. Computer Graphics Forum, 40(4):87–99, July 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chen:2023:WDT

[CSK^{+}23] Eric Ming Chen, Jin Sun, Apoorv Khandelwal, Dani Lischinski, Noah Snavely, and Hadar Averbuch-Elor. What’s in a decade? Transforming faces through time. Computer Graphics Forum, 42(2):281–291, May 2023. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cascaval:2022:DCP

[CSQ^{+}22] D. Cascaval, M. Shalah, P. Quinn, R. Bodik, M. Agrawala, and A. Schulz. Differentiable 3D CAD programs for bidirectional editing. Computer Graphics Forum, 41(2):309–323, May 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chong:2021:IOG

[CSSI21] Toby Chong, I-Chao Shen, Issei Sato, and Takeo Igarashi. Interactive optimization of generative image modelling using sequential subspace search and content-based guidance. Computer Graphics Forum, 40(1):279–292, February 2021. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chen:2020:CAM

[CTJ^{+}20] L. Chen, W. Tang, N. W. John, T. R. Wan, and J. J. Zhang. Context-aware mixed reality: A learning-based framework for semantic-level interaction. Computer Graphics Forum, 39(1): 484–496, February 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Corda:2020:RTD

[CTL^{+}20] F. Corda, J. M. Thiery, M. Livesu, E. Puppo, T. Boubekeur, and R. Scateni. Real-time deformation with coupled cages and skeletons. Computer Graphics Forum, 39(6):19–32, September 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Colas:2022:IFI

[CvTZ^{+}22] A. Colas, W. van Toll, K. Zibrek, L. Hoyet, A.-H. Olivier, and J. Pettré. Interaction fields: Intuitive sketch-based steering behaviors for crowd simulation. Computer Graphics Forum, 41(2): 521–534, May 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chu:2022:CEC

[CW22] Yiyao Chu and Wencheng Wang. Out-of-core extraction of curve skeletons for large volumetric models. Computer Graphics Forum, 41(7):1–12, October 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cheng:2022:IRI

[CWW^{+}22] Hao Cheng, Junhong Wang, Yun Wang, Bongshin Lee, Haidong Zhang, and Dongmei Zhang. Investigating the role and interplay of narrations and animations in data videos. Computer Graphics Forum, 41(3):527–539, June 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chen:2021:UUO

[CXW^{+}21] Luanmin Chen, Juzhan Xu, Chuan Wang, Haibin Huang, Hui Huang, and Ruizhen Hu. UprightRL: Upright orientation estimation of 3D shapes via reinforcement learning. Computer Graphics Forum, 40(7):265–275, October 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cheng:2022:QFI

[CXWZ22] Haojie Cheng, Chunxiao Xu, Jiajun Wang, and Lingxiao Zhao. Quad-fisheye image stitching for monoscopic panorama reconstruction. Computer Graphics Forum, 41(6):94–109, September 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Cai:2022:PBI

[CYD^{+}22] G. Cai, K. Yan, Z. Dong, I. Gkioulekas, and S. Zhao. Physicsbased inverse rendering using combined implicit and explicit geometries. Computer Graphics Forum, 41(4):129–138, July 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chandran:2022:FAD

[CZG^{+}22a] Prashanth Chandran, Gaspard Zoss, Markus Gross, Paulo Gotardo, and Derek Bradley. Facial animation with disentangled identity and motion using transformers. Computer Graphics Forum, 41(8):267–277, December 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Chandran:2022:STT

[CZG^{+}22b] Prashanth Chandran, Gaspard Zoss, Markus Gross, Paulo Gotardo, and Derek Bradley. Shape transformers: Topologyindependent 3D shape models using transformers. Computer Graphics Forum, 41(2):195–207, May 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Carrigan:2020:EPF

[CZGM20] E. Carrigan, E. Zell, C. Guiard, and R. McDonnell. Expression packing: As-few-as-possible training expressions for blendshape transfer. Computer Graphics Forum, 39(2):219–233, May 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dib:2021:PFR

[DBA^{+}21] A. Dib, G. Bharaj, J. Ahn, C. Thébault, P. Gosselin, M. Romeo, and L. Chevallier. Practical face reconstruction via differentiable ray tracing. Computer Graphics Forum, 40(2):153–164, May 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Duan:2020:DSD

[DBP20] Z. Duan, J. Bieron, and P. Peers. Deep separation of direct and global components from a single photograph under structured lighting. Computer Graphics Forum, 39(7):459–470, October 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Donati:2022:CFM

[DCMO22] Nicolas Donati, Etienne Corman, Simone Melzi, and Maks Ovsjanikov. Complex functional maps: a conformal link between tangent bundles. Computer Graphics Forum, 41(1):317–334, February 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Deschaintre:2020:GFT

[DDB20] Valentin Deschaintre, George Drettakis, and Adrien Bousseau. Guided fine-tuning for large-scale material transfer. Computer Graphics Forum, 39(4):91–105, July 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

dEon:2021:ABM

[d’E21] Eugene d’Eon. An analytic BRDF for materials with spherical Lambertian scatterers. Computer Graphics Forum, 40(4):153– 161, July 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dennig:2021:PQM

[DFB^{+}21] Frederik L. Dennig, Maximilian T. Fischer, Michael Blumenschein, Johannes Fuchs, Daniel A. Keim, and Evanthia Dimara. ParSetgnostics: Quality metrics for parallel sets. Computer Graphics Forum, 40(3):375–386, June 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Diaz:2021:TNV

[DFV21] J. Díaz, M. Fort, and P. Vázquez. TourVis: Narrative visualization of multi-stage bicycle races. Computer Graphics Forum, 40(3):531–542, June 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Diolatzis:2020:PPP

[DGJ^{+}20] Stavros Diolatzis, Adrien Gruson, Wenzel Jakob, Derek Nowrouzezahrai, and George Drettakis. Practical product path guiding using linearly transformed cosines. Computer Graphics Forum, 39(4):23–33, July 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dvorak:2023:PBE

[DKVV23] Jan Dvořák, Zuzana Káčereková, Petr Vaněček, and Libor Váša. Priority-based encoding of triangle mesh connectivity for a known geometry. Computer Graphics Forum, 42(1):60–71, February 2023. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Delanoy:2022:GFI

[DLC^{+}22] J. Delanoy, M. Lagunas, J. Condor, D. Gutierrez, and B. Masia. A generative framework for image-based editing of material appearance using perceptual attributes. Computer Graphics Forum, 41(1):453–464, February 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dielen:2021:LDF

[DLLK21] Alexander Dielen, Isaak Lim, Max Lyon, and Leif Kobbelt. Learning direction fields for quad mesh generation. Computer Graphics Forum, 40(5):181–191, August 2021. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dai:2021:TLW

[DLS21] Yutong Dai, Hao Lu, and Chunhua Shen. Towards light-weight portrait matting via parameter sharing. Computer Graphics Forum, 40(1):151–164, February 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dou:2022:CAI

[DLX^{+}22] Zhiyang Dou, Cheng Lin, Rui Xu, Lei Yang, Shiqing Xin, Taku Komura, and Wenping Wang. Coverage axis: Inner point selection for 3D shape skeletonization. Computer Graphics Forum, 41(2):419–432, May 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Doi:2021:GIA

[DMT21] K. Doi, Y. Morimoto, and R. Tsuruno. Global illuminationaware stylised shading. Computer Graphics Forum, 40(7):11–20, October 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

DiBartolomeo:2022:SMT

[DPB^{+}22a] Sara Di Bartolomeo, Alexis Pister, Paolo Buono, Catherine Plaisant, Cody Dunne, and Jean-Daniel Fekete. Six methods for transforming layered hypergraphs to apply layered graph layout algorithms. Computer Graphics Forum, 41(3):259–270, June 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dorda:2022:DSO

[DPB^{+}22b] D. Dorda, D. Peter, D. Borer, N. B. Huber, I. Sailer, M. Gross, B. Solenthaler, and B. Thomaszewski. Differentiable simulation for outcome-driven orthognathic surgery planning. Computer Graphics Forum, 41(8):53–61, December 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dumery:2022:EGL

[DPM^{+}22] C. Dumery, F. Protais, S. Mestrallet, C. Bourcier, and F. Ledoux. Evocube: a genetic labelling framework for polycube-maps. Computer Graphics Forum, 41(6):467–479, September 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dodik:2022:PGU

[DPÖM22] Ana Dodik, Marios Papas, Cengiz Öztireli, and Thomas Müller. Path guiding using spatio-directional mixture models. Computer Graphics Forum, 41(1):172–189, February 2022. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Diehl:2021:HTC

[DPR^{+}21] Alexandra Diehl, Rodrigo Pelorosso, Juan Ruiz, Renato Pajarola, M. Eduard Gröller, and Stefan Bruckner. Hornero: Thunderstorms characterization using visual analytics. Computer Graphics Forum, 40(3):299–310, June 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Neto:2020:VAA

[dQNSVE20] José F. de Queiroz Neto, Emanuele Santos, Creto Augusto Vidal, and David S. Ebert. A visual analytics approach to facilitate crime hotspot analysis. Computer Graphics Forum, 39(3):139– 151, June 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dongbo:2022:HSI

[DR22] Shi Dongbo and Chen Renjie. Harmonic shape interpolation on multiply-connected planar domains. Computer Graphics Forum, 41(5):1–11, August 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dolonius:2020:UFT

[DSA20] D. Dolonius, E. Sintorn, and U. Assarsson. UV-free texturing using sparse voxel DAGs. Computer Graphics Forum, 39(2):121– 132, May 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Djavaherpour:2021:DPS

[DSMA^{+}21] H. Djavaherpour, F. Samavati, A. Mahdavi-Amiri, F. Yazdanbakhsh, S. Huron, R. Levy, Y. Jansen, and L. Oehlberg. Data to physicalization: a survey of the physical rendering process. Computer Graphics Forum, 40(3):569–598, June 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dupuy:2021:HRR

[DV21] J. Dupuy and K. Vanhoey. A halfedge refinement rule for parallel Catmull–Clark subdivision. Computer Graphics Forum, 40(8): 57–70, December 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Desrichard:2022:SLP

[DVP22] François Desrichard, David Vanderhaeghe, and Mathias Paulin. Shadow layers for participating media. Computer Graphics Forum, 41(1):190–200, February 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dhanoa:2022:PMD

[DWH^{+}22] Vaishali Dhanoa, Conny Walchshofer, Andreas Hinterreiter, Holger Stitz, Eduard Groeller, and Marc Streit. A process model for dashboard onboarding. Computer Graphics Forum, 41(3):501– 513, June 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Dai:2022:DFR

[DX22] Pinxuan Dai and Ning Xie. Deep flow rendering: View synthesis via layer-aware reflection flow. Computer Graphics Forum, 41(4): 139–148, July 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Das:2020:QIC

[DXG^{+}20] Subhajit Das, Shenyu Xu, Michael Gleicher, Remco Chang, and Alex Endert. QUESTO: Interactive construction of objective functions for classification tasks. Computer Graphics Forum, 39 (3):153–165, June 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Deng:2022:SNR

[DYDZ22] Bailin Deng, Yuxin Yao, Roberto M. Dyke, and Juyong Zhang. A survey of non-rigid 3D registration. Computer Graphics Forum, 41(2):559–589, May 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Du:2021:LPG

[DZN^{+}21] Dong Du, Heming Zhu, Yinyu Nie, Xiaoguang Han, Shuguang Cui, Yizhou Yu, and Ligang Liu. Learning part generation and assembly for sketching man-made objects. Computer Graphics Forum, 40(1):222–233, February 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Even:2023:NLR

[EBB23] Melvin Even, Pierre Bénard, and Pascal Barla. Non-linear rough 2D animation using transient embeddings. Computer Graphics Forum, 42(2):411–425, May 2023. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ehmel:2021:TVC

[EBD21] F. Ehmel, V. Brüggemann, and M. Dörk. Topography of violence: Considerations for ethical and collaborative visualization design. Computer Graphics Forum, 40(3):13–24, June 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Elavsky:2022:HAM

[EBM22] Frank Elavsky, Cynthia Bennett, and Dominik Moritz. How accessible is my visualization? Evaluating visualization accessibility with chartability. Computer Graphics Forum, 41(3):57–70, June 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Eom:2020:SCA

[ECC^{+}20] Haegwang Eom, Byungkuk Choi, Kyungmin Cho, Sunjin Jung, Seokpyo Hong, and Junyong Noh. Synthesizing character animation with smoothly decomposed motion layers. Computer Graphics Forum, 39(1):595–606, February 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Eichmann:2020:OEM

[EEE^{+}20] P. Eichmann, D. Edge, N. Evans, B. Lee, M. Brehmer, and C. White. Orchard: Exploring multivariate heterogeneous networks on mobile phones. Computer Graphics Forum, 39(3):115– 126, June 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Einabadi:2021:DNM

[EGH21] Farshad Einabadi, Jean-Yves Guillemaut, and Adrian Hilton. Deep neural models for illumination estimation and relighting: a survey. Computer Graphics Forum, 40(6):315–331, September 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Evers:2021:UAV

[EHL21] Marina Evers, Karim Huesmann, and Lars Linsen. Uncertaintyaware visualization of regional time series correlation in spatiotemporal ensembles. Computer Graphics Forum, 40(3):519–530, June 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Endo:2020:DSI

[EK20] Y. Endo and Y. Kanamori. Diversifying semantic image synthesis and editing via class- and layer-wise VAEs. Computer Graphics Forum, 39(7):519–530, October 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Epperson:2022:LAH

[ELW^{+}22] Will Epperson, Doris Jung-Lin Lee, Leijie Wang, Kunal Agarwal, Aditya G. Parameswaran, Dominik Moritz, and Adam Perer. Leveraging analysis history for improved in situ visualization recommendation. Computer Graphics Forum, 41(3):145–155, June 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Erleben:2020:MMA

[EMAK20] K. Erleben, M. Macklin, S. Andrews, and P. G. Kry. The matchstick model for anisotropic friction cones. Computer Graphics Forum, 39(1):450–461, February 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Eirich:2022:RDS

[EMJ^{+}22] J. Eirich, M. Münch, D. Jäckle, M. Sedlmair, J. Bonart, and T. Schreck. RfX: a design study for the interactive exploration of a random forest to enhance testing procedures for electrical engines. Computer Graphics Forum, 41(6):302–315, September 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Eulzer:2021:VCB

[EMKL21] P. Eulzer, M. Meuschke, C. M. Klingner, and K. Lawonn. Visualizing carotid blood flow simulations for stroke prevention. Computer Graphics Forum, 40(3):435–446, June 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Eulzer:2022:VMS

[EMML22] P. Eulzer, M. Meuschke, G. Mistelbauer, and K. Lawonn. Vessel maps: a survey of map-like visualizations of the cardiovascular system. Computer Graphics Forum, 41(3):645–673, June 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Endo:2022:UCL

[End22] Y. Endo. User-controllable latent transformer for StyleGAN image layout editing. Computer Graphics Forum, 41(7):395–406, October 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Eichner:2020:MPD

[EST20] Christian Eichner, Heidrun Schumann, and Christian Tominski. Making parameter dependencies of time-series segmentation visually understandable. Computer Graphics Forum, 39(1):607– 622, February 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Filoscia:2020:BRO

[FAG^{+}20] I. Filoscia, T. Alderighi, D. Giorgi, L. Malomo, M. Callieri, and P. Cignoni. Book review: Optimizing Object Decomposition to Reduce Visual Artifacts in 3D Printing. Computer Graphics Forum, 39(2):423–434, May 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Frohler:2022:SCV

[FAP^{+}22] B. Fröhler, C. Anthes, F. Pointecker, J. Friedl, D. Schwajda, A. Riegler, S. Tripathi, C. Holzmann, M. Brunner, H. Jodlbauer, H.-C. Jetter, and C. Heinzl. A survey on cross-virtuality analytics. Computer Graphics Forum, 41(1):465–494, February 2022.

CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Fratarcangeli:2020:FNL

[FBG^{+}20] M. Fratarcangeli, D. Bradley, A. Gruber, G. Zoss, and T. Beeler. Fast nonlinear least squares optimization of large-scale semisparse problems. Computer Graphics Forum, 39(2):247–259, May 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Fu:2020:FAR

[FCD20] Tong Fu, Raphaëlle Chaine, and Julie Digne. FAKIR: An algorithm for revealing the anatomy and pose of statues from raw point sets. Computer Graphics Forum, 39(7):375–385, October 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Firmino:2022:PDM

[FFJ22] Arthur Firmino, Jeppe Revall Frisvad, and Henrik Wann Jensen. Progressive denoising of Monte Carlo rendered images. Computer Graphics Forum, 41(2):1–11, May 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Franke:2021:TWF

[FFM^{+}21] Linus Franke, Laura Fink, Jana Martschinke, Kai Selgrad, and Marc Stamminger. Time-warped foveated rendering for virtual reality headsets. Computer Graphics Forum, 40(1):110–123, February 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Fu:2022:DAS

[FGZ^{+}22] Yanping Fu, Zhenyu Gai, Haifeng Zhao, Shaojie Zhang, Ying Shan, Yang Wu, and Jin Tang. Depth-aware shadow removal. Computer Graphics Forum, 41(7):455–464, October 2022. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Fellegara:2020:EHP

[FIDF20] Riccardo Fellegara, Federico Iuricich, Leila De Floriani, and Ulderico Fugacci. Efficient homology-preserving simplification of high-dimensional simplicial shapes. Computer Graphics Forum, 39(1):244–259, February 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Falster:2020:CBS

[FJF20] V. Falster, A. Jarabo, and J. R. Frisvad. Computing the bidirectional scattering of a microstructure using scalar diffraction theory and path tracing. Computer Graphics Forum, 39(7):231– 242, October 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Frisvad:2020:SMA

[FJM^{+}20] J. R. Frisvad, S. A. Jensen, J. S. Madsen, A. Correia, L. Yang, S. K. S. Gregersen, Y. Meuret, and P.-E. Hansen. Survey of models for acquiring the optical properties of translucent materials. Computer Graphics Forum, 39(2):729–755, May 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Futschik:2021:SST

[FKL^{+}21] D. Futschik, M. Kučera, M. Lukáč, Z. Wang, E. Shechtman, and D. Sýkora. STALP: Style transfer with auxiliary limited pairing. Computer Graphics Forum, 40(2):563–573, May 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Fan:2020:BRS

[FLH^{+}20a] Zhaoxin Fan, Hongyan Liu, Jun He, Qi Sun, and Xiaoyong Du. Book review: SRNet: A 3D Scene Recognition Network using Static Graph and Dense Semantic Fusion. Computer Graphics Forum, 39(7):301–311, October 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Fan:2020:GBO

[FLH^{+}20b] Zhaoxin Fan, Hongyan Liu, Jun He, Qi Sun, and Xiaoyong Du. A graph-based one-shot learning method for point cloud recognition. Computer Graphics Forum, 39(7):313–323, October 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Franke:2021:VAS

[FMKK21] M. Franke, H. Martin, S. Koch, and K. Kurzhals. Visual analysis of spatio-temporal phenomena with 1D projections. Computer Graphics Forum, 40(3):335–347, June 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Frey:2021:TDE

[Fre21] S. Frey. Temporally dense exploration of moving and deforming shapes. Computer Graphics Forum, 40(1):7–21, February 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Frey:2022:OGL

[Fre22] S. Frey. Optimizing grid layouts for level-of-detail exploration of large data collections. Computer Graphics Forum, 41(3):247– 258, June 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Fondevilla:2021:FTD

[FRH^{+}21] Amelie Fondevilla, Damien Rohmer, Stefanie Hahmann, Adrien Bousseau, and Marie-Paule Cani. Fashion transfer: Dressing 3D characters from stylized fashion sketches. Computer Graphics Forum, 40(6):466–483, September 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Frey:2022:VAT

[FSK^{+}22] Steffen Frey, Stefan Scheller, Nikolaos Karadimitriou, Dongwon Lee, Guido Reina, Holger Steeb, and Thomas Ertl. Visual analysis of two-phase flow displacement processes in porous media. Computer Graphics Forum, 41(1):243–256, February 2022. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Fischer:2021:CVA

[FSS^{+}21] M. T. Fischer, D. Seebacher, R. Sevastjanova, D. A. Keim, and M. El-Assady. CommAID: Visual analytics for communication analysis through interactive dynamics modeling. Computer Graphics Forum, 40(3):25–36, June 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Frezzato:2022:SGM

[FTA22] A. Frezzato, A. Tangri, and S. Andrews. Synthesizing get-up motions for physics-based characters. Computer Graphics Forum, 41(8):207–218, December 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Filip:2021:FCE

[FVMK21] J. Filip, R. Vávra, F. J. Maile, and M. Kolafová. Framework for capturing and editing of anisotropic effect coatings. Computer Graphics Forum, 40(1):68–80, February 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Fraboni:2022:VMV

[FWBI22] Basile Fraboni, Antoine Webanck, Nicolas Bonneel, and Jean-Claude Iehl. Volumetric multi-view rendering. Computer Graphics Forum, 41(2):379–392, May 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Fan:2021:RTM

[FWHB21] Hangming Fan, Rui Wang, Yuchi Huo, and Hujun Bao. Real-time Monte Carlo denoising with weight sharing kernel prediction network. Computer Graphics Forum, 40(4):15–27, July 2021. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Fu:2020:IDP

[FYFL20] Qiang Fu, Hai Yan, Hongbo Fu, and Xueming Li. Interactive design and preview of colored snapshots of indoor scenes. Computer Graphics Forum, 39(7):543–552, October 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Guehl:2020:SPT

[GAD^{+}20] P. Guehl, R. Allègre, J.-M. Dischler, B. Benes, and E. Galin. Semi-procedural textures using point process texture basis functions. Computer Graphics Forum, 39(4):159–171, July 2020. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gieseke:2021:SCM

[GAM^{+}21] Lena Gieseke, Paul Asente, Radomír Měch, Bedrich Benes, and Martin Fuchs. A survey of control mechanisms for creative pattern generation. Computer Graphics Forum, 40(2):585–609, May 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gleicher:2020:BIC

[GBYH20] Michael Gleicher, Aditya Barve, Xinyi Yu, and Florian Heimerl. Boxer: Interactive comparison of classifier results. Computer Graphics Forum, 39(3):181–193, June 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Guo:2022:CGT

[GCDT22] Yu-Xiao Guo, Guojun Chen, Yue Dong, and Xin Tong. Classifier guided temporal supersampling for real-time rendering. Computer Graphics Forum, 41(7):237–246, October 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Geiger:2021:COP

[GCH^{+}21] J. Geiger, S. Cornelsen, J.-H. Haunert, P. Kindermann, T. Mchedlidze, M. Nöllenburg, Y. Okamoto, and A. Wolff. ClusterSets: Optimizing planar clusters in categorical point data. Computer Graphics Forum, 40(3):471–481, June 2021. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gadhave:2022:RIA

[GCL22] K. Gadhave, Z. Cutler, and A. Lex. Reusing interactive analysis workflows. Computer Graphics Forum, 41(3):133–144, June 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Goude:2021:EBC

[GCLB21] Ific Goudé, Rémi Cozot, Olivier Le Meur, and Kadi Bouatouch. Example-based colour transfer for 3D point clouds. Computer Graphics Forum, 40(6):428–446, September 2021. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gong:2020:BRC

[GCZ^{+}20] Xun Gong, Ping Chen, Zhemin Zhang, Ke Chen, Yue Xiang, and Xin Li. Book review: A Cross-Dimension Annotations Method for 3D Structural Facial Landmark Extraction. Computer Graphics Forum, 39(1):623–636, February 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ghosh:2023:IID

[GDG^{+}23] Anindita Ghosh, Rishabh Dabral, Vladislav Golyanik, Christian Theobalt, and Philipp Slusallek. IMoS: Intent-driven fullbody motion synthesis for human–object interactions. Computer Graphics Forum, 42(2):1–12, May 2023. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Guo:2021:GSR

[GE21] J. Guo and E. Eisemann. Geometric sample reweighting for Monte Carlo integration. Computer Graphics Forum, 40(7):109– 119, October 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Guo:2020:NEE

[GEE20] Jerry Jinfeng Guo, Martin Eisemann, and Elmar Eisemann. Next event estimation++: Visibility mapping for efficient light transport simulation. Computer Graphics Forum, 39(7):205–217, October 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ghorbani:2023:ZZS

[GFH^{+}23] Saeed Ghorbani, Ylva Ferstl, Daniel Holden, Nikolaus F. Troje, and Marc-André Carbonneau. ZeroEGGS: Zero-shot examplebased gesture generation from speech. Computer Graphics Forum, 42(1):206–216, February 2023. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gruber:2020:ISD

[GFZ^{+}20] A. Gruber, M. Fratarcangeli, G. Zoss, R. Cattaneo, T. Beeler, M. Gross, and D. Bradley. Interactive sculpting of digital faces using an anatomical modeling paradigm. Computer Graphics Forum, 39(5):93–102, August 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gitlina:2020:PMR

[GGD^{+}20] Y. Gitlina, G. C. Guarnera, D. S. Dhillon, J. Hansen, A. Lattas, D. Pai, and A. Ghosh. Practical measurement and reconstruction of spectral skin reflectance. Computer Graphics Forum, 39(4): 75–89, July 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gagnon:2021:PED

[GGMP21] Jonathan Gagnon, Julián E. Guzmán, David Mould, and Eric Paquette. Patch erosion for deformable lapped textures on 3D fluids. Computer Graphics Forum, 40(2):367–374, May 2021. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gamboa:2020:ETE

[GGN20] Luis E. Gamboa, Adrien Gruson, and Derek Nowrouzezahrai. An efficient transport estimator for complex layered materials. Computer Graphics Forum, 39(2):363–371, May 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Galin:2020:STU

[GGPP20] Eric Galin, Eric Guérin, Axel Paris, and Adrien Peytavie. Segment tracing using local Lipschitz bounds. Computer Graphics Forum, 39(2):545–554, May 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Grittmann:2021:CAM

[GGS21] Pascal Grittmann, Iliyan Georgiev, and Philipp Slusallek. Correlation-aware multiple importance sampling for bidirectional rendering algorithms. Computer Graphics Forum, 40(2):231–238, May 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gotsman:2021:LDP

[GH21] C. Gotsman and K. Hormann. On landmark distances in polygons. Computer Graphics Forum, 40(5):275–287, August 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gagniere:2020:HLE

[GHMR^{+}20] S. Gagniere, D. Hyde, A. Marquez-Razon, C. Jiang, Z. Ge, X. Han, Q. Guo, and J. Teran. A hybrid Lagrangian/Eulerian collocated velocity advection and projection method for fluid simulation. Computer Graphics Forum, 39(8):1–14, December 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Guo:2020:BIF

[GHYZ20] Y. Guo, M. Hašan, L. Yan, and S. Zhao. A Bayesian inference framework for procedural material parameter estimation. Computer Graphics Forum, 39(7):255–266, October 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gao:2022:ETB

[GK22] Hong Gao and Enkelejda Kasneci. Eye-tracking-based prediction of user experience in VR locomotion using machine learning. Computer Graphics Forum, 41(7):589–599, October 2022. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gaillard:2022:ADP

[GKG^{+}22] Mathieu Gaillard, Vojtěch Krs, Giorgio Gori, Radomír Měch, and Bedrich Benes. Automatic differentiable procedural modeling.

Computer Graphics Forum, 41(2):289–307, May 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Golla:2020:TUP

[GKK^{+}20] Tim Golla, Tom Kneiphof, Heiner Kuhlmann, Michael Weinmann, and Reinhard Klein. Temporal upsampling of point cloud sequences by optimal transport for plant growth visualization. Computer Graphics Forum, 39(6):167–179, September 2020. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Garrison:2022:TOV

[GKV^{+}22] Laura A. Garrison, Ivan Kolesar, Ivan Viola, Helwig Hauser, and Stefan Bruckner. Trends & opportunities in visualization for physiology: a multiscale overview. Computer Graphics Forum, 41(3):609–643, June 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gong:2020:BRD

[GLLL20] Xun Gong, Xin Li, Tianrui Li, and Yongqing Liang. Book review: A Discriminative Multi-Channel Facial Shape (MCFS) Representation and Feature Extraction for 3D Human Faces. Computer Graphics Forum, 39(6):66–81, September 2020. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Goel:2022:IMM

[GMH22] Aman Goel, Qianhui Men, and Edmond S. L. Ho. Interaction mix and match: Synthesizing close interaction using conditional hierarchical GAN with multi-hot class embedding. Computer Graphics Forum, 41(8):327–338, December 2022. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gathani:2022:GBA

[GMOB22] Sneha Gathani, Shayan Monadjemi, Alvitta Ottley, and Leilani Battle. A grammar-based approach for applying visualization taxonomies to interaction logs. Computer Graphics Forum, 41 (3):489–500, June 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Germano:2021:RTF

[GOG21] Rafael L. Germano, Manuel M. Oliveira, and Eduardo S. L. Gastal. Real-time frequency adjustment of images and videos. Computer Graphics Forum, 40(2):23–37, May 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Guerin:2022:GTA

[GPM^{+}22] Eric Guérin, Adrien Peytavie, Simon Masnou, Julie Digne, Basile Sauvage, James Gain, and Eric Galin. Gradient terrain authoring. Computer Graphics Forum, 41(2):85–95, May 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gutierrez:2020:BRD

[GRGH20] J. Gutierrez, J. Rabin, B. Galerne, and T. Hurtut. Book review: On Demand Solid Texture Synthesis Using Deep 3D Networks. Computer Graphics Forum, 39(1):511–530, February 2020. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Griffiths:2022:OOS

[GRP22] David Griffiths, Tobias Ritschel, and Julien Philip. OutCast: Outdoor single-image relighting with cast shadows. Computer Graphics Forum, 41(2):179–193, May 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Grenier:2022:CMN

[GSDT22] C. Grenier, B. Sauvage, J.-M. Dischler, and S. Thery. Colormapped noise vector fields for generating procedural micropatterns. Computer Graphics Forum, 41(7):477–487, October 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gillmann:2021:UAV

[GSWS21] Christina Gillmann, Dorothee Saur, Thomas Wischgoll, and Gerik Scheuermann. Uncertainty-aware visualization in medical imaging — a survey. Computer Graphics Forum, 40(3):665–689, June 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gruen:2023:MAB

[GvdVS23] J. Gruen, G. van der Voort, and T. Schultz. Model averaging and bootstrap consensus-based uncertainty reduction in diffusion MRI tractography. Computer Graphics Forum, 42(1):217–230, February 2023. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Garcia:2021:CMB

[GVF^{+}21] M. Garcia, R. Vergne, M.-A. Farhat, P. Bénard, C. Noûs, and J. Thollot. Coherent mark-based stylization of 3D scenes at the compositing stage. Computer Graphics Forum, 40(2):39–49, May 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gonda:2021:VVI

[GWB^{+}21] Felix Gonda, Xueying Wang, Johanna Beyer, Markus Hadwiger, Jeff W. Lichtman, and Hanspeter Pfister. VICE: Visual identification and correction of neural circuit errors. Computer Graphics Forum, 40(3):447–458, June 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ghorbani:2020:PCM

[GWE^{+}20] S. Ghorbani, C. Wloka, A. Etemad, M. A. Brubaker, and N. F. Troje. Probabilistic character motion synthesis using a hierarchical deep latent variable model. Computer Graphics Forum, 39(8): 225–239, December 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gruson:2020:SMC

[GWH20] Adrien Gruson, Rex West, and Toshiya Hachisuka. Stratified Markov chain Monte Carlo light transport. Computer Graphics Forum, 39(2):351–362, May 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gao:2022:GEE

[GWM22] Lin Gao, Lei Wu, and Xiangxu Meng. EL-GAN: Edge-enhanced generative adversarial network for layout-to-image generation. Computer Graphics Forum, 41(7):407–418, October 2022. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Guo:2022:SSB

[GZG^{+}22] Bao Guo, Yuhe Zhang, Jian Gao, Chunhui Li, and Yao Hu. SGLBP: Subgraph-based local binary patterns for feature extraction on point clouds. Computer Graphics Forum, 41(6):51–66, September 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gao:2020:ALS

[GZL^{+}20a] Yang Gao, Quancheng Zhang, Shuai Li, Aimin Hao, and Hong Qin. Accelerating liquid simulation with an improved data-driven method. Computer Graphics Forum, 39(6):180–191, September 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ge:2020:CHL

[GZL^{+}20b] T. Ge, Y. Zhao, B. Lee, D. Ren, B. Chen, and Y. Wang. Canis: A high-level language for data-driven chart animations. Computer Graphics Forum, 39(3):607–617, June 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Gao:2023:HHD

[GZLL23] Jian Gao, Yuhe Zhang, Zehua Liu, and Siyi Li. HDRNet: High-dimensional regression network for point cloud registration. Computer Graphics Forum, 42(1):33–46, February 2023. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hauser:2022:Ea

[HA22] Helwig Hauser and Pierre Alliez. Editorial. Computer Graphics Forum, 41(1):5–6, February 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hanocka:2022:LCC

[HACOG22] R. Hanocka, J. Assa, D. Cohen-Or, and R. Giryes. Learning camera control in dynamic scenes from limited demonstrations. Computer Graphics Forum, 41(1):427–437, February 2022. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hauser:2020:X

[HB20] Helwig Hauser and Bedrich Benes. 2019 editorial v2. Computer Graphics Forum, 39(1):5–6, February 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hauser:2021:E

[HB21] Helwig Hauser and Bedrich Benes. Editorial. Computer Graphics Forum, 40(1):5–6, February 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hergl:2021:VTF

[HBK^{+}21] Chiara Hergl, Christian Blecha, Vanessa Kretzschmar, Felix Raith, Fabian Günther, Markus Stommel, Jochen Jankowai, Ingrid Hotz, Thomas Nagel, and Gerik Scheuermann. Visualization of tensor fields in mechanics. Computer Graphics Forum, 40(6): 135–161, September 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hu:2022:SSP

[HCL^{+}22] Fei Hu, Honghua Chen, Xuequan Lu, Zhe Zhu, Jun Wang, Weiming Wang, Fu Lee Wang, and Mingqiang Wei. SPCNet: Stepwise point cloud completion network. Computer Graphics Forum, 41 (7):153–164, October 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hopkins:2020:VSS

[HCS20] Aspen K. Hopkins, Michael Correll, and Arvind Satyanarayan. VisuaLint: Sketchy in situ annotations of chart construction errors. Computer Graphics Forum, 39(3):219–228, June 2020. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Healey:2021:VAT

[HDP^{+}21] Christopher G. Healey, Gowtham Dinakaran, Kalpesh Padia, Shaoliang Nie, J. Riley Benson, Dave Caira, Dean Shaw, Gary Catalfu, and Ravi Devarajan. Visual analytics of text conversation sentiment and semantics. Computer Graphics Forum, 40(6): 484–499, September 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Heitz:2020:CIC

[Hei20] E. Heitz. Can’t invert the CDF? the triangle-cut parameterization of the region under the curve. Computer Graphics Forum, 39(4):121–132, July 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hu:2020:DDR

[HGC^{+}20] Bingyang Hu, Jie Guo, Yanjun Chen, Mengtian Li, and Yanwen Guo. DeepBRDF: A deep representation for manipulating measured BRDF. Computer Graphics Forum, 39(2):157–166, May 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hu:2022:CMA

[HHG^{+}22] Yiwei Hu, Miloš Hašan, Paul Guerrero, Holly Rushmeier, and Valentin Deschaintre. Controlling material appearance by examples. Computer Graphics Forum, 41(4):117–128, July 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Herter:2021:TVV

[HHH^{+}21] Felix Herter, Hans-Christian Hege, Markus Hadwiger, Verena Lepper, and Daniel Baum. Thin-volume visualization on curved domains. Computer Graphics Forum, 40(3):147–157, June 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Huang:2022:MBH

[HHH22] W. Huang, M. B. Hullin, and J. Hanika. A microfacet-based hair scattering model. Computer Graphics Forum, 41(4):79–91, July 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hografer:2020:SAM

[HHS20] Marius Hogräfer, Magnus Heitzler, and Hans-Jörg Schulz. The state of the art in map-like visualization. Computer Graphics Forum, 39(3):647–674, June 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hillaire:2020:SPR

[Hil20] Sébastien Hillaire. A scalable and production ready sky and atmosphere rendering technique. Computer Graphics Forum, 39(4): 13–22, July 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Huo:2020:SGB

[HJL^{+}20] Y. C. Huo, S. H. Jin, T. Liu, W. Hua, R. Wang, and H. J. Bao. Spherical Gaussian-based lightcuts for glossy interreflections. Computer Graphics Forum, 39(6):192–203, September 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Huang:2021:VDL

[HJZ^{+}21] Xinyi Huang, Suphanut Jamonnak, Ye Zhao, Tsung Heng Wu, and Wei Xu. A visual designer of layer-wise relevance propagation models. Computer Graphics Forum, 40(3):227–238, June 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hoque:2022:CQA

[HKM22] E. Hoque, P. Kavehzadeh, and A. Masry. Chart question answering: State of the art and future directions. Computer Graphics Forum, 41(3):555–572, June 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Huesmann:2022:SDN

[HL22] Karim Huesmann and Lars Linsen. SimilarityNet: a deep neural network for similarity analysis within spatio-temporal ensembles. Computer Graphics Forum, 41(3):379–389, June 2022. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hu:2020:BRC

[HLJ^{+}20] Hailong Hu, Zhong Li, Xiaogang Jin, Zhigang Deng, Minhong Chen, and Yi Shen. Book review: Curve Skeleton Extraction From 3D Point Clouds Through Hybrid Feature Point Shifting and Clustering. Computer Graphics Forum, 39(6):111–132, September 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hall:2022:STV

[HLJ^{+}22] Brian D. Hall, Yang Liu, Yvonne Jansen, Pierre Dragicevic, Fanny Chevalier, and Matthew Kay. A survey of tasks and visualizations in multiverse analysis reports. Computer Graphics Forum, 41(1):402–426, February 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hajisharif:2020:SSC

[HMGU20] Saghi Hajisharif, Ehsan Miandji, Christine Guillemot, and Jonas Unger. Single sensor compressive light field video camera. Computer Graphics Forum, 39(2):463–474, May 2020. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hasselgren:2020:NTA

[HMS^{+}20] J. Hasselgren, J. Munkberg, M. Salvi, A. Patney, and A. Lefohn. Neural temporal adaptive sampling and denoising. Computer Graphics Forum, 39(2):147–155, May 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Huang:2022:PPS

[HMSS22] Xingchang Huang, Pooran Memari, Hans-Peter Seidel, and Gurprit Singh. Point-pattern synthesis using Gabor and random filters. Computer Graphics Forum, 41(4):169–179, July 2022. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Han:2023:PVA

[HNWC23] Dongyun Han, Abdullah-Al-Raihan Nayeem, Jason Windett, and Isaac Cho. PDViz: a visual analytics approach for state policy data. Computer Graphics Forum, 42(1):305–318, February 2023. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hertz:2023:MDP

[HPG^{+}23] A. Hertz, O. Perel, R. Giryes, O. Sorkine-Hornung, and D. Cohen-Or. Mesh draping: Parametrization-free neural mesh transfer. Computer Graphics Forum, 42(1):72–85, February 2023. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hwang:2022:TMS

[HPKI22] Jaepyung Hwang, Gangrae Park, Taesoo Kwon, and Shin Ishii. Transition motion synthesis for object interaction based on learning transition strategies. Computer Graphics Forum, 41(6):37– 50, September 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hart:2020:PPS

[HPM^{+}20] D. Hart, M. Pharr, T. Müller, W. Lopes, M. McGuire, and P. Shirley. Practical product sampling by fitting and composing warps. Computer Graphics Forum, 39(4):149–158, July 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

He:2022:FNC

[HPO22] Zhongyun He, Jesús Pérez, and Miguel A. Otaduy. Fast numerical coarsening with local factorizations. Computer Graphics Forum, 41(8):9–17, December 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hwang:2021:POG

[HPSK21] Jae-Pyung Hwang, Gangrae Park, Il Hong Suh, and Taesoo Kwon. Primitive object grasping for finger motion synthesis. Computer Graphics Forum, 40(1):266–278, February 2021. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Huang:2020:COZ

[HRWO20] Ruqi Huang, Jing Ren, Peter Wonka, and Maks Ovsjanikov. Consistent Zoom Out: Efficient spectral map synchronization. Computer Graphics Forum, 39(5):265–278, August 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hofmann:2020:BRE

[HS20] Lutz Hofmann and Filip Sadlo. Book review: Extraction of Distinguished Hyperbolic Trajectories for 2D Time-Dependent Vector Field Topology. Computer Graphics Forum, 39(3):303–316, June 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hofmann:2021:LET

[HS21] Lutz Hofmann and Filip Sadlo. Local extraction of 3D timedependent vector field topology. Computer Graphics Forum, 40 (3):111–122, June 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Huang:2022:RTD

[HSG^{+}22] Tao Huang, Yadong Song, Jie Guo, Chengzhi Tao, Zijing Zong, Xihao Fu, Hongshan Li, and Yanwen Guo. Real-time deep radiance reconstruction from imperfect caches. Computer Graphics Forum, 41(7):267–278, October 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hladky:2021:SET

[HSS21] J. Hladky, H. P. Seidel, and M. Steinberger. SnakeBinning: Efficient temporally coherent triangle packing for shading streaming. Computer Graphics Forum, 40(2):475–488, May 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hartwig:2022:LHV

[HSvO^{+}22] S. Hartwig, M. Schelling, C. v. Onzenoodt, P.-P. Vázquez, P. Hermosilla, and T. Ropinski. Learning human viewpoint preferences from sparsely annotated models. Computer Graphics Forum, 41 (6):453–466, September 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hass:2020:AIG

[HT20] J. Hass and M. Trnkova. Approximating isosurfaces by guaranteed-quality triangular meshes. Computer Graphics Forum, 39(5):29–40, August 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hua:2023:ESI

[HTFW23] Q. Hua, V. Tázlar, A. Fichet, and A. Wilkie. Efficient storage and importance sampling for fluorescent reflectance. Computer Graphics Forum, 42(1):47–59, February 2023. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hong:2022:LIV

[HTI22] Jiayi Hong, Alain Trubuil, and Tobias Isenberg. LineageD: an interactive visual system for plant cell lineage assignments based on correctable machine learning. Computer Graphics Forum, 41 (3):195–207, June 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hauptfleisch:2020:BRS

[HTT^{+}20] F. Hauptfleisch, O. Texler, A. Texler, J. Krivánek, and D. Sýkora. Book review: StyleProp: Real-time Example-based Stylization of 3D Models. Computer Graphics Forum, 39(7):575–586, October 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Han:2022:SLS

[HW22] Jun Han and Chaoli Wang. SurfNet: Learning surface representations via graph convolutional network. Computer Graphics Forum, 41(3):109–120, June 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

He:2021:MWU

[HWC21] Zhenbang He, Yunhai Wang, and Zhanglin Cheng. Manhattanworld urban building reconstruction by fitting cubes. Computer Graphics Forum, 40(7):289–300, October 2021. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hanika:2022:OMS

[HWD22] Johannes Hanika, Andrea Weidlich, and Marc Droske. Oncemore scattered next event estimation for volume rendering. Computer Graphics Forum, 41(4):17–28, July 2022. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Han:2021:DPL

[HWDL21] Fangzhou Han, Can Wang, Hao Du, and Jing Liao. Deep portrait lighting enhancement with 3D guidance. Computer Graphics Forum, 40(4):177–188, July 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Han:2022:STK

[HWJ22] Yi Han, He Wang, and Xiaogang Jin. Spatio-temporal keyframe control of traffic simulation using coarse-to-fine optimization. Computer Graphics Forum, 41(7):541–552, October 2022. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Han:2022:RZZ

[HWW^{+}22] Xianjun Han, Xue Wang, Huabin Wang, Xuejun Li, and Hongyu Yang. Ref-ZSSR: Zero-shot single image superresolution with reference image. Computer Graphics Forum, 41(7):311–322, October 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Han:2021:LPB

[HXA21] Chengguizi Han, Tao Xue, and Mridul Aanjaneya. A Lagrangian particle-based formulation for coupled simulation of fracture and diffusion in thin membranes. Computer Graphics Forum, 40(7): 97–108, October 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Huang:2022:MMO

[HXW^{+}22] A. Huang, Q. Xie, Z. Wang, D. Lu, M. Wei, and J. Wang. MODNet: Multi-offset point cloud denoising network customized for multiscale patches. Computer Graphics Forum, 41(7):109–119, October 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hsiao:2023:IGG

[HYC^{+}23] Kai-Wen Hsiao, Yong-Liang Yang, Yung-Chih Chiu, Min-Chun Hu, Chih-Yuan Yao, and Hung-Kuo Chu. Img2Logo: Generating golden ratio logos from images. Computer Graphics Forum, 42(2):37–49, May 2023. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Hu:2022:MSP

[HYW^{+}22] Yiwei Hu, Yazhen Yuan, Rui Wang, Zhuo Yang, and Hujun Bao. Multirate shading with piecewise interpolatory approximation. Computer Graphics Forum, 41(7):255–266, October 2022. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Heep:2022:SSB

[HZ22] M. Heep and E. Zell. ShadowPatch: Shadow based segmentation for reliable depth discontinuities in photometric stereo. Computer Graphics Forum, 41(7):635–646, October 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Huang:2020:BRO

[HZSQ20] Hong Huang, Fan Zhong, Yuqing Sun, and Xueying Qin. Book review: An Occlusion-aware Edge-Based Method for Monocular 3D Object Tracking using Edge Confidence. Computer Graphics Forum, 39(7):399–409, October 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ikemakhen:2021:BHC

[IA21] A. Ikemakhen and T. Ahanchaou. Blending of hyperbolic closed curves. Computer Graphics Forum, 40(5):71–79, August 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Igarashi:2022:PAA

[II22] Yuki Igarashi and Takeo Igarashi. Pixel art adaptation for handicraft fabrication. Computer Graphics Forum, 41(7):489–494, October 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Inoue:2020:RAO

[IIHG^{+}20] N. Inoue, D. Ito, Y. Hold-Geoffroy, L. Mai, B. Price, and T. Yamasaki. RGB2AO: Ambient occlusion generation from RGB images. Computer Graphics Forum, 39(2):451–462, May 2020. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ivanova:2023:SAF

[IWH23] D. Ivanova, J. Williamson, and P. Henderson. Simulating analogue film damage to analyse and improve artefact restoration on high-resolution scans. Computer Graphics Forum, 42(2):133– 148, May 2023. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jhang:2022:SMB

[JC22] Jia-Wun Jhang and Chun-Fa Chang. Specular manifold bisection sampling for caustics rendering. Computer Graphics Forum, 41 (7):247–254, October 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jakob:2021:OLC

[JG21] J. Jakob and M. Guthe. Optimizing LBVH-Construction and hierarchy-traversal to accelerate k NN queries on point clouds using the GPU. Computer Graphics Forum, 40(1):124–137, February 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jahan:2021:SGL

[JGvK21] Tansin Jahan, Yanran Guan, and Oliver van Kaick. Semanticsguided latent space exploration for shape generation. Computer Graphics Forum, 40(2):115–126, May 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Johnson:2022:CMA

[JH22] Colin Johnson and Brandon Haworth. Cognitive model of agent exploration with vision and signage understanding. Computer Graphics Forum, 41(8):143–154, December 2022. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jeschke:2020:MPW

[JHC^{+}20] S. Jeschke, C. Hafner, N. Chentanez, M. Macklin, M. Müller-Fischer, and C. Wojtan. Making procedural water waves boundary-aware. Computer Graphics Forum, 39(8):47–54, December 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jung:2020:SMB

[JHD20] A. Jung, J. Hanika, and C. Dachsbacher. Spectral mollification for bidirectional fluorescence. Computer Graphics Forum, 39(2): 373–384, May 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jiang:2022:RRF

[JJZ^{+}22] Diqiong Jiang, Yiwei Jin, Fang-Lue Zhang, Yu-Kun Lai, Risheng Deng, Ruofeng Tong, and Min Tang. Reconstructing recognizable 3D face shapes based on 3D morphable models. Computer Graphics Forum, 41(6):348–364, September 2022. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jiang:2021:DMM

[JL21] Y. Jiang and Y. Lan. A dynamic mixture model for nonequilibrium multiphase fluids. Computer Graphics Forum, 40 (7):85–95, October 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jeong:2022:IAD

[JLB22] Sangwon Jeong, Shusen Liu, and Matthew Berger. Interactively assessing disentanglement in GANs. Computer Graphics Forum, 41(3):85–95, June 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jiang:2020:DFM

[JLDH20] Y. Jiang, C. Li, S. Deng, and S. M. Hu. A divergence-free mixture model for multiphase fluids. Computer Graphics Forum, 39(8): 69–77, December 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jimenez:2022:GTV

[JM22] Edgar Jiménez and Rodrigo Macías. Graphical tools for visualization of missing data in large longitudinal phenomena. Computer Graphics Forum, 41(1):438–452, February 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Juarez-Perez:2020:DSF

[JPK20] A. Juarez-Perez and M. Kallmann. Detection and synthesis of full-body environment interactions for virtual humans. Computer Graphics Forum, 39(1):274–286, February 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jiang:2022:SOE

[JSG^{+}22] J. Jiang, X. Shen, Y. Gong, Z. Fan, Y. Liu, G. Xing, X. Ren, and Y. Zhang. A second-order explicit pressure projection method for Eulerian fluid simulation. Computer Graphics Forum, 41(8):

95–105, December 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jiang:2022:SBD

[JSL22] Jie Jiang, Hock Soon Seah, and Hong Ze Liew. Stroke-based drawing and inbetweening with boundary strokes. Computer Graphics Forum, 41(1):257–269, February 2022. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jourdan:2022:CDS

[JSVB22] David Jourdan, Mélina Skouras, Etienne Vouga, and Adrien Bousseau. Computational design of self-actuated surfaces by printing plastic ribbons on stretched fabric. Computer Graphics Forum, 41(2):493–506, May 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jaunet:2020:DUD

[JVW20] T. Jaunet, R. Vuillemot, and C. Wolf. DRLViz: Understanding decisions and memory in deep reinforcement learning. Computer Graphics Forum, 39(3):49–61, June 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jin:2020:PBF

[JZGF20] N. Jin, Y. Zhu, Z. Geng, and R. Fedkiw. A pixel-based framework for data-driven clothing. Computer Graphics Forum, 39(8): 135–144, December 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Jiang:2022:LMS

[JZN^{+}22] Hao Jiang, Qing Zhang, Yongwei Nie, Lei Zhu, and Wei-Shi Zheng. Learning multi-scale deep image prior for high-quality unsupervised image denoising. Computer Graphics Forum, 41 (7):323–334, October 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ji:2020:DRN

[JZZ^{+}20] Zhongping Ji, Chengqin Zhou, Qiankan Zhang, Yu-Wei Zhang, and Wenping Wang. A deep residual network for geometric decontouring. Computer Graphics Forum, 39(7):27–41, October 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kerim:2021:NRV

[KAC^{+}21] Abdulrahman Kerim, Cem Aslan, Ufuk Celikcan, Erkut Erdem, and Aykut Erdem. NOVA: Rendering virtual worlds with humans for computer vision tasks. Computer Graphics Forum, 40(6): 258–272, September 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kerbiriou:2022:DER

[KADM22] Glenn Kerbiriou, Quentin Avril, Fabien Danieau, and Maud Marchal. Detailed eye region capture and animation. Computer Graphics Forum, 41(8):279–282, December 2022. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kwiatkowski:2022:SRL

[KAK^{+}22] Ariel Kwiatkowski, Eduardo Alvarado, Vicky Kalogeiton, C. Karen Liu, Julien Pettré, Michiel van de Panne, and Marie-Paule Cani. A survey on reinforcement learning methods in character animation. Computer Graphics Forum, 41(2):613–639, May 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Koschier:2022:SSM

[KBST22] Dan Koschier, Jan Bender, Barbara Solenthaler, and Matthias Teschner. A survey on SPH methods in computer graphics. Computer Graphics Forum, 41(2):737–760, May 2022. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kalshetti:2022:LSA

[KC22] P. Kalshetti and P. Chaudhuri. Local scale adaptation to hand shape model for accurate and robust hand tracking. Computer Graphics Forum, 41(8):219–229, December 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kazhdan:2020:PSR

[KCRH20] Misha Kazhdan, Ming Chuang, Szymon Rusinkiewicz, and Hugues Hoppe. Poisson surface reconstruction with envelope constraints. Computer Graphics Forum, 39(5):173–182, August 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Khartabil:2021:DEV

[KCW^{+}21] D. Khartabil, C. Collins, S. Wells, B. Bach, and J. Kennedy. Design and evaluation of visualization techniques to facilitate argument exploration. Computer Graphics Forum, 40(6):447–465, September 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kwan:2021:AIC

[KF21] K. C. Kwan and H. Fu. Automatic image checkpoint selection for guider-follower pedestrian navigation. Computer Graphics Forum, 40(1):357–368, February 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kohlbrenner:2021:GSI

[KFDA21] M. Kohlbrenner, U. Finnendahl, T. Djuren, and M. Alexa. Gauss stylization: Interactive artistic mesh modeling based on preferred surface normals. Computer Graphics Forum, 40(5):33–43, August 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kraus:2022:IAA

[KFS^{+}22] Matthias Kraus, Johannes Fuchs, Björn Sommer, Karsten Klein, Ulrich Engelke, Daniel Keim, and Falk Schreiber. Immersive analytics with abstract 3D visualizations: a survey. Computer Graphics Forum, 41(1):201–229, February 2022. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kim:2022:DRS

[KHW^{+}22] Byungsoo Kim, Xingchang Huang, Laura Wuelfroth, Jingwei Tang, Guillaume Cordonnier, Markus Gross, and Barbara Solenthaler. Deep reconstruction of 3D smoke densities from artist sketches. Computer Graphics Forum, 41(2):97–110, May 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kim:2020:FEF

[Kim20] Theodore Kim. A finite element formulation of baraff-witkin cloth. Computer Graphics Forum, 39(8):171–179, December 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kips:2022:RTV

[KJB^{+}22] R. Kips, R. Jiang, S. Ba, B. Duke, M. Perrot, P. Gori, and I. Bloch. Real-time virtual-try-on from a single example image through deep inverse graphics and learned differentiable renderers. Computer Graphics Forum, 41(2):29–40, May 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kim:2021:AVD

[KJRK21] N. W. Kim, S. C. Joyner, A. Riegelhuth, and Y. Kim. Accessible visualization: Design space, opportunities, and challenges. Computer Graphics Forum, 40(3):173–188, June 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kim:2021:DLB

[KJS^{+}21] Seonghyeon Kim, Sunjin Jung, Kwanggyoon Seo, Roger Blanco i Ribera, and Junyong Noh. Deep learning-based unsupervised human facial retargeting. Computer Graphics Forum, 40(7):45– 55, October 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kim:2021:TBE

[KK21] Dongjoon Kim and Heewon Kye. Z-thickness blending: Effective fragment merging for multi-fragment rendering. Computer Graphics Forum, 40(7):149–160, October 2021. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Khulusi:2020:SVM

[KKM^{+}20] R. Khulusi, J. Kusnick, C. Meinecke, C. Gillmann, J. Focht, and S. Jänicke. A survey on visualizations for musical data. Computer Graphics Forum, 39(6):82–110, September 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kim:2021:STR

[KKN^{+}21] Hyomin Kim, Jungeon Kim, Hyeonseo Nam, Jaesik Park, and Seungyong Lee. Spatiotemporal texture reconstruction for dynamic objects using a single RGB-D camera. Computer Graphics Forum, 40(2):523–535, May 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Koike:2020:AEL

[KMA20] T. Koike, S. Morishima, and R. Ando. Asynchronous Eulerian liquid simulation. Computer Graphics Forum, 39(2):1–8, May 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Khanteimouri:2022:RBG

[KMC22] P. Khanteimouri, M. Mandad, and M. Campen. Rational Bézier guarding. Computer Graphics Forum, 41(5):89–99, August 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kim:2021:DPT

[KMH21] Hyeok Kim, Dominik Moritz, and Jessica Hullman. Design patterns and trade-offs in responsive visualization for communication. Computer Graphics Forum, 40(3):459–470, June 2021. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kirgo:2021:WBH

[KMP^{+}21] Maxime Kirgo, Simone Melzi, Giuseppe Patanè, Emanuele Rodolà, and Maks Ovsjanikov. Wavelet-based heat kernel derivatives: Towards informative localized shape analysis. Computer Graphics Forum, 40(1):165–179, February 2021. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kobrtek:2022:CMO

[KMTH22] Jozef Kobrtek, Tomas Milet, Michal Tóth, and Adam Herout. Comparison of modern omnidirectional precise shadowing techniques versus ray tracing. Computer Graphics Forum, 41(1):106– 121, February 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kuriyama:2022:CBS

[KMTM22] Shigeru Kuriyama, Tomohiko Mukai, Takafumi Taketomi, and Tomoyuki Mukasa. Context-based style transfer of tokenized gestures. Computer Graphics Forum, 41(8):305–315, December 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kikuchi:2021:MVC

[KOYSS21] K. Kikuchi, M. Otani, K. Yamaguchi, and E. Simo-Serra. Modeling visual containment for Web page layout optimization. Computer Graphics Forum, 40(7):33–44, October 2021. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Karciauskas:2022:PAB

[KP22] K. Karčiauskas and J. Peters. Point-augmented bi-cubic subdivision surfaces. Computer Graphics Forum, 41(7):13–23, October 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Karciauskas:2023:EGS

[KP23] K. Karčiauskas and J. Peters. Evolving guide subdivision. Computer Graphics Forum, 42(2):321–332, May 2023. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kim:2020:MRB

[KPKH20] SangBin Kim, Inbum Park, Seongsu Kwon, and JungHyun Han. Motion retargetting based on dilated convolutions and skeletonspecific loss functions. Computer Graphics Forum, 39(2):497– 507, May 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kopanas:2021:PBN

[KPLD21] Georgios Kopanas, Julien Philip, Thomas Leimkühler, and George Drettakis. Point-based neural rendering with per-view optimization. Computer Graphics Forum, 40(4):29–43, July 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kakar:2021:SDE

[KQL^{+}21] T. Kakar, X. Qin, T. La, S. K. Sahoo, S. De, E. A. Rundensteiner, and L. Harrison. SumRe: Design and evaluation of a gistbased summary visualization for incident reports triage. Computer Graphics Forum, 40(3):263–274, June 2021. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kreskowski:2022:EDI

[KRF22] A. Kreskowski, G. Rendle, and B. Froehlich. Efficient direct isosurface rasterization of scalar volumes. Computer Graphics Forum, 41(7):215–226, October 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Karer:2020:CFM

[KSHL20] B. Karer, I. Scheler, H. Hagen, and H. Leitte. ConceptGraph: A formal model for interpretation and reasoning during visual analysis. Computer Graphics Forum, 39(6):5–18, September 2020.

CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kee:2023:OBS

[KUKH23] Min Hyung Kee, Kiwon Um, HyunMo Kang, and JungHyun Han. An optimization-based SPH solver for simulation of hyperelastic solids. Computer Graphics Forum, 42(2):225–233, May 2023. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kushner:2021:LRO

[KUS^{+}21] Sarah Kushner, Risa Ulinski, Karan Singh, David I. W. Levin, and Alec Jacobson. Levitating rigid objects with hidden rods and wires. Computer Graphics Forum, 40(2):221–230, May 2021. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kluge:2020:PRI

[KWK20] M. Kluge, T. Weyrich, and A. Kolb. Progressive refinement imaging. Computer Graphics Forum, 39(1):360–374, February 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liu:2022:ZDA

[LADH22] Sean J. Liu, Maneesh Agrawala, Stephen DiVerdi, and Aaron Hertzmann. ZoomShop: Depth-aware editing of photographic composition. Computer Graphics Forum, 41(2):57–70, May 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Loizou:2020:BRL

[LAK20] Marios Loizou, Melinos Averkiou, and Evangelos Kalogerakis. Book review: Learning Part Boundaries from 3D Point Clouds. Computer Graphics Forum, 39(5):183–195, August 2020. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Leon:2020:CCV

[LB20] G. Molina León and A. Breiter. Co-creating visualizations: A first evaluation with social science researchers. Computer Graphics Forum, 39(3):291–302, June 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lamb:2022:MRF

[LBB22] N. Lamb, S. Banerjee, and N. K. Banerjee. MendNet: Restoration of fractured shapes using learned occupancy functions. Computer Graphics Forum, 41(5):65–78, August 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

LaRosa:2023:SAV

[LBB^{+}23] B. La Rosa, G. Blasilli, R. Bourqui, D. Auber, G. Santucci, R. Capobianco, E. Bertini, R. Giot, and M. Angelini. State of the art of visual analytics for eXplainable deep learning. Computer Graphics Forum, 42(1):319–355, February 2023. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lemonari:2022:AVC

[LBC^{+}22] Marilena Lemonari, Rafael Blanco, Panayiotis Charalambous, Nuria Pelechano, Marios Avraamides, Julien Pettré, and Yiorgos Chrysanthou. Authoring virtual crowds: a survey. Computer Graphics Forum, 41(2):677–701, May 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lavoue:2021:PQB

[LBFS21] Guillaume Lavoué, Nicolas Bonneel, Jean-Philippe Farrugia, and Cyril Soler. Perceptual quality of BRDF approximations: dataset and metrics. Computer Graphics Forum, 40(2):327–338, May 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lyon:2020:CML

[LBK20] M. Lyon, D. Bommes, and L. Kobbelt. Cost minimizing local anisotropic quad mesh refinement. Computer Graphics Forum, 39(5):163–172, August 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lee:2021:IIM

[LBKP21] Yonghyeon Lee, Jonghyuk Baek, Young Min Kim, and Frank Chongwoo Park. IMAT: The iterative medial axis transform. Computer Graphics Forum, 40(6):162–181, September 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Latif:2021:DUV

[LCB21] Shahid Latif, Siming Chen, and Fabian Beck. A deeper understanding of visualization-text interplay in geographic data-driven stories. Computer Graphics Forum, 40(3):311–322, June 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lejemble:2021:SED

[LCBM21] T. Lejemble, D. Coeurjolly, L. Barthe, and N. Mellado. Stable and efficient differential estimators on oriented point clouds. Computer Graphics Forum, 40(5):205–216, August 2021. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lai:2020:FSS

[LCG^{+}20] Junyu Lai, Yangang Chen, Yu Gu, Christopher Batty, and Justin W. L. Wan. Fast and scalable solvers for the fluid pressure equations with separating solid boundary conditions. Computer Graphics Forum, 39(2):23–33, May 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lu:2021:CCH

[LCI^{+}21] Yucheng Lu, Luyu Cheng, Tobias Isenberg, Chi-Wing Fu, Guoning Chen, Hui Liu, Oliver Deussen, and Yunhai Wang. Curve complexity heuristic KD-trees for neighborhood-based exploration of 3D curves. Computer Graphics Forum, 40(2):461–474, May 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lyon:2021:QLC

[LCK21a] M. Lyon, M. Campen, and L. Kobbelt. Quad layouts via constrained T-mesh quantization. Computer Graphics Forum, 40(2): 305–314, May 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lyon:2021:SQL

[LCK21b] M. Lyon, M. Campen, and L. Kobbelt. Simpler quad layouts using relaxed singularities. Computer Graphics Forum, 40(5): 169–180, August 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liu:2020:SID

[LCO20] Zhengliang Liu, R. Jordan Crouser, and Alvitta Ottley. Survey on individual differences in visualization. Computer Graphics Forum, 39(3):693–712, June 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liu:2023:UTW

[LCP^{+}23] Mengya Liu, Ajad Chhatkuli, Janis Postels, Luc Van Gool, and Federico Tombari. Unsupervised template warp consistency for implicit surface correspondences. Computer Graphics Forum, 42(2):77–87, May 2023. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Levi:2022:SPS

[Lev22] Zohar Levi. Seamless parametrization of spheres with controlled singularities. Computer Graphics Forum, 41(1):57–68, February 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Luongo:2020:BRM

[LFD^{+}20] A. Luongo, V. Falster, M. B. Doest, M. M. Ribo, E. R. Eiriksson, D. B. Pedersen, and J. R. Frisvad. Book review: Microstructure Control in 3D Printing with Digital Light Processing. Computer Graphics Forum, 39(1):347–359, February 2020. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Li:2022:DAP

[LFH^{+}22] Tingting Li, Yunfei Fu, Xiaoguang Han, Hui Liang, Jian Jun Zhang, and Jian Chang. DiffusionPointLabel: Annotated point cloud generation with diffusion model. Computer Graphics Forum, 41(7):131–139, October 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liu:2023:LLS

[LFR23] Chen Liu, Michael Fischer, and Tobias Ritschel. Learning to learn and sample BRDFs. Computer Graphics Forum, 42(2): 201–211, May 2023. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Licorish:2021:ACN

[LFS21] C. Licorish, N. Faraj, and B. Summa. Adaptive compositing and navigation of variable resolution images. Computer Graphics Forum, 40(1):138–150, February 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lian:2022:CSC

[LG22] Zhouhui Lian and Yichen Gao. CVFont: Synthesizing Chinese vector fonts via deep layout inferring. Computer Graphics Forum, 41(6):212–225, September 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lo:2022:MVW

[LGS^{+}22] Leo Yu-Ho Lo, Ayush Gupta, Kento Shigyo, Aoyu Wu, Enrico Bertini, and Huamin Qu. Misinformed by visualization: What do we learn from misinformative visualizations? Computer Graphics Forum, 41(3):515–525, June 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Li:2022:IED

[LGW^{+}22] Rui-Zeng Li, Jia-Peng Guo, Qi Wang, Shuangming Chai, Ligang Liu, and Xiao-Ming Fu. Interactive editing of discrete Chebyshev nets. Computer Graphics Forum, 41(2):111–120, May 2022. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Luvizon:2023:SAM

[LHG^{+}23] D. C. Luvizon, M. Habermann, V. Golyanik, A. Kortylewski, and C. Theobalt. Scene-aware 3D multi-human motion capture from a single camera. Computer Graphics Forum, 42(2):371–383, May 2023. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liu:2021:RLE

[LHJ21] Hongli Liu, Honglei Han, and Min Jiang. Rank-1 lattices for efficient path integral estimation. Computer Graphics Forum, 40(2):91–102, May 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Li:2022:MHR

[LHJ^{+}22] Yuan Li, Xiangyang He, Yankai Jiang, Huan Liu, Yubo Tao, and Lin Hai. MeshFormer: High-resolution mesh segmentation with graph transformer. Computer Graphics Forum, 41(7):37– 49, October 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Li:2021:ASM

[LHL^{+}21] Qinsong Li, Ling Hu, Shengjun Liu, Dangfu Yang, and Xinru Liu. Anisotropic spectral manifold wavelet descriptor. Computer Graphics Forum, 40(1):81–96, February 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Leonard:2021:LMS

[LHW21] L. Leonard, K. Höhlein, and R. Westermann. Learning multiplescattering solutions for sphere-tracing of volumetric subsurface effects. Computer Graphics Forum, 40(2):165–178, May 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liu:2022:WCR

[LHX^{+}22] Zhenyuan Liu, Jingyu Hu, Hao Xu, Peng Song, Ran Zhang, Bernd Bickel, and Chi-Wing Fu. Worst-case rigidity analysis and optimization for assemblies with mechanical joints. Computer Graphics Forum, 41(2):507–519, May 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Li:2022:IND

[LHZ^{+}22] Moran Li, Haibin Huang, Yi Zheng, Mengtian Li, Nong Sang, and Chongyang Ma. Implicit neural deformation for sparse-view face reconstruction. Computer Graphics Forum, 41(7):601–610, October 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liu:2021:NDS

[LJ21] Hsueh-Ti Derek Liu and Alec Jacobson. Normal-driven spherical shape analogies. Computer Graphics Forum, 40(5):45–55, August 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liu:2022:TSP

[LJJ22] Yang Liu, Shaojie Jiao, and Wojciech Jarosz. Temporally sliced photon primitives for time-of-flight rendering. Computer Graphics Forum, 41(4):29–40, July 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lu:2021:CNR

[LJLB21] Y. Lu, K. Jiang, J. A. Levine, and M. Berger. Compressive neural representations of volumetric scalar fields. Computer Graphics Forum, 40(3):135–146, June 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lee:2021:HRO

[LJSL21] Gi Beom Lee, Moonsoo Jeong, Yechan Seok, and Sungkil Lee. Hierarchical raster occlusion culling. Computer Graphics Forum, 40(2):489–495, May 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lichtenberg:2020:PFE

[LL20] N. Lichtenberg and K. Lawonn. Parameterization, feature extraction and binary encoding for the visualization of tree-like structures. Computer Graphics Forum, 39(1):497–510, February 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Li:2020:SSE

[LLC^{+}20] Q. Li, Y. J. Liu, L. Chen, X. C. Yang, Y. Peng, X. R. Yuan, and M. L. L. Wijerathne. SEEVis: A smart emergency evacuation plan visualization system with data-driven shot designs. Computer Graphics Forum, 39(3):523–535, June 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liu:2020:BRR

[LLDT20] S-L. Liu, Y. Liu, L-F. Dong, and X. Tong. Book review: RAS: A Data-Driven Rigidity-Aware Skinning Model For 3D Facial Animation. Computer Graphics Forum, 39(1):581–594, February 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Loschner:2020:HOT

[LLJ^{+}20] Fabian Löschner, Andreas Longva, Stefan Jeske, Tassilo Kugelstadt, and Jan Bender. Higher-order time integration for deformable solids. Computer Graphics Forum, 39(8):157–169, December 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lai:2021:VCM

[LLL^{+}21] G. Lai, F. F. Leymarie, W. Latham, T. Arita, and R. Suzuki. Virtual creature morphology — a review. Computer Graphics Forum, 40(2):659–681, May 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liu:2022:FMS

[LLL^{+}22] Xiaokang Liu, Chenran Li, Lin Lu, Oliver Deussen, and Changhe Tu. Fabricable multi-scale Wang tiles. Computer Graphics Forum, 41(5):149–159, August 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Leon:2022:MMC

[LLLB22] G. Molina León, M. Lischka, W. Luo, and A. Breiter. Mobile and multimodal? A comparative evaluation of interactive workplaces for visual data exploration. Computer Graphics Forum, 41(3): 417–428, June 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lescoat:2020:SMS

[LLT^{+}20a] Thibault Lescoat, Hsueh-Ti Derek Liu, Jean-Marc Thiery, Alec Jacobson, Tamy Boubekeur, and Maks Ovsjanikov. Spectral mesh simplification. Computer Graphics Forum, 39(2):315–324, May 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Li:2020:WVV

[LLT^{+}20b] Q. Li, Q. Q. Liu, C. F. Tang, Z. W. Li, S. C. Wei, X. R. Peng, M. H. Zheng, T. J. Chen, and Q. Yang. Warehouse vis: A visual analytics approach to facilitating warehouse location selection for business districts. Computer Graphics Forum, 39(3):483–495, June 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liu:2020:PFD

[LLZ^{+}20] Hao-Yu Liu, Zhong-Yuan Liu, Zheng-Yu Zhao, Ligang Liu, and Xiao-Ming Fu. Practical fabrication of discrete Chebyshev nets. Computer Graphics Forum, 39(7):13–26, October 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lejemble:2020:PAM

[LMBM20] T. Lejemble, C. Mura, L. Barthe, and N. Mellado. Persistence analysis of multi-scale planar structure graph in point clouds. Computer Graphics Forum, 39(2):35–50, May 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Laccone:2020:ADC

[LMF^{+}20] Francesco Laccone, Luigi Malomo, Maurizio Froli, Paolo Cignoni, and Nico Pietroni. Automatic design of cable-tensioned glass shells. Computer Graphics Forum, 39(1):260–273, February 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Li:2020:IPL

[LNRH20] Z. Li, G. Nawratil, F. Rist, and M. Hensel. Invertible paradoxic loop structures for transformable design. Computer Graphics Forum, 39(2):261–275, May 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lekschas:2020:PEI

[LPH^{+}20] Fritz Lekschas, Brant Peterson, Daniel Haehn, Eric Ma, Nils Gehlenborg, and Hanspeter Pfister. Peax: Interactive visual pattern search in sequential data using unsupervised deep representation learning. Computer Graphics Forum, 39(3):167–179, June 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liem:2020:SEV

[LPW20a] J. Liem, C. Perin, and J. Wood. Structure and empathy in visual data storytelling: Evaluating their influence on attitude. Computer Graphics Forum, 39(3):277–289, June 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lu:2020:BRI

[LPW^{+}20b] Feixiang Lu, Haotian Peng, Hongyu Wu, Jun Yang, Xinhang Yang, Ruizhi Cao, Liangjun Zhang, Ruigang Yang, and Bin Zhou. Book review: InstanceFusion: Real-time Instance-level 3D Reconstruction Using a Single RGBD Camera. Computer Graphics Forum, 39(7):433–445, October 2020. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lachaud:2020:ICC

[LRTC20] J.-O. Lachaud, P. Romon, B. Thibert, and D. Coeurjolly. Interpolated corrected curvature measures for polygonal surfaces. Computer Graphics Forum, 39(5):41–54, August 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lutz:2021:CGN

[LSD21] Nicolas Lutz, Basile Sauvage, and Jean-Michel Dischler. Cyclostationary Gaussian noise: theory and synthesis. Computer Graphics Forum, 40(2):239–250, May 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lutz:2023:PAT

[LSD23] Nicolas Lutz, Basile Sauvage, and Jean-Michel Dischler. Preserving the autocovariance of texture tilings using importance sampling. Computer Graphics Forum, 42(2):347–358, May 2023. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Luo:2021:TBF

[LSE21] Xuejiao Luo, Leonardo Scandolo, and Elmar Eisemann. Texture browser: Feature-based texture exploration. Computer Graphics Forum, 40(3):99–109, June 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Li:2021:MAN

[LSK21] Tianxing Li, Rui Shi, and Takashi Kanai. MultiResGNet: Approximating nonlinear deformation via multi-resolution graphs. Computer Graphics Forum, 40(2):537–548, May 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Li:2023:DAD

[LSK23] T. Li, R. Shi, and T. Kanai. Detail-aware deep clothing animations infused with multi-source attributes. Computer Graphics Forum, 42(1):231–244, February 2023. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lengauer:2021:BDR

[LSP^{+}21] Stefan Lengauer, Ivan Sipiran, Reinhold Preiner, Tobias Schreck, and Benjamin Bustos. A benchmark dataset for repetitive pattern recognition on textured 3D surfaces. Computer Graphics Forum, 40(5):1–8, August 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Luo:2022:RSS

[LSSW22] Shoutong Luo, Zhengxing Sun, Yunhan Sun, and Yi Wang. Resolution-switchable 3D semantic scene completion. Computer Graphics Forum, 41(7):121–130, October 2022. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lin:2021:HAH

[LSY21] D. Lin, L. Seiler, and C. Yuksel. Hardware adaptive high-order interpolation for real-time graphics. Computer Graphics Forum, 40(8):1–16, December 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liu:2020:HIS

[LTDG20] Songrun Liu, Jianchao Tan, Zhigang Deng, and Yotam Gingold. Hyperspectral inverse skinning. Computer Graphics Forum, 39 (6):49–65, September 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lobo:2020:FDC

[LTH20] M-J. Lobo, A. C. Telea, and C. Hurter. Feature driven combination of animated vector field visualizations. Computer Graphics Forum, 39(3):429–441, June 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Li:2022:CPC

[LTY^{+}22] Y. D. Li, M. Tang, Y. Yang, Z. Huang, R. F. Tong, S. C. Yang, Y. Li, and D. Manocha. N-Cloth: Predicting 3D cloth deformation with mesh-based networks. Computer Graphics Forum, 41(2):547–558, May 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liu:2021:TDS

[LWB^{+}21] Sinuo Liu, Xiaokun Wang, Xiaojuan Ban, Yanrui Xu, Jing Zhou, Jiří Kosinka, and Alexandru C. Telea. Turbulent details simulation for SPH fluids via vorticity refinement. Computer Graphics Forum, 40(1):54–67, February 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Li:2020:ABL

[LWH^{+}20] Shi Li, Rui Wang, Yuchi Huo, Wenting Zheng, Wei Hua, and Hujun Bao. Automatic band-limited approximation of shaders using mean-variance statistics in clamped domain. Computer Graphics Forum, 39(7):181–192, October 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lohfink:2020:FCT

[LWL^{+}20] Anna-Pia Lohfink, Florian Wetzels, Jonas Lukasczyk, Gunther H. Weber, and Christoph Garth. Fuzzy contour trees: Alignment and joint layout of multiple contour trees. Computer Graphics Forum, 39(3):343–355, June 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lin:2021:PBM

[LWY^{+}21] Weiheng Lin, Beibei Wang, Jian Yang, Lu Wang, and Ling-Qi Yan. Path-based Monte Carlo denoising using a three-scale neural network. Computer Graphics Forum, 40(1):369–381, February 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Li:2020:TSP

[LWZ^{+}20] Simin Li, Qiang Wen, Shuang Zhao, Zixun Sun, and Shengfeng He. Two-stage photograph cartoonization via line tracing. Computer Graphics Forum, 39(7):587–599, October 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liu:2022:WLE

[LXY^{+}22] Shengjun Liu, Haojun Xu, Dong-Ming Yan, Ling Hu, Xinru Liu, and Qinsong Li. WTFM layer: an effective map extractor for unsupervised shape correspondence. Computer Graphics Forum, 41(7):51–61, October 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Li:2020:RCI

[LXYM20] Jianping Kelvin Li, Shenyu Xu, Yecong (Chris) Ye, and Kwan-Liu Ma. Resolving conflicting insights in asynchronous collaborative visual analysis. Computer Graphics Forum, 39(3):497–509, June 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lin:2020:MPS

[LY20] Jie Lin and Chuan-Kai Yang. A multi-person selfie system via augmented reality. Computer Graphics Forum, 39(7):553–564, October 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Lan:2021:VAD

[LYA^{+}21] Fangfei Lan, Michael Young, Lauren Anderson, Anders Ynnerman, Alexander Bock, Michelle A. Borkin, Angus G. Forbes, Juna A. Kollmeier, and Bei Wang. Visualization in astrophysics: Developing new methods, discovering our universe, and educating the Earth. Computer Graphics Forum, 40(3):635–663, June 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Liu:2022:FBK

[LZ22] Huan Liu and Hui Zhang. A flip-book of knot diagrams for visualizing surfaces in 4-space. Computer Graphics Forum, 41(3): 345–354, June 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Luan:2021:USS

[LZBD21] Fujun Luan, Shuang Zhao, Kavita Bala, and Zhao Dong. Unified shape and SVBRDF recovery using differentiable Monte Carlo rendering. Computer Graphics Forum, 40(4):101–113, July 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

May:2023:COI

[MA23] C. May and D. Aliaga. CubeGAN: Omnidirectional image synthesis using generative adversarial networks. Computer Graphics Forum, 42(2):213–224, May 2023. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Macedo:2020:RBS

[MAA20] M. C. F. Macedo, A. L. Apolinário, Jr., and K. A. Agüero. Revectorization-based soft shadow mapping. Computer Graphics Forum, 39(1):389–404, February 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Morais:2022:EHV

[MAS22] L. Morais, N. Andrade, and D. Sousa. Exploring how visualization design and situatedness evoke compassion in the wild. Computer Graphics Forum, 41(3):441–452, June 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Michel:2020:RTM

[MB20] É. Michel and T. Boubekeur. Real time multiscale rendering of dense dynamic stackings. Computer Graphics Forum, 39(7):169– 179, October 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mathew:2022:SVC

[MBA22] C. D. T. Mathew, B. Benes, and D. G. Aliaga. Sketching vocabulary for crowd motion. Computer Graphics Forum, 41(8): 119–130, December 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Marques:2022:GPR

[MBB22] R. Marques, C. Bouville, and K. Bouatouch. Gaussian process for radiance functions on the S^{2} sphere. Computer Graphics Forum, 41(6):67–81, September 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mariani:2020:GAS

[MCBR20] G. Mariani, L. Cosmo, A. M. Bronstein, and E. Rodolà. Generating adversarial surfaces via band-limited perturbations. Computer Graphics Forum, 39(5):253–264, August 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mandal:2023:RFG

[MCC23] A. Mandal, P. Chaudhuri, and S. Chaudhuri. Remeshing-free graph-based finite element method for fracture simulation. Computer Graphics Forum, 42(1):117–134, February 2023. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mahajan:2022:VDS

[MCK^{+}22] Shambhavi Mahajan, Bonnie Chen, Alireza Karduni, Yea-Seul Kim, and Emily Wall. VIBE: a design space for VIsual Belief Elicitation in data journalism. Computer Graphics Forum, 41 (3):477–488, June 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

McNutt:2021:WTC

[McN21] A. McNutt. What are table cartograms good for anyway? An algebraic analysis. Computer Graphics Forum, 40(3):61–73, June 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Maggiordomo:2021:TDP

[MCT21] Andrea Maggiordomo, Paolo Cignoni, and Marco Tarini. Texture defragmentation for photo-reconstructed 3D models. Computer Graphics Forum, 40(2):65–78, May 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mikula:2022:IAI

[MDBH22] Natalia Mikula, Tom Dörffel, Daniel Baum, and Hans-Christian Hege. An interactive approach for identifying structure definitions. Computer Graphics Forum, 41(3):321–332, June 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mezieres:2022:HVL

[MDVP22] Pierre Mézières, François Desrichard, David Vanderhaeghe, and Mathias Paulin. Harmonics virtual lights: Fast projection of luminance field on spherical harmonics for efficient rendering. Computer Graphics Forum, 41(6):182–195, September 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Molenaar:2023:ECH

[ME23] M. Molenaar and E. Eisemann. Editing compressed highresolution voxel scenes with attributes. Computer Graphics Forum, 42(2):235–243, May 2023. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Macklin:2020:PDD

[MEM^{+}20] M. Macklin, K. Erleben, M. Müller, N. Chentanez, S. Jeschke, and T. Y. Kim. Primal/Dual descent methods for dynamics. Computer Graphics Forum, 39(8):89–100, December 2020. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Miller:2022:ADS

[MFH^{+}22] Matthias Miller, Daniel Fürst, Hanna Hauptmann, Daniel A. Keim, and Mennatallah El-Assady. Augmenting digital sheet music through visual analytics. Computer Graphics Forum, 41 (1):301–316, February 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Modi:2021:EEM

[MFJ^{+}21] V. Modi, L. Fulton, A. Jacobson, S. Sueda, and D. I. W. Levin. EMU: Efficient muscle simulation in deformation space. Computer Graphics Forum, 40(1):234–248, February 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Munkberg:2020:NDL

[MH20] Jacob Munkberg and Jon Hasselgren. Neural denoising with layer embeddings. Computer Graphics Forum, 39(4):1–12, July 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Metzer:2022:ZIV

[MHG^{+}22] G. Metzer, R. Hanocka, R. Giryes, N. J. Mitra, and D. Cohen-Or. Z2P: Instant visualization of point clouds. Computer Graphics Forum, 41(2):461–471, May 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mourot:2022:SDL

[MHL^{+}22] Lucas Mourot, Ludovic Hoyet, François Le Clerc, François Schnitzler, and Pierre Hellier. A survey on deep learning for skeleton-based human animation. Computer Graphics Forum, 41(1):122–157, February 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mourot:2022:UDL

[MHLH22] Lucas Mourot, Ludovic Hoyet, François Le Clerc, and Pierre Hellier. UnderPressure: Deep learning for foot contact detection, ground reaction force estimation and footskate cleanup. Computer Graphics Forum, 41(8):195–206, December 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Meidiana:2020:STF

[MHT^{+}20] Amyra Meidiana, Seok-Hee Hong, Marnijati Torkel, Shijun Cai, and Peter Eades. Sublinear time force computation for big complex network visualization. Computer Graphics Forum, 39(3): 579–591, June 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mitchell:2022:DLT

[Mit22] Scott A. Mitchell. Deterministic linear time for maximal Poissondisk sampling using chocks without rejection or approximation. Computer Graphics Forum, 41(5):101–111, August 2022. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Matute:2020:HSC

[ML20] J. Matute and L. Linsen. Hinted star coordinates for mixed data. Computer Graphics Forum, 39(1):117–133, February 2020. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Matute:2022:EDT

[ML22] José Matute and Lars Linsen. Evaluating data-type heterogeneity in interactive visual analyses with parallel axes. Computer Graphics Forum, 41(1):335–349, February 2022. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic). See erratum [Ano22b].

Mayer:2021:VVE

[MLD^{+}21] B. Mayer, K. Lawonn, K. Donnay, B. Preim, and M. Meuschke. VEHICLE: Validation and exploration of the hierarchical integration of conflict event data. Computer Graphics Forum, 40(3): 1–12, June 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mancinelli:2021:PCC

[MLP21] C. Mancinelli, M. Livesu, and E. Puppo. Practical computation of the cut locus on discrete surfaces. Computer Graphics Forum, 40(5):261–273, August 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mezieres:2022:RAS

[MMBP22] P. Mézières, N. Mellado, L. Barthe, and M. Paulin. Recursive analytic spherical harmonics gradient for spherical lights. Computer Graphics Forum, 41(2):393–406, May 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Muller:2020:DRB

[MMC^{+}20] Matthias Müller, Miles Macklin, Nuttapong Chentanez, Stefan Jeschke, and Tae-Yong Kim. Detailed rigid body simulation with extended position based dynamics. Computer Graphics Forum, 39(8):101–112, December 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Moschella:2022:LSU

[MMC^{+}22] Luca Moschella, Simone Melzi, Luca Cosmo, Filippo Maggioli, Or Litany, Maks Ovsjanikov, Leonidas Guibas, and Emanuele Rodolà. Learning spectral unions of partial deformable 3D shapes. Computer Graphics Forum, 41(2):407–417, May 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Muller:2022:PBS

[MMCJ22] Matthias Müller, Miles Macklin, Nuttapong Chentanez, and Stefan Jeschke. Physically based shape matching. Computer Graphics Forum, 41(8):1–7, December 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mukherjee:2021:NST

[MMH^{+}21] Sabyasachi Mukherjee, Sayan Mukherjee, Binh-Son Hua, Nobuyuki Umetani, and Daniel Meister. Neural sequence transformation. Computer Graphics Forum, 40(7):131–140, October 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Majercik:2022:DDG

[MMK^{+}22] Zander Majercik, Thomas Müller, Alexander Keller, Derek Nowrouzezahrai, and Morgan McGuire. Dynamic diffuse global illumination resampling. Computer Graphics Forum, 41(1):158– 171, February 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Maggioli:2022:MMM

[MMMR22] F. Maggioli, R. Marin, S. Melzi, and E. Rodolà. MoMaS: Mold manifold simulation for real-time procedural texturing. Computer Graphics Forum, 41(7):519–527, October 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Maggioli:2021:OFP

[MMO^{+}21] F. Maggioli, S. Melzi, M. Ovsjanikov, M. M. Bronstein, and E. Rodolà. Orthogonalized Fourier polynomials for signal approximation and transfer. Computer Graphics Forum, 40(2):435–447, May 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mas:2020:SEA

[MMP20] Albert Mas, Ignacio Martin, and Gustavo Patow. Simulating the evolution of ancient fortified cities. Computer Graphics Forum, 39(1):650–671, February 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Marin:2020:BRF

[MMRC20] R. Marin, S. Melzi, E. Rodolà, and U. Castellani. Book review: FARM: Functional Automatic Registration Method for 3D Human Bodies. Computer Graphics Forum, 39(1):160–173, February 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mercier:2020:LBM

[MN20] Olivier Mercier and Derek Nowrouzezahrai. Local bases for model-reduced smoke simulations. Computer Graphics Forum, 39 (2):9–22, May 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Magnet:2023:SEF

[MO23] Robin Magnet and Maks Ovsjanikov. Scalable and efficient functional map computations on dense meshes. Computer Graphics Forum, 42(2):89–101, May 2023. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Meister:2021:SBV

[MOB^{+}21] Daniel Meister, Shinji Ogaki, Carsten Benthin, Michael J. Doyle, Michael Guthe, and Jiří Bittner. A survey on bounding volume hierarchies for ray tracing. Computer Graphics Forum, 40(2): 683–712, May 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Marin:2022:SCR

[MOW22] D. Marin, S. Ohrhallinger, and M. Wimmer. SIGDT: 2D curve reconstruction. Computer Graphics Forum, 41(7):25–36, October 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Matusovic:2023:IDP

[MPE23] Marko Matusovic, Amal Dev Parakkat, and Elmar Eisemann. Interactive depixelization of pixel art through spring simulation. Computer Graphics Forum, 42(2):51–60, May 2023. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mura:2021:WEF

[MPSM21] Claudio Mura, Renato Pajarola, Konrad Schindler, and Niloy Mitra. Walk2Map: Extracting floor plans from indoor walk trajectories. Computer Graphics Forum, 40(2):375–388, May 2021.

CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Marschner:2020:HMR

[MPZS20] Z. Marschner, D. Palmer, P. Zhang, and J. Solomon. Hexahedral mesh repair via sum-of-squares relaxation. Computer Graphics Forum, 39(5):133–147, August 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mistelbauer:2021:IMP

[MRB^{+}21] G. Mistelbauer, C. Rössl, K. Bäumler, B. Preim, and D. Fleischmann. Implicit modeling of patient-specific aortic dissections with elliptic Fourier descriptors. Computer Graphics Forum, 40 (3):423–434, June 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mitchel:2021:EEC

[MRCK21] Thomas W. Mitchel, Szymon Rusinkiewicz, Gregory S. Chirikjian, and Michael Kazhdan. ECHO: Extended convolution histogram of orientations for local surface description. Computer Graphics Forum, 40(1):180–194, February 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Miller:2022:CVA

[MRKEA22] Matthias Miller, Julius Rauscher, Daniel A. Keim, and Mennatallah El-Assady. CorpusVis: Visual analysis of digital sheet music collections. Computer Graphics Forum, 41(3):283–294, June 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Martin:2022:MSI

[MRR^{+}22] Rosalie Martin, Arthur Roullier, Romain Rouffet, Adrien Kaiser, and Tamy Boubekeur. MaterIA: Single image high-resolution material capture in the wild. Computer Graphics Forum, 41(2): 163–177, May 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mueller-Roemer:2020:ASL

[MRSF20] J. S. Mueller-Roemer, A. Stork, and D. Fellner. Analysis of schedule and layout tuning for sparse matrices with compound entries on GPUs. Computer Graphics Forum, 39(6):133–143, September 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Morgenroth:2020:BRE

[MRWE20] D. Morgenroth, S. Reinhardt, D. Weiskopf, and B. Eberhardt. Book review: Efficient 2D Simulation on Moving 3D Surfaces. Computer Graphics Forum, 39(8):27–38, December 2020. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Milef:2023:VPP

[MSK23] N. Milef, S. Sueda, and N. Khademi Kalantari. Variational pose prediction with dynamic sample selection from sparse tracking signals. Computer Graphics Forum, 42(2):359–369, May 2023. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

McDonald:2021:LTE

[MSY^{+}21] T. McDonald, R. Shrestha, X. Yi, H. Bhatia, D. Chen, D. Goswami, V. Pascucci, T. Turbyville, and P.-T. Bremer. Leveraging topological events in tracking graphs for understanding particle diffusion. Computer Graphics Forum, 40(3):251–262, June 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mokady:2022:JJK

[MTB^{+}22] R. Mokady, R. Tzaban, S. Benaim, A. H. Bermano, and D. Cohen-Or. JOKR: Joint keypoint representation for unsupervised video retargeting. Computer Graphics Forum, 41(6): 245–257, September 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Masood:2021:VAE

[MTL^{+}21] T. Bin Masood, S. S. Thygesen, M. Linares, A. I. Abrikosov, V. Natarajan, and I. Hotz. Visual analysis of electronic densities and transitions in molecules. Computer Graphics Forum, 40(3): 287–298, June 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Meng:2022:MIM

[MvdEV22] Linhao Meng, Stef van den Elzen, and Anna Vilanova. ModelWise: Interactive model comparison for model diagnosis, improvement and selection. Computer Graphics Forum, 41(3): 97–108, June 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mestres:2021:LLA

[MVNT21] Nolan Mestres, Romain Vergne, Camille Noûs, and Joëlle Thollot. Local light alignment for multi-scale shape depiction. Computer Graphics Forum, 40(2):575–584, May 2021. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Morth:2020:RIV

[MWLH^{+}20] E. Mörth, K. Wagner-Larsen, E. Hodneland, C. Krakstad, I. S. Haldorsen, S. Bruckner, and N. N. Smit. RadEx: Integrated visual exploration of multiparametric studies for radiomic tumor profiling. Computer Graphics Forum, 39(7):611–622, October 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mao:2022:FCS

[MWM^{+}22] Tianlu Mao, Ji Wang, Ruoyu Meng, Qinyuan Yan, Shaohua Liu, and Zhaoqi Wang. A fusion crowd simulation method: Integrating data with dynamics, personality with common. Computer Graphics Forum, 41(8):131–142, December 2022. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ma:2020:DRC

[MWS^{+}20a] Z. Ma, A. Walzer, C. Schumacher, R. Rust, F. Gramazio, M. Kohler, and M. Bächer. Designing robotically-constructed metal frame structures. Computer Graphics Forum, 39(2):411– 422, May 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Mlakar:2020:SSL

[MWS^{+}20b] D. Mlakar, M. Winter, P. Stadlbauer, H.-P. Seidel, M. Steinberger, and R. Zayer. Subdivision-specialized linear algebra kernels for static and dynamic mesh connectivity on the GPU. Computer Graphics Forum, 39(2):335–349, May 2020. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ma:2020:SNS

[MWZ^{+}20] Qian Ma, Guangshun Wei, Yuanfeng Zhou, Xiao Pan, Shiqing Xin, and Wenping Wang. SRF-Net: Spatial relationship feature network for tooth point cloud classification. Computer Graphics Forum, 39(7):267–277, October 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ma:2021:LEM

[MYT^{+}21] Li-Ke Ma, Zeshi Yang, Xin Tong, Baining Guo, and KangKang Yin. Learning and exploring motor skills with spacetime bounds. Computer Graphics Forum, 40(2):251–263, May 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ma:2020:GIL

[MZAD^{+}20] Chao Ma, Ye Zhao, Shamal Al-Dohuki, Jing Yang, Xinyue Ye, Farah Kamw, and Md Amiruzzaman. GTMapLens: Interactive lens for geo-text data browsing on map. Computer Graphics Forum, 39(3):469–481, June 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Narita:2022:TCM

[NA22] F. Narita and R. Ando. Tiled characteristic maps for tracking detailed liquid surfaces. Computer Graphics Forum, 41(8): 231–242, December 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Nasikun:2020:LST

[NBH20] Ahmad Nasikun, Christopher Brandt, and Klaus Hildebrandt. Locally supported tangential vector, n-vector, and tensor fields. Computer Graphics Forum, 39(2):203–217, May 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Nie:2020:KAC

[NBPM20] K. Nie, P. Baltzer, B. Preim, and G. Mistelbauer. Knowledgeassisted comparative assessment of breast cancer using dynamic contrast-enhanced magnetic resonance imaging. Computer Graphics Forum, 39(3):13–23, June 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Nardini:2021:AIC

[NCB^{+}21] Pascal Nardini, Min Chen, Michael Böttinger, Gerik Scheuermann, and Roxana Bujack. Automatic improvement of continuous colormaps in Euclidean colorspaces. Computer Graphics Forum, 40(3):361–373, June 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Nie:2021:FRE

[NHSS21] Xiaoying Nie, Yong Hu, Zhiyuan Su, and Xukun Shen. Fluid reconstruction and editing from a monocular video based on the SPH model with external force guidance. Computer Graphics Forum, 40(6):62–76, September 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Naruniec:2020:HRN

[NHSW20] J. Naruniec, L. Helminger, C. Schroers, and R. M. Weber. Highresolution neural face swapping for visual effects. Computer Graphics Forum, 39(4):173–184, July 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Nabata:2020:TSR

[NID20] K. Nabata, K. Iwasaki, and Y. Dobashi. Two-stage resampling for bidirectional path tracing with multiple light sub-paths. Computer Graphics Forum, 39(7):219–230, October 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Nyatsanga:2023:CRD

[NKA^{+}23] S. Nyatsanga, T. Kucherenko, C. Ahuja, G. E. Henter, and M. Neff. A comprehensive review of data-driven co-speech gesture generation. Computer Graphics Forum, 42(2):569–596, May 2023. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Nysjo:2020:RAI

[NMN20] F. Nysjö, F. Malmberg, and I. Nyström. RayCaching: Amortized isosurface rendering for virtual reality. Computer Graphics Forum, 39(1):220–230, February 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Neff:2022:MHS

[NMSS22] T. Neff, J. H. Mueller, M. Steinberger, and D. Schmalstieg. Meshlets and how to shade them: a study on texture-space shading. Computer Graphics Forum, 41(2):277–287, May 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Navarro:2021:SDM

[NODI21] Pablo Navarro, J. Ignacio Orlando, Claudio Delrieux, and Emmanuel Iarussi. SketchZooms: Deep multi-view descriptors for matching line drawings. Computer Graphics Forum, 40(1):410– 423, February 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Nguyen:2020:EEA

[NQHH20] F. Nguyen, X. Qiao, J. Heer, and J. Hullman. Exploring the effects of aggregation choices on untrained visualization users’ generalizations from data. Computer Graphics Forum, 39(6):33– 48, September 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Nir:2022:CCL

[NRS22] Oron Nir, Gal Rapoport, and Ariel Shamir. CAST: Character labeling in animation using self-supervision by tracking. Computer Graphics Forum, 41(2):135–145, May 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Neff:2021:DTR

[NSP^{+}21] T. Neff, P. Stadlbauer, M. Parger, A. Kurz, J. H. Mueller, C. R. A. Chaitanya, A. Kaplanyan, and M. Steinberger. DONeRF: Towards real-time rendering of compact neural radiance fields using depth oracle networks. Computer Graphics Forum, 40(4): 45–59, July 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Nuvoli:2021:ASS

[NTM^{+}21] S. Nuvoli, A. Tola, A. Muntoni, N. Pietroni, E. Gobbetti, and R. Scateni. Automatic surface segmentation for seamless fabrication using 4-axis milling machines. Computer Graphics Forum, 40(2):191–203, May 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Niu:2020:BRW

[NYB^{+}20] Chengjie Niu, Yang Yu, Zhenwei Bian, Jun Li, and Kai Xu. Book review: Weakly Supervised Part-wise 3D Shape Reconstruction from Single-View RGB Images. Computer Graphics Forum, 39 (7):447–457, October 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Noeckel:2021:FAR

[NZCS21] James Noeckel, Haisen Zhao, Brian Curless, and Adriana Schulz. Fabrication-aware reverse engineering for carpentry. Computer Graphics Forum, 40(5):301–314, August 2021. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Nguyen:2021:PBP

[NZL^{+}21] Duong B. Nguyen, Lei Zhang, Robert S. Laramee, David Thompson, Rodolfo Ostilla Monico, and Guoning Chen. Physics-based pathline clustering and exploration. Computer Graphics Forum, 40(1):22–37, February 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Overby:2021:GIG

[OKN21] Matthew Overby, Danny Kaufman, and Rahul Narain. Globally injective geometry optimization with non-injective steps. Computer Graphics Forum, 40(5):111–123, August 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Oh:2021:TST

[OL21] Young Jin Oh and In-Kwon Lee. Two-step temporal interpolation network using forward advection for efficient smoke simulation. Computer Graphics Forum, 40(2):355–365, May 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ouyang:2021:RGP

[OLK^{+}21] Y. Ouyang, S. Liu, M. Kettunen, M. Pharr, and J. Pantaleoni. ReSTIR GI: Path resampling for real-time path tracing. Computer Graphics Forum, 40(8):17–29, December 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Oppermann:2020:OVL

[OM20] Michael Oppermann and Tamara Munzner. Ocupado: Visualizing location-based counts over time across buildings. Computer Graphics Forum, 39(3):127–138, June 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Otto:2022:LDG

[ONH^{+}22] C. Otto, J. Naruniec, L. Helminger, T. Etterlin, G. Mignone, P. Chandran, G. Zoss, C. Schroers, M. Gross, P. Gotardo, D. Bradley, and R. Weber. Learning dynamic 3D geometry and texture for video face swapping. Computer Graphics Forum, 41 (7):611–622, October 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Oliveira:2022:NBS

[OP22] Felipe Oliveira and Afonso Paiva. Narrow-band screen-space fluid rendering. Computer Graphics Forum, 41(6):82–93, September 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ohrhallinger:2021:PCR

[OPP^{+}21] S. Ohrhallinger, J. Peethambaran, A. D. Parakkat, T. K. Dey, and R. Muthuganapathy. 2D points curve reconstruction survey and benchmark. Computer Graphics Forum, 40(2):611–632, May 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ouyang:2020:AAN

[OPY^{+}20] Wenqing Ouyang, Yue Peng, Yuxin Yao, Juyong Zhang, and Bailin Deng. Anderson acceleration for nonconvex ADMM based on Douglas-Rachford splitting. Computer Graphics Forum, 39 (5):221–239, August 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Oztireli:2020:CTV

[Özt20] A. Cengiz Öztireli. A comprehensive theory and variational framework for anti-aliasing sampling patterns. Computer Graphics Forum, 39(4):133–148, July 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Piccolotto:2022:VPS

[PBM^{+}22] N. Piccolotto, M. Bögl, C. Muehlmann, K. Nordhausen, P. Filzmoser, and S. Miksch. Visual parameter selection for spatial blind source separation. Computer Graphics Forum, 41(3):157– 168, June 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ponchio:2020:BRE

[PCDS20] F. Ponchio, M. Callieri, M. Dellepiane, and R. Scopigno. Book review: Effective Annotations Over 3D Models. Computer Graphics Forum, 39(1):89–105, February 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Platkevic:2021:FSA

[PCS21] A. Platkevič, C. Curtis, and D. Sýkora. Fluidymation: Stylizing animations using natural dynamics of artistic media. Computer Graphics Forum, 40(7):21–32, October 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Peters:2021:BIS

[Pet21] Christoph Peters. BRDF importance sampling for linear lights. Computer Graphics Forum, 40(8):31–39, December 2021. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Postolache:2020:PAD

[PFCR20] E. Postolache, M. Fumero, L. Cosmo, and E. Rodolà. A parametric analysis of discrete Hamiltonian functional maps. Computer Graphics Forum, 39(5):103–118, August 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Paris:2021:SGC

[PGP^{+}21] A. Paris, E. Guérin, A. Peytavie, P. Collon, and E. Galin. Synthesizing geologically coherent cave networks. Computer Graphics Forum, 40(7):277–287, October 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Papp:2022:EQC

[PHP22] György Papp, Miklós Hoffmann, and Ildikó Papp. Embedding QR code onto triangulated meshes using horizon based ambient occlusion. Computer Graphics Forum, 41(1):29–45, February 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Panine:2022:NIS

[PKO22] Mikhail Panine, Maxime Kirgo, and Maks Ovsjanikov. Nonisometric shape matching via functional maps on landmarkadapted bases. Computer Graphics Forum, 41(6):394–417, September 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Preim:2020:SVA

[PL20] Bernhard Preim and Kai Lawonn. A survey of visual analytics for public health. Computer Graphics Forum, 39(1):543–580, February 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Peng:2020:FUQ

[PLL^{+}20] Hao Peng, Peiqing Liu, Lin Lu, Andrei Sharf, Lin Liu, Dani Lischinski, and Baoquan Chen. Fabricable unobtrusive 3D-QR-Codes with directional light. Computer Graphics Forum, 39(5):

15–27, August 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pieras:2023:CSC

[PMB^{+}23] M. Pieras, R. Marroquim, D. Broekens, E. Eisemann, and A. Vilanova. ComVis-Sail: Comparative sailing performance visualization for coaching. Computer Graphics Forum, 42(1):86–100, February 2023. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Parakkat:2022:DPP

[PMC22a] Amal Dev Parakkat, Pooran Memari, and Marie-Paule Cani. Delaunay painting: Perceptual image colouring from raster contours with gaps. Computer Graphics Forum, 41(6):166–181, September 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pegoraro:2022:LSM

[PMC^{+}22b] M. Pegoraro, S. Melzi, U. Castellani, R. Marin, and E. Rodolà. Localized shape modelling with global coherence: an inverse spectral approach. Computer Graphics Forum, 41(5):13–24, August 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Perez-Messina:2022:TGT

[PMCEA^{+}22] I. Pérez-Messina, D. Ceneda, M. El-Assady, S. Miksch, and F. Sperrle. A typology of guidance tasks in mixed-initiative visual analytics environments. Computer Graphics Forum, 41(3): 465–476, June 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pintore:2020:BRS

[PMG^{+}20] Giovanni Pintore, Claudio Mura, Fabio Ganovelli, Lizeth Fuentes-Perez, Renato Pajarola, and Enrico Gobbetti. Book review: State-of-the-art in Automatic 3D Reconstruction of Structured Indoor Environments. Computer Graphics Forum, 39(2): 667–699, May 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pister:2023:CVQ

[PPF23] A. Pister, C. Prieur, and J.-D. Fekete. ComBiNet: Visual query and comparison of bipartite multivariate dynamic social networks. Computer Graphics Forum, 42(1):290–304, February 2023. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pizana:2020:BMN

[PRCO20] José M. Pizana, Alejandro Rodríguez, Gabriel Cirio, and Miguel A. Otaduy. A bending model for nodal discretizations of yarn-level cloth. Computer Graphics Forum, 39(8):181–189, December 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Piochowiak:2021:SVR

[PRD21] M. Piochowiak, T. Rapp, and C. Dachsbacher. Stochastic volume rendering of multi-phase SPH data. Computer Graphics Forum, 40(1):97–109, February 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Preiner:2020:ANL

[PSK^{+}20] R. Preiner, J. Schmidt, K. Krösl, T. Schreck, and G. Mistelbauer. Augmenting node-link diagrams with topographic attribute maps. Computer Graphics Forum, 39(3):369–381, June 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Palmer:2021:FFO

[PSS21] D. Palmer, O. Stein, and J. Solomon. Frame field operators. Computer Graphics Forum, 40(5):231–245, August 2021. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Probst:2023:MFC

[PT23] T. Probst and M. Teschner. Monolithic friction and contact handling for rigid bodies and fluids using SPH. Computer Graphics Forum, 42(1):155–179, February 2023. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pan:2021:DDS

[PWBD21] Junjun Pan, Siyuan Wang, Junxuan Bai, and Ju Dai. Diverse dance synthesis via keyframes with transformer controllers. Computer Graphics Forum, 40(7):71–83, October 2021. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ponton:2022:CMM

[PYAP22] J. L. Ponton, H. Yun, C. Andujar, and N. Pelechano. Combining motion matching and orientation prediction to animate avatars for consumer-grade VR devices. Computer Graphics Forum, 41 (8):107–118, December 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Park:2020:MSF

[PYL20] Hwangpil Park, Ri Yu, and Jehee Lee. Multi-segment foot for human modelling and simulation. Computer Graphics Forum, 39 (1):637–649, February 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Pan:2022:PEL

[PZL^{+}22] Haoran Pan, Jun Zhou, Yuanpeng Liu, Xuequan Lu, Weiming Wang, Xuefeng Yan, and Mingqiang Wei. SO(3)-Pose: SO(3)- equivarianceg learning for 6D object pose estimation. Computer Graphics Forum, 41(7):371–381, October 2022. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Qiu:2021:LPD

[QFWH21] Simeng Qiu, Qiang Fu, Congli Wang, and Wolfgang Heidrich. Linear polarization demosaicking for monochrome and colour polarization focal plane arrays. Computer Graphics Forum, 40(6): 77–89, September 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Qu:2022:AML

[QLL^{+}22] Dezhan Qu, Cheng Lv, Yiming Lin, Huijie Zhang, and Rong Wang. AirLens: Multi-level visual exploration of air quality evolution in urban agglomerations. Computer Graphics Forum, 41(3):223–234, June 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Qiu:2021:REE

[QLWQ21] Sheng Qiu, Chen Li, Changbo Wang, and Hong Qin. A rapid, end-to-end, generative model for gaseous phenomena from limited views. Computer Graphics Forum, 40(6):242–257, September 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Qiao:2020:MSI

[QLZ^{+}20] Yu Qiao, Yuhao Liu, Qiang Zhu, Xin Yang, Yuxin Wang, Qiang Zhang, and Xiaopeng Wei. Multi-scale information assembly for image matting. Computer Graphics Forum, 39(7):565–574, October 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Qu:2020:ADG

[QMSL20] Hang Qu, Omid Mashayekhi, Chinmayee Shah, and Philip Levis. Accelerating distributed graphical fluid simulations with micropartitioning. Computer Graphics Forum, 39(1):375–388, February 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Qi:2022:BFW

[QSBJ22] Yang Qi, Dario Seyb, Benedikt Bitterli, and Wojciech Jarosz. A bidirectional formulation for walk on spheres. Computer Graphics Forum, 41(4):51–62, July 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Qin:2020:BRP

[QZL^{+}20] Hongxing Qin, Songshan Zhang, Qihuang Liu, Li Chen, and Baoquan Chen. Book review: PointSkelCNN: Deep Learning-Based 3D Human Skeleton Extraction from Point Clouds. Computer Graphics Forum, 39(7):363–374, October 2020. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Qin:2022:RRP

[QZLC22] Hongxing Qin, Yucheng Zhang, Zhentao Liu, and Baoquan Chen. Rigid registration of point clouds based on partial optimal transport. Computer Graphics Forum, 41(6):365–378, September 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rainer:2022:NPR

[RBRD22] Gilles Rainer, Adrien Bousseau, Tobias Ritschel, and George Drettakis. Neural precomputed radiance transfer. Computer Graphics Forum, 41(2):365–378, May 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rosset:2023:IDC

[RCDB23] Nicolas Rosset, Guillaume Cordonnier, Regis Duvigneau, and Adrien Bousseau. Interactive design of 2D car profiles with aerodynamic feedback. Computer Graphics Forum, 42(2):427–437, May 2023. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rathore:2021:TVE

[RCPW21] Archit Rathore, Nithin Chalapathi, Sourabh Palande, and Bei Wang. TopoAct: Visually exploring the shape of activations in deep learning. Computer Graphics Forum, 40(1):382–397, February 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Reischl:2020:PBR

[RDG20] M. Reischl, E. Derzapf, and M. Guthe. Physically based real-time rendering of teeth and partial restorations. Computer Graphics Forum, 39(1):106–116, February 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ritchie:2023:NMC

[RGJ^{+}23] Daniel Ritchie, Paul Guerrero, R. Kenny Jones, Niloy J. Mitra, Adriana Schulz, Karl D. D. Willis, and Jiajun Wu. Neurosymbolic models for computer graphics. Computer Graphics Forum, 42(2):545–568, May 2023. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rainer:2020:UNE

[RGJW20] Gilles Rainer, Abhijeet Ghosh, Wenzel Jakob, and Tim Weyrich. Unified neural encoding of BTFs. Computer Graphics Forum, 39(2):167–178, May 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rosen:2020:BRI

[RGR20] Matthew Rosen, Keith W. Godin, and Nikunj Raghuvanshi. Book review: Interactive sound propagation for dynamic scenes using 2D wave simulation. Computer Graphics Forum, 39(8):39– 46, December 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Raidou:2020:SDP

[RGW20] Renata G. Raidou, M. Eduard Gröller, and Hsiang-Yun Wu. Slice and dice: A physicalization workflow for anatomical edutainment. Computer Graphics Forum, 39(7):623–634, October 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ryu:2021:FDM

[RKL^{+}21] Hoseok Ryu, Minseok Kim, Seungwhan Lee, Moon Seok Park, Kyoungmin Lee, and Jehee Lee. Functionality-driven musculature retargeting. Computer Graphics Forum, 40(1):341–356, February 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Reinbold:2020:BRV

[RKW20] Christian Reinbold, Alexander Kumpf, and Rüdiger Westermann. Book review: Visualizing the Stability of 2D Point Sets from Dimensionality Reduction Techniques. Computer Graphics Forum, 39(1):333–346, February 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rakotosaona:2020:PLD

[RLG^{+}20] Marie-Julie Rakotosaona, Vittorio La Barbera, Paul Guerrero, Niloy J. Mitra, and Maks Ovsjanikov. PointCleanNet: Learning to denoise and remove outliers from dense point clouds. Computer Graphics Forum, 39(1):185–203, February 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ren:2020:AAE

[RLHS20a] Dongdong Ren, Jinbao Li, Meng Han, and Minglei Shu. Not all areas are equal: A novel separation-restoration-fusion network for image raindrop removal. Computer Graphics Forum, 39(7): 495–505, October 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ren:2020:SNS

[RLHS20b] Dongdong Ren, Jinbao Li, Meng Han, and Minglei Shu. SCGA-Net: Skip connections global attention network for image restoration. Computer Graphics Forum, 39(7):507–518, October 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Riehmann:2020:SCT

[RLR^{+}20] P. Riehmann, G. Molina León, J. Reibert, F. Echtler, and B. Froehlich. Short-contact touch-manipulation of scatterplot matrices on wall displays. Computer Graphics Forum, 39(3):265– 276, June 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rulff:2022:URL

[RMH^{+}22] Joao Rulff, Fabio Miranda, Maryam Hosseini, Marcos Lage, Mark Cartwright, Graham Dove, Juan Bello, and Claudio T. Silva. Urban rhapsody: Large-scale exploration of urban soundscapes. Computer Graphics Forum, 41(3):209–221, June 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Romeo:2020:MFS

[RMSQ20] M. Romeo, C. Monteagudo, and D. Sánchez-Quirós. Muscle and fascia simulation with extended position based dynamics. Computer Graphics Forum, 39(1):134–146, February 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ren:2021:DOS

[RMWO21] Jing Ren, Simone Melzi, Peter Wonka, and Maks Ovsjanikov. Discrete optimization for shape matching. Computer Graphics Forum, 40(5):81–96, August 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Romero:2020:MEN

[ROCP20] Cristian Romero, Miguel A. Otaduy, Dan Casas, and Jesus Perez. Modeling and estimation of nonlinear skin mechanics for animated avatars. Computer Graphics Forum, 39(2):77–88, May 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rocha:2020:PSI

[ROG20] Luís Cláudio Gouveia Rocha, Manuel M. Oliveira, and Eduardo S. L. Gastal. Prefilters for sharp image display. Computer Graphics Forum, 39(2):435–449, May 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ronfard:2021:FDC

[Ron21] Rémi Ronfard. Film directing for computer games and animation. Computer Graphics Forum, 40(2):713–730, May 2021. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rodriguez-Pardo:2023:HWI

[RPPMCG23] Carlos Rodriguez-Pardo, Melania Prieto-Martin, Dan Casas, and Elena Garces. How will it drape like? Capturing fabric mechanics from depth images. Computer Graphics Forum, 42(2):149–160, May 2023. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rittig:2021:NAS

[RSB^{+}21] Tobias Rittig, Denis Sumin, Vahid Babaei, Piotr Didyk, Alexey Voloboy, Alexander Wilkie, Bernd Bickel, Karol Myszkowski, Tim Weyrich, and Jaroslav Křivánek. Neural acceleration of scattering-aware color 3D printing. Computer Graphics Forum, 40(2):205–219, May 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Reda:2021:CNP

[RSGP21] Khairi Reda, Amey A. Salvi, Jack Gray, and Michael E. Papka. Color nameability predicts inference accuracy in spatial visualizations. Computer Graphics Forum, 40(3):49–60, June 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rubio-Sanchez:2021:OAD

[RSLSRÁ21] M. Rubio-Sánchez, D. J. Lehmann, A. Sanchez, and J. L. Rojo-Álvarez. Optimal axes for data value estimation in star coordinates and radial axes plots. Computer Graphics Forum, 40(3): 483–494, June 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Riso:2022:PPP

[RSP22] Marzia Riso, Davide Sforza, and Fabio Pellacini. pOp: Parameter optimization of differentiable vector patterns. Computer Graphics Forum, 41(4):161–168, July 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Rohmer:2021:VSR

[RTK^{+}21] Damien Rohmer, Marco Tarini, Niranjan Kalyanasundaram, Faezeh Moshfeghifar, Marie-Paule Cani, and Victor Zordan. Velocity skinning for real-time stylized skeletal animation. Computer Graphics Forum, 40(2):549–561, May 2021. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Reinbold:2021:PSS

[RW21] Christian Reinbold and Rüdiger Westermann. Parameterized splitting of summed volume tables. Computer Graphics Forum, 40(3):123–134, June 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ren:2022:MMI

[RZZ^{+}22] Haocheng Ren, Hao Zhang, Jia Zheng, Jiaxiang Zheng, Rui Tang, Yuchi Huo, Hujun Bao, and Rui Wang. MINERVAS: Massive INterior EnviRonments VirtuAl Synthesis. Computer Graphics Forum, 41(7):63–74, October 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sharma:2020:GBT

[SAK20] O. Sharma, T. Arora, and A. Khattar. Graph-based transfer function for volume rendering. Computer Graphics Forum, 39 (1):76–88, February 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sorger:2021:ENE

[SAK^{+}21] J. Sorger, A. Arleo, P. Kán, W. Knecht, and M. Waldner. Egocentric network exploration for immersive analytics. Computer Graphics Forum, 40(7):241–252, October 2021. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Schmidt:2022:TAD

[SBB^{+}22] P. Schmidt, J. Born, D. Bommes, M. Campen, and L. Kobbelt. TinyAD: Automatic differentiation in geometry processing made simple. Computer Graphics Forum, 41(5):113–124, August 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Stanko:2020:IGS

[SBBB20] Tibor Stanko, Mikhail Bessmeltsev, David Bommes, and Adrien Bousseau. Integer-grid sketch simplification and vectorization. Computer Graphics Forum, 39(5):149–161, August 2020. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sane:2020:SSP

[SBGC20] Sudhanshu Sane, Roxana Bujack, Christoph Garth, and Hank Childs. A survey of seed placement and streamline selection techniques. Computer Graphics Forum, 39(3):785–809, June 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sugimoto:2022:SOD

[SBH22] R. Sugimoto, C. Batty, and T. Hachisuka. Surface-only dynamic deformables using a boundary element method. Computer Graphics Forum, 41(8):75–86, December 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sorgente:2023:SIM

[SBMS23] T. Sorgente, S. Biasotti, G. Manzini, and M. Spagnuolo. A survey of indicators for mesh quality assessment. Computer Graphics Forum, 42(2):461–483, May 2023. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sharp:2020:LNT

[SC20] Nicholas Sharp and Keenan Crane. A Laplacian for nonmanifold triangle meshes. Computer Graphics Forum, 39(5):69–80, August 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Song:2021:HBP

[SCCZ21] Guoxian Song, Tat-Jen Cham, Jianfei Cai, and Jianmin Zheng. Half-body portrait relighting with overcomplete lighting representation. Computer Graphics Forum, 40(6):371–381, September 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Shen:2020:ZCE

[SCHC20] I-Chao Shen, Ming-Shiuan Chen, Chun-Kai Huang, and Bing-Yu Chen. ZomeFab: Cost-effective hybrid fabrication with zometools. Computer Graphics Forum, 39(1):322–332, February 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Shi:2021:SSL

[SCZ21] Yahao Shi, Xinyu Cao, and Bin Zhou. Self-supervised learning of part mobility from point cloud sequence. Computer Graphics Forum, 40(6):104–116, September 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sharma:2022:PRF

[SDR^{+}22] G. Sharma, B. Dash, A. RoyChowdhury, M. Gadelha, M. Loizou, L. Cao, R. Wang, E. G. Learned-Miller, S. Maji, and E. Kalogerakis. PriFit: Learning to fit primitives improves few shot point cloud segmentation. Computer Graphics Forum, 41(5):39–50, August 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sperrle:2021:SHC

[SEAG^{+}21] F. Sperrle, M. El-Assady, G. Guo, R. Borgo, D. Horng Chau, A. Endert, and D. Keim. A survey of human-centered evaluations in human-centered machine learning. Computer Graphics Forum, 40(3):543–568, June 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Storm-Furru:2020:VTG

[SFB20] S. Storm-Furru and S. Bruckner. VA-TRAC: Geospatial trajectory analysis for monitoring, identification, and verification in fishing vessel operations. Computer Graphics Forum, 39(3):101– 114, June 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Schatz:2021:VAL

[SFMS^{+}21] Karsten Schatz, Juan José Franco-Moreno, Marco Schäfer, Alexander S. Rose, Valerio Ferrario, Jürgen Pleiss, Pere-Pau Vázquez, Thomas Ertl, and Michael Krone. Visual analysis of large-scale protein-ligand interaction data. Computer Graphics Forum, 40(6):394–408, September 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Schneegans:2022:RTR

[SGAG22] S. Schneegans, J. Gilg, V. Ahlers, and A. Gerndt. Real-time rendering of eclipses without incorporation of atmospheric effects. Computer Graphics Forum, 41(7):279–289, October 2022. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sereno:2022:HTT

[SGBI22] Mickael Sereno, Stéphane Gosset, Lonni Besançon, and Tobias Isenberg. Hybrid Touch/Tangible spatial selection in augmented reality. Computer Graphics Forum, 41(3):403–415, June 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sohns:2023:DBV

[SGL23] Jan-Tobias Sohns, Christoph Garth, and Heike Leitte. Decision boundary visualization for counterfactual reasoning. Computer Graphics Forum, 42(1):7–20, February 2023. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Santesteban:2020:SDD

[SGOC20] Igor Santesteban, Elena Garces, Miguel A. Otaduy, and Dan Casas. SoftSMPL: Data-driven modeling of nonlinear soft-tissue dynamics for parametric humans. Computer Graphics Forum, 39(2):65–75, May 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sloup:2021:ORT

[SH21] J. Sloup and V. Havran. Optimizing ray tracing of trimmed NURBS surfaces on the GPU. Computer Graphics Forum, 40 (7):161–172, October 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Schussler:2022:PGV

[SHJD22] Vincent Schüßler, Johannes Hanika, Alisa Jung, and Carsten Dachsbacher. Path guiding with vertex triplet distributions. Computer Graphics Forum, 41(4):1–15, July 2022. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sassen:2020:NDS

[SHR20] Josua Sassen, Klaus Hildebrandt, and Martin Rumpf. Nonlinear deformation synthesis via sparse principal geodesic analysis. Computer Graphics Forum, 39(5):119–132, August 2020. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Siddiqui:2021:PAU

[SHV21] Faizan Siddiqui, Thomas Höllt, and Anna Vilanova. A progressive approach for uncertainty visualization in diffusion tensor imaging. Computer Graphics Forum, 40(3):411–422, June 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Schelling:2021:EVL

[SHVR21] M. Schelling, P. Hermosilla, P.-P. Vázquez, and T. Ropinski. Enabling viewpoint learning through dynamic label generation. Computer Graphics Forum, 40(2):413–423, May 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sagrista:2020:VAF

[SJS20] Antoni Sagristà, Stefan Jordan, and Filip Sadlo. Visual analysis of the finite-time Lyapunov exponent. Computer Graphics Forum, 39(3):331–342, June 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sevastjanova:2022:LVE

[SKB^{+}22] R. Sevastjanova, A. Kalouli, C. Beck, H. Hauptmann, and M. El-Assady. LMFingerprints: Visual explanations of language model embedding spaces through layerwise contextualization scores. Computer Graphics Forum, 41(3):295–307, June 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Schindler:2022:NPA

[SKRW22] M. Schindler, T. Korpitsch, R. G. Raidou, and H.-Y. Wu. Nested papercrafts for anatomical and biological edutainment. Computer Graphics Forum, 41(3):541–553, June 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Schutz:2021:RPC

[SKW21] Markus Schütz, Bernhard Kerbl, and Michael Wimmer. Rendering point clouds with compute shaders and vertex order optimization. Computer Graphics Forum, 40(4):115–126, July 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Subias:2023:WMA

[SL23] J. Daniel Subias and M. Lagunas. In-the-wild material appearance editing using perceptual attributes. Computer Graphics Forum, 42(2):333–345, May 2023. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sietzen:2021:IAC

[SLB^{+}21] Stefan Sietzen, Mathias Lechner, Judy Borowski, Ramin Hasani, and Manuela Waldner. Interactive analysis of CNN robustness. Computer Graphics Forum, 40(7):253–264, October 2021. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sahoo:2022:NFM

[SLB22] S. Sahoo, Y. Lu, and M. Berger. Neural flow map reconstruction. Computer Graphics Forum, 41(3):391–402, June 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sondagger:2022:RTV

[SLCL22] Hyeongseok Son†, JunyongLee†, SunghyunCho, andSeungyongLee.Real− timevideodeblurringvialightweightmotioncompensation.ComputerGraphicsForum, 41 (7) : 177 − −188, October2022.CODENCGF ODY.ISSN0167 −

7055(print), 1467 − 8659(electronic).

Saffo:2020:DCD

[SLJ^{+}20] David Saffo, Aristotelis Leventidis, Twinkle Jain, Michelle A. Borkin, and Cody Dunne. Data comets: Designing a visualization tool for analyzing autonomous aerial vehicle logs with grounded evaluation. Computer Graphics Forum, 39(3): 455–468, June 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Song:2020:BRS

[SLL20] Mofei Song, Yu Liu, and Xiao Fan Liu. Book review: Semi-Supervised 3D Shape Recognition via Multimodal Deep Co-training. Computer Graphics Forum, 39 (7):279–289, October 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sulzer:2021:SSR

[SLMV21] R. Sulzer, L. Landrieu, R. Marlet, and B. Vallet. Scalable surface reconstruction with Delaunay-graph neural networks. Computer Graphics Forum, 40(5):157– 167, August 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Shen:2021:CMV

[SLS^{+}21] I-Chao Shen, Kuan-Hung Liu, Li-Wen Su, Yu-Ting Wu, and Bing-Yu Chen. ClipFlip: Multi-view clipart design. Computer Graphics Forum, 40(1):327– 340, February 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Song:2022:JHO

[SMM22] H.-X. Song, T.-J. Mu, and R. R. Martin. Joint hand and object pose estimation from a single RGB image using high-level 2D constraints. Computer Graphics Forum, 41(7):383–394, October 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Schutz:2020:PRT

[SMOW20] Markus Schütz, Gottfried Mandlburger, Johannes Otepka, and Michael Wimmer. Progressive real-time rendering of one billion points without hierarchical acceleration structures. Computer Graphics Forum, 39(2):51–64, May 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Schertzer:2022:FRT

[SMRB22] Jérémie Schertzer, Corentin Mercier, Sylvain Rousseau, and Tamy Boubekeur. Fiblets for real-time rendering of massive brain tractograms. Computer Graphics Forum, 41(2):447–460, May 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Stadlbauer:2020:IMC

[SMS^{+}20] P. Stadlbauer, D. Mlakar, H.-P. Seidel, M. Steinberger, and R. Zayer. Interactive modeling of cellular structures on surfaces with application to additive manufacturing. Computer Graphics Forum, 39(2):277–289, May 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Seo:2022:SEP

[SOL^{+}22] Kwanggyoon Seo, Seoung Wug Oh, Jingwan Lu, Joon-Young Lee, Seonghyeon Kim, and Junyong Noh. StylePortraitVideo: Editing portrait videos with expression optimization. Computer Graphics Forum, 41(7):165–175, October 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Schutz:2020:FCO

[SOW20] Markus Schütz, Stefan Ohrhallinger, and Michael Wimmer. Fast out-of-core octree generation for massive point clouds. Computer Graphics Forum, 39(7): 155–167, October 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Schoentgen:2020:PBL

[SPDM20] Arnaud Schoentgen, Pierre Poulin, Emmanuelle Darles, and Philippe Meseure. Particle-based liquid control using animation templates. Computer Graphics Forum, 39(8):79–88, December 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Schmidt:2023:SMA

[SPK23] P. Schmidt, D. Pieper, and L. Kobbelt. Surface maps via adaptive triangulations. Computer Graphics Forum, 42(2):103–117, May 2023. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Serpa:2020:BTF

[SR20] Ygor Rebouças Serpa and Maria Andréia Formico Rodrigues. Broadmark: A testing framework for broad-phase collision detection algorithms. Computer Graphics Forum, 39(1):436–449, February 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sun:2020:ZCN

[SRC^{+}20] Zhiyu Sun, Ethan Rooke, Jerome Charton, Yusen He, Jia Lu, and Stephen Baek. ZerNet: Convolutional neural networks on arbitrary surfaces via Zernike local tangent space estimation. Computer Graphics Forum, 39(6):204–216, September 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Stephenson:2021:CPG

[SRK21] Mark Stephenson, Ram Rangan, and Stephen W. Keckler. Cooperative profile guided optimizations. Computer Graphics Forum, 40(8):71–83, December 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Shekhar:2021:IPE

[SRM^{+}21] Sumit Shekhar, Max Reimann, Maximilian Mayer, Amir Semmo, Sebastian Pasewaldt, Jürgen Döllner, and Matthias Trapp. Interactive photo editing on smartphones via intrinsic decomposition. Computer Graphics Forum, 40(2): 497–510, May 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sztrajman:2021:NBR

[SRRW21] Alejandro Sztrajman, Gilles Rainer, Tobias Ritschel, and Tim Weyrich. Neural BRDF representation and importance sampling. Computer Graphics Forum, 40(6):332–346, September 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Silvennoinen:2021:MBD

[SS21] Ari Silvennoinen and Peter-Pike Sloan. Moving basis decomposition for precomputed light transport. Computer Graphics Forum, 40(4):127–137, July 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Singh:2020:FAC

[SSC^{+}20] Gurprit Singh, Kartic Subr, David Coeurjolly, Victor Ostromoukhov, and Wojciech Jarosz. Fourier analysis of correlated Monte Carlo importance sampling. Computer Graphics Forum, 39(1):7–19, February 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sperrle:2021:LCU

[SSKEA21] F. Sperrle, H. Schäfer, D. Keim, and M. El-Assady. Learning contextualized user preferences for co-adaptive guidance in mixed-initiative topic model refinement. Computer Graphics Forum, 40(3):215–226, June 2021. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Solteszova:2020:MLT

[SSS^{+}20a] V. Solteszova, N. N. Smit, S. Stoppel, R. Grüner, and S. Bruckner. Memento: Localized time-warping for spatio-temporal selection. Computer Graphics Forum, 39(1):231–243, February 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sun:2020:ETS

[SSS20b] Yunxin Sun, Tamar Shinar, and Craig Schroeder. Effective time step restrictions for explicit MPM simulation. Computer Graphics Forum, 39(8):55–67, December 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

South:2022:EUL

[SSV^{+}22] Laura South, David Saffo, Olga Vitek, Cody Dunne, and Michelle A. Borkin. Effective use of Likert scales in visualization evaluations: a systematic review. Computer Graphics Forum, 41(3):43–55, June 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Shi:2021:AAA

[SSX^{+}21] D. Shi, F. Sun, X. Xu, X. Lan, D. Gotz, and N. Cao. AutoClips: an automatic approach to video generation from data facts. Computer Graphics Forum, 40 (3):495–505, June 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sondag:2022:VAC

[STX^{+}22] M. Sondag, C. Turkay, K. Xu, L. Matthews, S. Mohr, and D. Archambault. Visual analytics of contact tracing policy simulations during an emergency response. Computer Graphics Forum, 41(3):29–41, June 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Subr:2021:QNN

[Sub21] Kartic Subr. Q-NET: a network for low-dimensional integrals of neural proxies. Computer Graphics Forum, 40(4):61–71, July 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Schreck:2020:PMA

[SW20] Camille Schreck and Chris Wojtan. A practical method for animating anisotropic elastoplastic materials. Computer Graphics Forum, 39(2):89–99, May 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Schreck:2022:CLS

[SW22] Camille Schreck and Chris Wojtan. Coupling 3D liquid simulation with 2D wave propagation for large scale water surface animation using the equivalent sources method. Computer Graphics Forum, 41(2):343–353, May 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Stevens:2020:BRH

[SWB^{+}20] Andrew H. Stevens, Colin Ware, Thomas Butkiewicz, David Rogers, and Greg Abram. Book review: Hairy Slices II: Depth Cues for Visualizing 3D Streamlines Through Cutting Planes. Computer Graphics Forum, 39(3):25–35, June 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Stein:2020:SDV

[SWJG20] Oded Stein, Max Wardetzky, Alec Jacobson, and Eitan Grinspun. A simple discretization of the vector Dirichlet energy. Computer Graphics Forum, 39(5): 81–92, August 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Shen:2022:SMS

[SWW^{+}22a] Yiyang Shen, Yongzhen Wang, Mingqiang Wei, Honghua Chen, Haoran Xie, Gary Cheng, and Fu Lee Wang. Semi-MoreGAN: Semi-supervised generative adversarial network for mixture of rain removal. Computer Graphics Forum, 41(7):443–454, October 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sun:2022:ETP

[SWW^{+}22b] Haoran Sun, Shiyi Wang, Wenhai Wu, Yao Jin, Hujun Bao, and Jin Huang. Efficient texture parameterization driven by perceptual-loss-on-screen. Computer Graphics Forum, 41(7):507–518, October 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Su:2022:UAU

[SXHA22] Haozhe Su, Tao Xue, Chengguizi Han, and Mridul Aanjaneya. A-ULMPM: an adaptively updated Lagrangian material point method for efficient physics simulation without numerical fracture. Computer Graphics Forum, 41(2):325–341, May 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Song:2021:TCR

[SXZS21] Chengfang Song, Chunxia Xiao, Yeting Zhang, and Haigang Sui. Thin cloud removal for single RGB aerial image. Computer Graphics Forum, 40(1):398– 409, February 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

She:2020:LTA

[SYG20] Y. She, Y. Yi, and J. Gu. Learning target-adaptive correlation filters for visual tracking. Computer Graphics Forum, 39(7):387–397, October 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sorokin:2021:LHS

[SYHL21] Maks Sorokin, Wenhao Yu, Sehoon Ha, and C. Karen Liu. Learning human search behavior from egocentric visual inputs. Computer Graphics Forum, 40 (2):389–398, May 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Son:2020:EMD

[SYKE20] S.-H. Son, S.-H. Yoon, M.-S. Kim, and G. Elber. Efficient minimum distance computation for solids of revolution. Computer Graphics Forum, 39(2):535–544, May 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Schottler:2021:VIG

[SYPB21] Sarah Schöttler, Yalong Yang, Hanspeter Pfister, and Benjamin Bach. Visualizing and interacting with geospatial networks: a survey and design space. Computer Graphics Forum, 40(6):5–33, September 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Sun:2021:DFV

[SZM^{+}21] Yifan Sun, Yixuan Zhang, Ali Mosallaei, Michael D. Shah, Cody Dunne, and David Kaeli. Daisen: a framework for visualizing detailed GPU execution. Computer Graphics Forum, 40(3):239–250, June 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Thonat:2021:VBR

[TAA^{+}21] T. Thonat, Y. Aksoy, M. Aittala, S. Paris, F. Durand, and G. Drettakis. Videobased rendering of dynamic stationary environments from unsynchronized inputs. Computer Graphics Forum, 40(4):73–86, July 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Tang:2021:HSD

[TACS21] Jingwei Tang, Vinicius C. Azevedo, Guillaume Cordonnier, and Barbara Solenthaler. Honey, I shrunk the domain: Frequency-aware force field reduction for efficient fluids optimization. Computer Graphics Forum, 40(2):339–353, May 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Trautner:2021:LWI

[TB21] T. Trautner and S. Bruckner. Line Weaver: Importance-driven order enhanced rendering of dense line charts. Computer Graphics Forum, 40(3):399–410, June 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Trettner:2021:GDC

[TBK21] P. Trettner, D. Bommes, and L. Kobbelt. Geodesic distance computation via virtual source propagation. Computer Graphics Forum, 40(5):247–260, August 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Thygesen:2022:LDE

[TBL^{+}22] Signe Sidwall Thygesen, Talha Bin Masood, Mathieu Linares, Vijay Natarajan, and Ingrid Hotz. Level of detail exploration of electronic transition ensembles using hierarchical clustering. Computer Graphics Forum, 41(3):333–344, June 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Trautner:2020:SPM

[TBSB20] T. Trautner, F. Bolte, S. Stoppel, and S. Bruckner. Sunspot plots: Model-based structure enhancement for dense scatter plots. Computer Graphics Forum, 39 (3):551–563, June 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Tennekes:2021:DSO

[TC21] Martijn Tennekes and Min Chen. Design space of origin-destination data visualization. Computer Graphics Forum, 40(3):323–334, June 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ture:2021:NSI

[TÇE^{+}21] Murat Türe, Mustafa Ege Çıklabakkal, Aykut Erdem, Erkut Erdem, Pinar Satılmış, and Ahmet Oguz Akyüz. From noon to sunset: Interactive rendering, relighting, and recolouring of landscape photographs by modifying solar position. Computer Graphics Forum, 40(6):500–515, September 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Troidl:2022:BCS

[TCG^{+}22] Jakob Troidl, Corrado Cali, Eduard Gröller, Hanspeter Pfister, Markus Hadwiger, and Johanna Beyer. Barrio: Customizable spatial neighborhood analysis and comparison for nanoscale brain structures. Computer Graphics Forum, 41 (3):183–194, June 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Tricard:2020:BRR

[TCL20] Thibault Tricard, Frédéric Claux, and Sylvain Lefebvre. Book review: Ribbed Support Vaults for 3D Printing of Hollowed Objects. Computer Graphics Forum, 39(1):147–159, February 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Tang:2022:SOC

[TCT22] Pengbin Tang, Stelian Coros, and Bernhard Thomaszewski. A second order cone programming approach for simulating biphasic materials. Computer Graphics Forum, 41(8):87–93, December 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Tessari:2023:SSB

[TDDB23] L. Tessari, A. Dittebrand, M. J. Doyle, and C. Benthin. Stochastic subsets for BVH construction. Computer Graphics Forum, 42(2):255–267, May 2023. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Tewari:2020:SAN

[TFT^{+}20] A. Tewari, O. Fried, J. Thies, V. Sitzmann, S. Lombardi, K. Sunkavalli, R. Martin-Brualla, T. Simon, J. Saragih, M. Nießner, R. Pandey, S. Fanello, G. Wetzstein, J.-Y. Zhu, C. Theobalt, M. Agrawala, E. Shechtman, D. B Goldman, and M. Zollhöfer. State of the art on neural rendering. Computer Graphics Forum, 39(2):701–727, May 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Trettner:2020:FRQ

[TK20] P. Trettner and L. Kobbelt. Fast and robust QEF minimization using probabilistic quadrics. Computer Graphics Forum, 39(2):325–334, May 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Trettner:2021:SQB

[TK21] P. Trettner and L. Kobbelt. Sampling from quadric-based CSG surfaces. Computer Graphics Forum, 40(8):41–56, December 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Tang:2023:PIN

[TKAS23] Jingwei Tang, Byungsoo Kim, Vinicius C. Azevedo, and Barbara Solenthaler. Physics-informed neural corrector for deformation-based fluid control. Computer Graphics Forum, 42(2):161–173, May 2023. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Tajima:2021:RHW

[TKE21] D. Tajima, Y. Kanamori, and Y. Endo. Relighting humans in the wild: Monocular full-body human relighting with domain adaptation. Computer Graphics Forum, 40(7):205–216, October 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Tretschk:2023:SAD

[TKR^{+}23] Edith Tretschk, Navami Kairanda, Mallikarjun B R, Rishabh Dabral, Adam Kortylewski, Bernhard Egger, Marc Habermann, Pascal Fua, Christian Theobalt, and Vladislav Golyanik. State of the art in dense monocular non-rigid 3D reconstruction. Computer Graphics Forum, 42(2):485–520, May 2023. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Tang:2021:WLY

[TL21] Shusen Tang and Zhouhui Lian. Write like you: Synthesizing your cursive online Chinese handwriting via metric-based meta learning. Computer Graphics Forum, 40(2):141–151, May 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Tian:2020:NUS

[TLC20] Yufeng Tian, Xin Li, and Falai Chen. Non-uniform subdivision surfaces with sharp features. Computer Graphics Forum, 39(6):232–242, September 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Thompson:2020:UDS

[TLLS20] J. Thompson, Z. Liu, W. Li, and J. Stasko. Understanding the design space and authoring paradigms for animated data graphics. Computer Graphics Forum, 39 (3):207–218, June 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Tapia:2021:PSR

[TRPO21] Javier Tapia, Cristian Romero, Jesús Pérez, and Miguel A. Otaduy. Parametric skeletons with reduced soft-tissue deformations. Computer Graphics Forum, 40 (6):34–46, September 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Tang:2020:ESS

[TSK^{+}20] Xuan Tang, Teseo Schneider, Shoaib Kamil, Aurojit Panda, Jinyang Li, and Daniele Panozzo. EGGS: Sparsity-specific code generation. Computer Graphics Forum, 39(5):209–219, August 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Takimoto:2022:DHA

[TST^{+}22] Yusuke Takimoto, Hiroyuki Sato, Hikari Takehara, Keishiro Uragaki, Takehiro Tawara, Xiao Liang, Kentaro Oku, Wataru Kishimoto, and Bo Zheng. Dressi: a hardware-agnostic differentiable renderer with reactive shader packing and soft rasterization. Computer Graphics Forum, 41(2):13–27, May 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Tewari:2022:ANR

[TTM^{+}22] A. Tewari, J. Thies, B. Mildenhall, P. Srinivasan, E. Tretschk, W. Yifan, C. Lassner, V. Sitzmann, R. Martin-Brualla, S. Lombardi, T. Simon, C. Theobalt, M. Nießner, J. T. Barron, G. Wetzstein, M. Zollhöfer, and V. Golyanik. Advances in neural rendering. Computer Graphics Forum, 41(2):703–735, May 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Tojo:2022:RPV

[TU22] Kenji Tojo and Nobuyuki Umetani. Recolorable posterization of volumetric radiance fields using visibility-weighted palette extraction. Computer Graphics Forum, 41(4):149–160, July 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Todova:2022:WGM

[TWF22] L. Tódová, A. Wilkie, and L. Fascione. Wide gamut moment-based constrained spectral uplifting. Computer Graphics Forum, 41(6):258–272, September 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Tyagi:2022:IWF

[TZP^{+}22] Anjul Tyagi, Jian Zhao, Pushkar Patel, Swasti Khurana, and Klaus Mueller. Infographics wizard: Flexible infographics authoring and design exploration. Computer Graphics Forum, 41(3):121–132, June 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Tan:2022:RTM

[TZX^{+}22] Haowen Tan, Junqiu Zhu, Yanning Xu, Xiangxu Meng, Lu Wang, and Ling-Qi Yan. Real-time microstructure rendering with MIP-Mapped normal map samples. Computer Graphics Forum, 41(1):495–506, February 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Unterguggenberger:2021:CMB

[UKPW21] J. Unterguggenberger, B. Kerbl, J. Pernsteiner, and M. Wimmer. Conservative meshlet bounds for robust culling of skinned meshes. Computer Graphics Forum, 40(7):57–69, October 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Usta:2022:TSM

[UPE22] B. Usta, S. Pont, and E. Eisemann. Targeting shape and material in lighting design. Computer Graphics Forum, 41(7):299–309, October 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ulmer:2021:PPV

[USK21] A. Ulmer, D. Sessler, and J. Kohlhammer. ProBGP: Progressive visual analytics of live BGP updates. Computer Graphics Forum, 40(3):37–48, June 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

vonBuelow:2022:FGM

[vBGF22] Max von Buelow, Stefan Guthe, and Dieter W. Fellner. Fine-grained memory profiling of GPGPU kernels. Computer Graphics Forum, 41(7):227–235, October 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Vevoda:2022:WSR

[VBRKW22] P. Vévoda, T. Bashford-Rogers, M. Kolářová, and A. Wilkie. A wide spectral range sky radiance model. Computer Graphics Forum, 41(7):291–298, October 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Vernier:2021:GSD

[VCT21] E. F. Vernier, J. L. D. Comba, and A. C. Telea. Guided stable dynamic projections. Computer Graphics Forum, 40(3):87–98, June 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

vandeRuit:2021:MPM

[vdRE21] M. van de Ruit and E. Eisemann. A multi-pass method for accelerated spectral sampling. Computer Graphics Forum, 40(7):141–148, October 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Vitsas:2023:PTB

[VEPG23] N. Vitsas, I. Evangelou, G. Papaioannou, and A. Gkaravelis. Parallel transformation of bounding volume hierarchies into oriented bounding box trees. Computer Graphics Forum, 42(2):245–254, May 2023. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Vernier:2020:QET

[VGdS^{+}20] E. F. Vernier, R. Garcia, I. P. da Silva, J. L. D. Comba, and A. C. Telea. Quantitative evaluation of time-dependent multidimensional projection techniques. Computer Graphics Forum, 39(3):241–252, June 2020. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Vimont:2020:IMS

[VGL^{+}20] Ulysse Vimont, James Gain, Maud Lastic, Guillaume Cordonnier, Babatunde Abiodun, and Marie-Paule Cani. Interactive meso-scale simulation of skyscapes. Computer Graphics Forum, 39(2):585–596, May 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Vuillamy:2022:SPP

[VLLA22] J. Vuillamy, A. Lieutier, F. Lafarge, and P. Alliez. Simplification of 2D polygonal partitions via point-line projective duality, and application to urban reconstruction. Computer Graphics Forum, 41(6):379–393, September 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

vanOnzenoodt:2021:BNP

[vOSRR21] Christian van Onzenoodt, Gurprit Singh, Timo Ropinski, and Tobias Ritschel. Blue noise plots. Computer Graphics Forum, 40(2):425–433, May 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Vitsas:2020:IGF

[VPGV20] N. Vitsas, G. Papaioannou, A. Gkaravelis, and A. A. Vasilakis. Illuminationguided furniture layout optimization. Computer Graphics Forum, 39(2):291– 301, May 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Vuillemot:2021:BOD

[VRBT21] R. Vuillemot, P. Rivière, A. Beignon, and A. Tabard. Boundary objects in design studies: Reflections on the collaborative creation of isochrone maps. Computer Graphics Forum, 40(3):349–360, June 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Vernier:2020:QCT

[VSC^{+}20] E. Vernier, M. Sondag, J. Comba, B. Speckmann, A. Telea, and K. Verbeek. Quantitative comparison of time-dependent treemaps. Computer Graphics Forum, 39(3):393–404, June 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Vidaurre:2020:FCG

[VSGC20] Raquel Vidaurre, Igor Santesteban, Elena Garces, and Dan Casas. Fully convolutional graph neural networks for parametric virtual try-on. Computer Graphics Forum, 39(8):145–156, December 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

vanToll:2021:AMC

[vTP21] W. van Toll and J. Pettré. Algorithms for microscopic crowd simulation: Advancements in the 2010s. Computer Graphics Forum, 40(2):731–754, May 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Vasilakis:2020:SMR

[VVP20] A. A. Vasilakis, K. Vardis, and G. Papaioannou. A survey of multifragment rendering. Computer Graphics Forum, 39(2):623–642, May 2020. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Vechev:2022:CDK

[VZTH22] V. Vechev, J. Zarate, B. Thomaszewski, and O. Hilliges. Computational design of kinesthetic garments. Computer Graphics Forum, 41(2):535–546, May 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2020:RID

[Wan20] Cuilan Wang. Representative isovalue detection and isosurface segmentation using novel isosurface measures. Computer Graphics Forum, 39(3):37–47, June 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wen:2020:BRA

[WBB^{+}20] Q. Wen, D. Bradley, T. Beeler, S. Park, O. Hilliges, J. Yong, and F. Xu. Book review: Accurate Real-time 3D Gaze Tracking Using a Lightweight Eyeball Calibration. Computer Graphics Forum, 39(2):475–485, May 2020. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2020:VDU

[WBL20] Yunzhe Wang, George Baciu, and Chenhui Li. Visualizing dynamics of urban regions through a geo-semantic graph-based method. Computer Graphics Forum, 39(1):405–419, February 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Whalen:2021:SSJ

[WBM21] E. Whalen, A. Beyene, and C. Mueller. SimJEB: Simulated jet engine bracket dataset. Computer Graphics Forum, 40(5):9–17, August 2021. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2020:FES

[WBZB20] M. Wang, D. Bradley, S. Zafeiriou, and T. Beeler. Facial expression synthesis using a global-local multilinear framework. Computer Graphics Forum, 39(2): 235–245, May 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wu:2022:SAS

[WCM22] Mengxi Wu, Yi-Jen Chiang, and Christopher Musco. Streaming approach to in situ selection of key time steps for time-varying volume data. Computer Graphics Forum, 41(3):309–320, June 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wei:2023:RPD

[WCR^{+}23] Jiayi Wei, Jiong Chen, Damien Rohmer, Pooran Memari, and Mathieu Desbrun. Robust pointset denoising of piecewise-smooth surfaces through line processes. Computer Graphics Forum, 42(2):175–189, May 2023. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2022:LSS

[WCS^{+}22] C. Wang, B. Chen, Hp. Seidel, K. Myszkowski, and A. Serrano. Learning a selfsupervised tone mapping operator via feature contrast masking loss. Computer Graphics Forum, 41(2):71–84, May 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wolf:2021:PBB

[WCSH21] Thomas Wolf, Victor Cornillère, and Olga Sorkine-Hornung. Physically-based book simulation with freeform developable surfaces. Computer Graphics Forum, 40(2):449–460, May 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wen:2020:CRR

[WCU^{+}20] Jiahao Wen, Jiong Chen, Nobuyuki Umetani, Hujun Bao, and Jin Huang. Cosserat rod with rh-adaptive discretization. Computer Graphics Forum, 39 (7):143–154, October 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2020:RTG

[WDH20] Beibei Wang, Hong Deng, and Nicolas Holzschuch. Real-time glints rendering with pre-filtered discrete stochastic microfacets. Computer Graphics Forum, 39(6):144–154, September 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Weier:2021:OPS

[WDH^{+}21] Philippe Weier, Marc Droske, Johannes Hanika, Andrea Weidlich, and Jiří Vorba. Optimised path space regularisation. Computer Graphics Forum, 40 (4):139–151, July 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2020:LEC

[WDK^{+}20] Bin Wang, Yuanmin Deng, Paul Kry, Uri Ascher, Hui Huang, and Baoquan Chen. Learning elastic constitutive material and damping models. Computer Graphics Forum, 39(7):81–91, October 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2022:GDR

[WDY^{+}22] Ziyu Wang, Yu Deng, Jiaolong Yang, Jingyi Yu, and Xin Tong. Generative deformable radiance fields for disentangled image synthesis of topology-varying objects. Computer Graphics Forum, 41(7):431–442, October 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2022:FER

[WFJ^{+}22] Bolun Wang, Zachary Ferguson, Xin Jiang, Marco Attene, Daniele Panozzo, and Teseo Schneider. Fast and exact root parity for continuous collision detection. Computer Graphics Forum, 41(2):355–363, May 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2021:APS

[WGH^{+}21] Yun Wang, Yi Gao, Ray Huang, Weiwei Cui, Haidong Zhang, and Dongmei Zhang. Animated presentation of static infographics with InfoMotion. Computer Graphics Forum, 40(3):507–518, June 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Walczak:2022:UPB

[WGT^{+}22] Lars Walczak, Joachim Georgii, Lennart Tautz, Mathias Neugebauer, Isaac Wamala, Simon Sündermann, Volkmar Falk, and Anja Hennemuth. Using positionbased dynamics for simulating mitral valve closure and repair procedures. Computer Graphics Forum, 41(1):270–287, February 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2020:AMC

[WH20] S. Wang and N. Holzschuch. Adaptive matrix completion for fast visibility computations with many lights rendering. Computer Graphics Forum, 39(4):47–58, July 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wu:2022:OGB

[WHPG22] Kui Wu, Xu He, Zherong Pan, and Xifeng Gao. Occluder generation for buildings in digital games. Computer Graphics Forum, 41(7):205–214, October 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Weiss:2022:FNR

[WHW22] S. Weiss, P. Hermüller, and R. Westermann. Fast neural representations for direct volume rendering. Computer Graphics Forum, 41(6):196–211, September 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wolff:2023:DPG

[WHZ^{+}23] Katja Wolff, Philipp Herholz, Verena Ziegler, Frauke Link, Nico Brügel, and Olga Sorkine-Hornung. Designing personalized garments with body movement. Computer Graphics Forum, 42(1):180–194, February 2023. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2022:STS

[WJK22] R. Wang, C. Jung, and Y. Kim. Seeing through sounds: Mapping auditory dimensions to data and charts for people with visual impairments. Computer Graphics Forum, 41(3):71–83, June 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wiewel:2020:LSS

[WKA^{+}20] S. Wiewel, B. Kim, V. C. Azevedo, B. Solenthaler, and N. Thuerey. Latent space subdivision: Stable and controllable time predictions for fluid flow. Computer Graphics Forum, 39(8):15–25, December 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Kim:2020:SAS

[wKH20] Seung wook Kim and JungHyun Han. Simulation of arbitrarily-shaped magnetic objects. Computer Graphics Forum, 39(7):119–130, October 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2022:ESS

[WL22] Q. Wang and R. S. Laramee. EHR STAR: The state-of-the-art in interactive EHR visualization. Computer Graphics Forum, 41(1):69–105, February 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wetzels:2022:BDI

[WLG22] Florian Wetzels, Heike Leitte, and Christoph Garth. Branch decompositionindependent edit distances for merge trees. Computer Graphics Forum, 41(3): 367–378, June 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wong:2021:RRD

[WLNM21] Yu-Shiang Wong, Changjian Li, Matthias Nießner, and Niloy J. Mitra. RigidFusion: RGB-D scene reconstruction with rigidly-moving objects. Computer Graphics Forum, 40(2):511–522, May 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2022:UUS

[WLT^{+}22] Haoran Wang, Jiaxin Li, Alexandru Telea, Jiří Kosinka, and Zizhao Wu. USTNet: Unsupervised shape-to-shape translation via disentangled representations. Computer Graphics Forum, 41(7):141–152, October 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wei:2022:SMA

[WM22] Hui Wei and Lingjiang Meng. A stereo matching algorithm for high-precision guidance in a weakly textured industrial robot environment dominated by planar facets. Computer Graphics Forum, 41(1):288–300, February 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Weinrauch:2023:VLS

[WMS^{+}23] A. Weinrauch, D. Mlakar, H. P. Seidel, M. Steinberger, and R. Zayer. A variational loop shrinking analogy for handle and tunnel detection and Reeb graph construction on surfaces. Computer Graphics Forum, 42(2):309–320, May 2023. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2020:CRT

[WMU^{+}20] Feng Wang, Nathan Marshak, Will Usher, Carsten Burstedde, Aaron Knoll, Timo Heister, and Chris R. Johnson. CPU ray tracing of tree-based adaptive mesh refinement data. Computer Graphics Forum, 39(3):1–12, June 2020. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wu:2020:STM

[WNT^{+}20] Hsiang-Yun Wu, Benjamin Niedermann, Shigeo Takahashi, Maxwell J. Roberts, and Martin Nöllenburg. A survey on transit map layout — from design, machine, and human perspectives. Computer Graphics Forum, 39(3):619–646, June 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2020:MHP

[WQL^{+}20] Min Wang, Feng Qiu, Wentao Liu, Chen Qian, Xiaowei Zhou, and Lizhuang Ma. Monocular human pose and shape reconstruction using part differentiable rendering. Computer Graphics Forum, 39(7):351–362, October 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2020:PHM

[WRZ20] Yangang Wang, Ruting Rao, and Changqing Zou. Personalized hand modeling from multiple postures with multi-view color images. Computer Graphics Forum, 39(7):339–350, October 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wu:2020:BRD

[WS20] Yunjie Wu and Zhengxing Sun. Book review: DFR: Differentiable Function Rendering for Learning 3D Generation from Images. Computer Graphics Forum, 39(5):241–252, August 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wei:2020:NBB

[WSN^{+}20] M. Wei, Z. Song, Y. Nie, J. Wu, Z. Ji, Y. Guo, H. Xie, J. Wang, and F. L. Wang. Normal-based bas-relief modelling via near-lighting photometric stereo. Computer Graphics Forum, 39(1):204–219, February 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2021:SAC

[WSP21] Ziqi Wang, Peng Song, and Mark Pauly. State of the art on computational design of assemblies with rigid parts. Computer Graphics Forum, 40(2):633–657, May 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2021:OPL

[WTB^{+}21] Qisi Wang, Yutian Tao, Eric Brandt, Court Cutting, and Eftychios Sifakis. Optimized processing of localized collisions in projective dynamics. Computer Graphics Forum, 40(6):382–393, September 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2020:BRL

[WTO^{+}20] Tong Wang, Wenyuan Tao, Chung-Ming Own, Xiantuo Lou, and Yuehua Zhao. Book review: The Layerizing VoxPoint Annular Convolutional Network for 3D Shape Classification. Computer Graphics Forum, 39(7):291–300, October 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2022:CSG

[WWF^{+}22] Jie Wang, Yongzhen Wang, Yidan Feng, Lina Gong, Xuefeng Yan, Haoran Xie, Fu Lee Wang, and Mingqiang Wei. Contrastive semantic-guided image smoothing network. Computer Graphics Forum, 41(7):335–346, October 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wolligandt:2021:MDG

[WWRT21] S. Wolligandt, T. Wilde, C. Rössl, and H. Theisel. A modified double gyre with ground truth hyperbolic trajectories for flow visualization. Computer Graphics Forum, 40(1):209–221, February 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2020:BRG

[WWW20] Junpeng Wang, Jun Wu, and Rüdiger Westermann. Book review: A Globally Conforming Lattice Structure for 2D Stress Tensor Visualization. Computer Graphics Forum, 39(3):417–427, June 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wen:2022:SRS

[WWZ^{+}22] Tao Wen, Beibei Wang, Lei Zhang, Jie Guo, and Nicolas Holzschuch. SVBRDF recovery from a single image with highlights using a pre-trained generative adversarial network. Computer Graphics Forum, 41(6):110–123, September 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2021:RGN

[WXL^{+}21] Kai Wang, Xianghao Xu, Leon Lei, Selena Ling, Natalie Lindsay, Angel X. Chang, Manolis Savva, and Daniel Ritchie. Roominoes: Generating novel 3D floor plans from existing 3D rooms. Computer Graphics Forum, 40(5):57–69, August 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Weiss:2020:BRI

[WYA^{+}20] Tomer Weiss, Ilkay Yildiz, Nitin Agarwal, Esra Ataer-Cansizoglu, and Jae-Woo Choi. Book review: Image-Driven Furniture Style for Interactive 3D Scene Modeling. Computer Graphics Forum, 39(7):57–68, October 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2020:BRR

[WYM^{+}20] Peihui Wang, Na Yuan, Yuewen Ma, Shiqing Xin, Ying He, Shuangmin Chen, Jian Xu, and Wenping Wang. Book review: Robust Computation of 3D Apollonius Diagrams. Computer Graphics Forum, 39(7):43–55, October 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2022:TBI

[WYZ^{+}22] Yongzhen Wang, Xuefeng Yan, Kaiwen Zhang, Lina Gong, Haoran Xie, Fu Lee Wang, and Mingqiang Wei. TogetherNet: Bridging image restoration and object detection together via dynamic enhancement learning. Computer Graphics Forum, 41(7):465–476, October 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2020:ACS

[WZB20] B. Wang, M. Zheng, and J. Barbič. Adjustable constrained soft-tissue dynamics. Computer Graphics Forum, 39(7):69–79, October 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Wang:2022:EEM

[WZX^{+}22] Luyuan Wang, Hanyuan Zhang, Qinjie Xiao, Hao Xu, Chunhua Shen, and Xiaogang Jin. Effective eyebrow matting with domain adaptation. Computer Graphics Forum, 41(7):347–358, October 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Xiao:2021:DMR

[XABR21] X. Xiao, P. Alliez, L. Busé, and L. Rineau. Delaunay meshing and repairing of NURBS models. Computer Graphics Forum, 40(5):125–142, August 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Xiao:2022:IRC

[XCC22] Yanyang Xiao, Juan Cao, and Zhonggui Chen. Image representation on curved optimal triangulation. Computer Graphics Forum, 41(6):23–36, September 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Xu:2020:CDO

[XFS^{+}20] Hao Xu, Tianwen Fu, Peng Song, Mingjun Zhou, Chi-Wing Fu, and Niloy J. Mitra. Computational design and optimization of non-circular gears. Computer Graphics Forum, 39(2):399–409, May 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Xiao:2020:EBC

[XHZ^{+}20] Chufeng Xiao, Chu Han, Zhuming Zhang, Jing Qin, Tien-Tsin Wong, Guoqiang Han, and Shengfeng He. Example-based colourization via dense encoding pyramids. Computer Graphics Forum, 39(1):20–33, February 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Xie:2021:IDF

[XK21] Kaixiang Xie and Paul G. Kry. Inverse dynamics filtering for sampling-based motion control. Computer Graphics Forum, 40(6):304–314, September 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Xu:2020:SMA

[XL20] Sen-Zhe Xu and Yu-Kun Lai. Simultaneous multi-attribute image-to-image translation using parallel latent transform networks. Computer Graphics Forum, 39(7):531–542, October 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Xie:2020:ACD

[XLKvdP20] Zhaoming Xie, Hung Yu Ling, Nam Hee Kim, and Michiel van de Panne. ALL-STEPS: Curriculum-driven learning of stepping stone skills. Computer Graphics Forum, 39(8):213–224, December 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Xu:2020:SAU

[XOW^{+}20] Kai Xu, Alvitta Ottley, Conny Walchshofer, Marc Streit, Remco Chang, and John Wenskovitch. Survey on the analysis of user interactions and visualization provenance. Computer Graphics Forum, 39(3):757–783, June 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Xu:2020:UIR

[XSW^{+}20] Zilin Xu, Qiang Sun, Lu Wang, Yanning Xu, and Beibei Wang. Unsupervised image reconstruction for gradient-domain volumetric rendering. Computer Graphics Forum, 39(7):193–203, October 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Xu:2022:EME

[XSZX22] Shaobin Xu, Minghui Sun, Zhengtai Zhang, and Hao Xue. Exploring multivariate event sequences with an interactive similarity builder. Computer Graphics Forum, 41(3):271–282, June 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Xie:2021:EMD

[XTL^{+}21] Peng Xie, Wenyuan Tao, Jie Li, Wentao Huang, and Siming Chen. Exploring multi-dimensional data via subset embedding. Computer Graphics Forum, 40 (3):75–86, June 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Xie:2022:NFV

[XTS^{+}22] Yiheng Xie, Towaki Takikawa, Shunsuke Saito, Or Litany, Shiqin Yan, Numair Khan, Federico Tombari, James Tompkin, Vincent Sitzmann, and Srinath Sridhar. Neural fields in visual computing and beyond. Computer Graphics Forum, 41(2):641–676, May 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Xia:2020:GPS

[XWHM20] Mengqi (Mandy) Xia, Bruce Walter, Christophe Hery, and Steve Marschner. Gaussian product sampling for rendering layered materials. Computer Graphics Forum, 39(1):420–435, February 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Xia:2020:SVA

[XXeL^{+}20] Meng Xia, Min Xu, Chuan en Lin, Ta Ying Cheng, Huamin Qu, and Xiaojuan Ma. SeqDynamics: Visual analytics for evaluating online problem-solving dynamics. Computer Graphics Forum, 39(3):511–522, June 2020. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yang:2022:LSP

[YBF22] Y. Yang, C. Barnes, and A. Finkelstein. Learning from shader program traces. Computer Graphics Forum, 41(2):41–56, May 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yang:2021:SIT

[YBRP21] Haiyan Yang, Rafael Ballester-Ripoll, and Renato Pajarola. SenVis: Interactive tensor-based sensitivity visualization. Computer Graphics Forum, 40(3): 275–286, June 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ye:2021:DRS

[YDPG21] Wenjie Ye, Yue Dong, Pieter Peers, and Baining Guo. Deep reflectance scanning: Recovering spatially-varying material appearance from a flash-lit video sequence. Computer Graphics Forum, 40(6):409–427, September 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yue:2021:IIQ

[YGW^{+}21] Xuanwu Yue, Qiao Gu, Deyun Wang, Huamin Qu, and Yong Wang. iQUANT: Interactive quantitative investment using sparse regression factors. Computer Graphics Forum, 40(3):189–200, June 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yuan:2021:CSV

[YGY^{+}21] Mengke Yuan, Bernard Ghanem, Dong-Ming Yan, Baoyuan Wu, Xiaopeng Zhang, and Peter Wonka. Customized summarizations of visual data collections. Computer Graphics Forum, 40(6):347–370, September 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yi:2022:MSA

[YJS^{+}22] Shinyoung Yi, Daniel S. Jeon, Ana Serrano, Se-Yoon Jeong, Hui-Yong Kim, Diego Gutierrez, and Min H. Kim. Modelling surround-aware contrast sensitivity for HDR displays. Computer Graphics Forum, 41(1):350–363, February 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yang:2021:LRT

[YKL21] Dongseok Yang, Doyeon Kim, and Sung-Hee Lee. LoBSTr: Real-time lower-body pose prediction from sparse upper-body tracking signals. Computer Graphics Forum, 40(2):265–275, May 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yang:2022:PHO

[YLC^{+}22] Jinlin Yang, Shibo Liu, Shuangming Chai, Ligang Liu, and Xiao-Ming Fu. Precise high-order meshing of 2D domains with rational Bézier curves. Computer Graphics Forum, 41(5):79–88, August 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ye:2022:NLG

[YLHX22] Yongjing Ye, Libin Liu, Lei Hu, and Shihong Xia. Neural3Points: Learning to generate physically realistic full-body motion for virtual reality users. Computer Graphics Forum, 41(8):183–194, December 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yu:2021:LAN

[YLL^{+}21] Hanning Yu, Wentao Liu, Chengjiang Long, Bo Dong, Qin Zou, and Chunxia Xiao. Luminance attentive networks for HDR image and panorama reconstruction. Computer Graphics Forum, 40(7):181–192, October 2021. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yang:2020:STA

[YLS20] Lei Yang, Shiqiu Liu, and Marco Salvi. A survey of temporal antialiasing techniques. Computer Graphics Forum, 39(2):607–621, May 2020. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yan:2021:SFC

[YMS^{+}21] Lin Yan, Talha Bin Masood, Raghavendra Sridharamurthy, Farhan Rasheed, Vijay Natarajan, Ingrid Hotz, and Bei Wang. Scalar field comparison with topological descriptors: Properties and applications for scientific visualization. Computer Graphics Forum, 40(3):599–633, June 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yan:2022:APS

[YPZ^{+}22] Ming Yan, Yuanyuan Pu, Pengzheng Zhao, Dan Xu, Hao Wu, Qiuxia Yang, and Ruxin Wang. Abstract painting synthesis via decremental optimization. Computer Graphics Forum, 41(7):419–430, October 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ye:2020:MEB

[YSLF20] Chunyang Ye, Jian-Ping Su, Ligang Liu, and Xiao-Ming Fu. Memory-efficient bijective parameterizations of very-large-scale models. Computer Graphics Forum, 39(7):1–12, October 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yang:2022:BML

[YT22] Xingchao Yang and Takafumi Taketomi. BareSkinNet: De-makeup and delighting via 3D face reconstruction. Computer Graphics Forum, 41(7):623–634, October 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yang:2023:MER

[YTK23] Xingchao Yang, Takafumi Taketomi, and Yoshihiro Kanamori. Makeup extraction of 3D representation via illumination-aware image decomposition. Computer Graphics Forum, 42(2):293–307, May 2023. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Ye:2021:CDB

[YUIH21] Z. Ye, N. Umetani, T. Igarashi, and T. Hoffmann. A curvature and densitybased generative representation of shapes. Computer Graphics Forum, 40(1): 38–53, February 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yang:2023:SDA

[YVK^{+}23] J. Yang, N. Vining, S. Kheradmand, N. Carr, L. Sigal, and A. Sheffer. Subpixel deblurring of anti-aliased raster clip-art. Computer Graphics Forum, 42 (2):61–76, May 2023. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yin:2020:LTS

[YY20] Zhiqi Yin and KangKang Yin. Linear time stable PD controllers for physicsbased character animation. Computer Graphics Forum, 39(8):191–200, December 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yang:2020:SBM

[YYH20] Yanzhe Yang, Jimei Yang, and Jessica Hodgins. Statistics-based motion synthesis for social conversations. Computer Graphics Forum, 39(8):201–212, December 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Yang:2022:LOP

[YZ22] Y. X. Yang and S. G. Zhang. Local offset point cloud transformer based implicit surface reconstruction. Computer Graphics Forum, 41(7):99–107, October 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhang:2021:DHE

[ZA21] Y. Zhang and T. O. Aydın. Deep HDR estimation with generative detail reconstruction. Computer Graphics Forum, 40(2):179–190, May 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhang:2022:PRG

[ZA22] Xiaowei Zhang and Daniel Aliaga. Procedural roof generation from a single satellite image. Computer Graphics Forum, 41(2):249–260, May 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhao:2021:PDD

[ZAB^{+}21] Tong Zhao, Pierre Alliez, Tamy Boubekeur, Laurent Busé, and Jean-Marc Thiery. Progressive discrete domains for implicit surface reconstruction. Computer Graphics Forum, 40(5):143–156, August 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zeng:2022:RTF

[ZCC22] Z. Zeng, S. Cotin, and H. Courtecuisse. Real-time FE simulation for large-scale problems using precondition-based contact resolution and isolated DOFs constraints. Computer Graphics Forum, 41(6):418–434, September 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhang:2023:LTH

[ZCH^{+}23] Yunbo Zhang, Alexander Clegg, Sehoon Ha, Greg Turk, and Yuting Ye. Learning to transfer in-hand manipulations using a greedy shape curriculum. Computer Graphics Forum, 42(2):25–36, May 2023. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zeng:2022:TSN

[ZCLJ22] D. Zeng, E. Chambers, D. Letscher, and T. Ju. Topological simplification of nested shapes. Computer Graphics Forum, 41(5):161–173, August 2022. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhang:2020:MAI

[ZDS20] P. Zhang, D. DeFord, and J. Solomon. Medial axis isoperimetric profiles. Computer Graphics Forum, 39(5):1–13, August 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zsolnai-Feher:2020:PME

[ZFWW20] Károly Zsolnai-Fehér, Peter Wonka, and Michael Wimmer. Photorealistic material editing through direct image manipulation. Computer Graphics Forum, 39 (4):107–120, July 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhang:2020:CFW

[ZGP^{+}20] Chenhao Zhang, Shanshan Gao, Xiao Pan, Yuting Wang, and Yuanfeng Zhou. Coarse to fine: Weak feature boosting network for salient object detection. Computer Graphics Forum, 39(7):411–420, October 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhou:2022:FAH

[ZHHK22] J. Zhou, G. J. Hettinga, S. Houwink, and J. Kosinka. Feature-adaptive and hierarchical subdivision gradient meshes. Computer Graphics Forum, 41(1): 389–401, February 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhao:2020:NPP

[ZHL^{+}20] Zipeng Zhao, Kemeng Huang, Chen Li, Changbo Wang, and Hong Qin. A novel plastic phase-field method for ductile fracture with GPU optimization. Computer Graphics Forum, 39(7):105–117, October 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhang:2021:FUL

[ZJA21] Jiayi Eris Zhang, Alec Jacobson, and Marc Alexa. Fast updates for least-squares rotational alignment. Computer Graphics Forum, 40(2):13–22, May 2021. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhao:2022:FGS

[ZJZ^{+}22] Y. Q. Zhao, Z. Jin, H. Y. Zhao, F. Zhang, Z. W. Tao, C. F. Dou, X. H. Xu, and D. H. Liu. Fine-grained scene graph generation with overlap region and geometrical center. Computer Graphics Forum, 41(7):359–370, October 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhou:2021:ASI

[ZK21a] Xilong Zhou and Nima Khademi Kalantari. Adversarial single-image SVBRDF estimation with hybrid training. Computer Graphics Forum, 40(2):315–325, May 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhu:2021:SSI

[ZK21b] Jialin Zhu and Tom Kelly. Seamless satellite-image synthesis. Computer Graphics Forum, 40(7):193–204, October 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zeng:2023:TTO

[ZK23] Libing Zeng and Nima Khademi Kalantari. Test-time optimization for video depth estimation using pseudo reference depth. Computer Graphics Forum, 42(1):195–205, February 2023. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhang:2020:HPT

[ZLL20] Jinsong Zhang, Xingzi Liu, and Kun Li. Human pose transfer by adaptive hierarchical deformation. Computer Graphics Forum, 39(7):325–337, October 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhang:2022:VLD

[ZLL^{+}22a] Zibo Zhang, Xueting Liu, Chengze Li, Huisi Wu, and Zhenkun Wen. Vectorizing line drawings of arbitrary thickness via boundary-based topology reconstruction. Computer Graphics Forum, 41(2):433–445, May 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zong:2022:RSR

[ZLL^{+}22b] Jonathan Zong, Crystal Lee, Alan Lundgard, JiWoong Jang, Daniel Hajas, and Arvind Satyanarayan. Rich screen reader experiences for accessible data visualization. Computer Graphics Forum, 41(3):15–27, June 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhao:2022:ESS

[ZLW^{+}22] Chongyao Zhao, Jinkeng Lin, Tianyu Wang, Hujun Bao, and Jin Huang. Efficient and stable simulation of inextensible Cosserat rods by a compact representation. Computer Graphics Forum, 41(7):567–578, October 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zheng:2022:SSI

[ZLWT22] X. Zheng, Y. Liu, P. Wang, and X. Tong. SDF-StyleGAN: Implicit SDF-based StyleGAN for 3D shape generation. Computer Graphics Forum, 41(5):52–63, August 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhang:2020:CGC

[ZLY^{+}20] Ling Zhang, Chengjiang Long, Qingan Yan, Xiaolong Zhang, and Chunxia Xiao. CLA-GAN: A context and lightness aware generative adversarial network for shadow removal. Computer Graphics Forum, 39(7):483–494, October 2020. CO-DEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zeng:2021:TRM

[ZLY^{+}21] Zheng Zeng, Shiqiu Liu, Jinglei Yang, Lu Wang, and Ling-Qi Yan. Temporally reliable motion vectors for real-time ray tracing. Computer Graphics Forum, 40 (2):79–90, May 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zell:2022:CFL

[ZM22] E. Zell and R. McDonnell. Compact facial landmark layouts for performance capture. Computer Graphics Forum, 41(2):121–133, May 2022. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhang:2021:DCD

[ZMV^{+}21] Xianyao Zhang, Marco Manzi, Thijs Vogels, Henrik Dahlberg, Markus Gross, and Marios Papas. Deep compositional denoising for high-quality Monte Carlo rendering. Computer Graphics Forum, 40(4):1–13, July 2021. CODEN CG-FODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhang:2022:AFS

[ZOM^{+}22] Xianyao Zhang, Melvin Ott, Marco Manzi, Markus Gross, and Marios Papas. Automatic feature selection for denoising volumetric renderings. Computer Graphics Forum, 41(4):63–77, July 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhenyuan:2023:DAD

[ZPH^{+}23] Liu Zhenyuan, Michal Piovarči, Christian Hafner, Raphaël Charrondière, and Bernd Bickel. Directionality-aware design of embroidery patterns. Computer Graphics Forum, 42(2):397–409, May 2023. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhong:2022:NNS

[ZPIS22] Shengze Zhong, Parinya Punpongsanon, Daisuke Iwai, and Kosuke Sato. NSTO: Neural synthesizing topology optimization for modulated structure generation. Computer Graphics Forum, 41(7):553–566, October 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhang:2020:PWD

[ZQW^{+}20] Ruisong Zhang, Weize Quan, Baoyuan Wu, Zhifeng Li, and Dong-Ming Yan. Pixel-wise dense detector for image inpainting. Computer Graphics Forum, 39 (7):471–482, October 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhao:2020:MPD

[ZSL^{+}20] Hui Zhao, Kehua Su, Chenchen Li, Boyu Zhang, Lei Yang, Na Lei, Xiaoling Wang, Steven J. Gortler, and Xianfeng Gu. Mesh parametrization driven by unit normal flow. Computer Graphics Forum, 39(1):34–49, February 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zheng:2022:EUB

[ZSL^{+}22] Yicun Zheng, Haoran Sun, Xinguo Liu, Hujun Bao, and Jin Huang. Economic upper bound estimation in Hausdorff distance computation for triangle meshes. Computer Graphics Forum, 41(1):46–56, February 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhang:2022:WNV

[ZSS22] P. Zhang, D. Smirnov, and J. Solomon. Wassersplines for neural vector fieldcontrolled animation. Computer Graphics Forum, 41(8):31–41, December 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhuang:2021:RTD

[ZSWL21] Tao Zhuang, Pengfei Shen, Beibei Wang, and Ligang Liu. Real-time denoising using BRDF pre-integration factorization. Computer Graphics Forum, 40(7): 173–180, October 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zheng:2022:GLD

[ZSY^{+}22] Chengyu Zheng, Ding Shi, Xuefeng Yan, Dong Liang, Mingqiang Wei, Xin Yang, Yanwen Guo, and Haoran Xie. GlassNet: Label decoupling-based three-stream neural network for robust image glass detection. Computer Graphics Forum, 41(1):377–388, February 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhou:2020:CCS

[ZTZC20] J. Zhou, C. Tu, D. Zorin, and M. Campen. Combinatorial construction of seamless parameter domains. Computer Graphics Forum, 39(2):179–190, May 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhang:2021:DDE

[ZWCM21] Meng Zhang, Tuanfeng Wang, Duygu Ceylan, and Niloy J. Mitra. Deep detail enhancement for any garment. Computer Graphics Forum, 40(2):399–411, May 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhang:2021:RMV

[ZWG^{+}21] Qitong Zhang, Lei Wang, Linlin Ge, Shan Luo, Taihao Zhu, Feng Jiang, Jimmy Ding, and Jieqing Feng. A robust multi-view system for high-fidelity human body shape reconstruction. Computer Graphics Forum, 40(5):19–31, August 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhang:2022:CRU

[ZWG^{+}22] Wen-Xiang Zhang, Qi Wang, Jia-Peng Guo, Shuangming Chai, Ligang Liu, and Xiao-Ming Fu. Constrained remeshing using evolutionary vertex optimization. Computer Graphics Forum, 41(2):237–247, May 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zohrevandi:2022:DES

[ZWLY22] E. Zohrevandi, C. A. L. Westin, J. Lundberg, and A. Ynnerman. Design and evaluation study of visual analytics decision support tools in air traffic control. Computer Graphics Forum, 41(1):230–242, February 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhou:2020:BRS

[ZWM^{+}20] J. Zhou, M. J. Wang, W. D. Mao, M. L. Gong, and X. P. Liu. Book review: SiamesePointNet: A Siamese Point Network Architecture for Learning 3D Shape Descriptor. Computer Graphics Forum, 39(1):309–321, February 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zohrevandi:2022:EEE

[ZWVL22] E. Zohrevandi, C. A. L. Westin, K. Vrotsou, and J. Lundberg. Exploring effects of ecological visual analytics interfaces on experts’ and novices’ decision-making processes: a case study in air traffic control. Computer Graphics Forum, 41(3): 453–464, June 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhang:2021:NMF

[ZWW^{+}21] Yu-Wei Zhang, Jinlei Wang, Wenping Wang, Yanzhao Chen, Hui Liu, Zhongping Ji, and Caiming Zhang. Neural modelling of flower bas-relief from 2D line drawing. Computer Graphics Forum, 40(6):288–303, September 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhou:2023:NVS

[ZWY23] Shizhe Zhou, Zezu Wang, and Dongwei Ye. Novel view synthesis of transparent object from a single image. Computer Graphics Forum, 42(1):21–32, February 2023. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhang:2020:GHQ

[ZXC^{+}20] Zhe Zhang, Panpan Xu, Jian Chang, Wencheng Wang, Chong Zhao, and Jian Jun Zhang. Generating high-quality superpixels in textured images. Computer Graphics Forum, 39(7):421–431, October 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhu:2020:DKD

[ZXJ^{+}20] Shilin Zhu, Zexiang Xu, Henrik Wann Jensen, Hao Su, and Ravi Ramamoorthi. Deep kernel density estimation for photon mapping. Computer Graphics Forum, 39(4):35–45, July 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhang:2021:WAH

[ZXS^{+}21] Jia-Qi Zhang, Xiang Xu, Zhi-Meng Shen, Ze-Huan Huang, Yang Zhao, Yan-Pei Cao, Pengfei Wan, and Miao Wang. Write-an-animation: High-level text-based animation editing with character-scene interaction. Computer Graphics Forum, 40(7):217–228, October 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhu:2020:GCC

[ZYCF20] Tianyu Zhu, Chunyang Ye, Shuangming Chai, and Xiao-Ming Fu. Greedy cut construction for parameterizations. Computer Graphics Forum, 39(2):191–202, May 2020. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhao:2021:AUD

[ZYP^{+}21] Yong Zhao, Le Yang, Ercheng Pei, Meshia Cédric Oveneke, Mitchel Alioscha-Perez, Longfei Li, Dongmei Jiang, and Hichem Sahli. Action unit driven facial expression synthesis from a single image with patch attentive GAN. Computer Graphics Forum, 40(6):47–61, September 2021. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).

Zhang:2022:LSW

[ZZF^{+}22] Di Zhang, Xiaoya Zhai, Xiao-Ming Fu, Heming Wang, and Ligang Liu. Largescale worst-case topology optimization. Computer Graphics Forum, 41(7):529– 540, October 2022. CODEN CGFODY. ISSN 0167-7055 (print), 1467-8659 (electronic).